### v1

# Optimal Rates for DP-SCO with a Single Epoch and Large Batches

Abstract
The most common algorithms for differentially private (DP) machine learning (ML) are all based on stochastic gradient descent, for example, DP-SGD . These algorithms achieve DP by treating each gradient as an independent private query. However, this independence can cause us to overpay in privacy loss because we don’t analyze the entire gradient trajectory. In this work, we propose a new DP algorithm, which we call Accelerated-DP-SRGD (DP stochastic recursive gradient descent), that enables us to break this independence and only pay for privacy in the gradient difference, i.e., in the new information at the current step. Our algorithm achieves the optimal DP-stochastic convex optimization (DP-SCO) error (up to polylog factors) using only a single epoch over the dataset, and converges at the Nesterov’s accelerated rate.
Our algorithm can be run in at most $\sqrt{n}$ batch gradient steps with batch size at least $\sqrt{n}$, unlike prior work which required $O(n)$ queries with mostly constant batch sizes. To achieve this, our algorithm combines three key ingredients, a variant of stochastic recursive gradients (SRG), accelerated gradient descent, and correlated noise generation from DP continual counting. Finally, we also show that our algorithm improves over existing SoTA on multi-class logistic regression on MNIST and CIFAR-10.
1 Introduction
Differentially private stochastic convex optimization (DP-SCO) and its close counterpart, DP empirical risk minimization (DP-ERM) are two of the most well studied problems in private machine learning [11, 8, 27, 7, 2, 10, 55, 4, 35, 29, 18, 40]. Optimal rates are known for nearly all classes of convex losses, e.g., Lipschitz (and smooth) [8, 27], and Lipschitz with strong convexity [29]. However, there is comparatively little work in optimizing the number of gradient calls [12] required to achieve these rates [35, 27, 29].
In this work, we show that it is possible to obtain optimal DP-SCO error (upto polylog $n,1/\delta$ factors) with a single epoch gradient based algorithm for $\ell_{2}$-Lipschitz and smooth losses. More precisely, we show the following:
There exists a gradient based DP algorithm that makes a single pass over the data set (a.k.a. single epoch  [^4] Formally, a gradient based algorithm $\mathcal{A}(D)\to\{s_{1},\ldots,s_{T}\}$ is called single epoch, if it computes the gradients on any data point $d\in D$ at exactly one state $s_{t}$. Without additional restrictions, a state is allowed to encode arbitrary amount of history, i.e., $s_{t}\supseteq\{s_{1},\ldots,s_{t-1}\}$.) and achieves optimal DP-SCO error (within polylog factors), with $T=n/B$ batch gradient steps and batch size $B=O(\sqrt{n})$ The algorithm uses exactly two gradients per training example [^5] Here, $n$ is the number of training examples..
Our work significantly improves the results of [27, 55] that required at least $T=\Omega(n)$ batch gradient steps (a.k.a. adaptive gradient oracle calls) with batch size $B=n/T$ to reach the optimal DP-SCO error. Further, unlike [27] the privacy guarantee of our algorithm does not need to assume convexity, which makes it empirically applicable in non-convex settings too without violating DP guarantees. We demonstrate this with an empirical comparison on multi-class logistic regression problems (for which the clipped loss is non-convex [48]), where we show improvements over the strong baseline algorithm of DP-FTRL (DP follow the regularized leader) [38, 23, 15, 14, 13] on standard benchmarks. (See Section 2 for more details on DP-SGD and DP-FTRL .)
Problem definition: There is a distribution of examples $\mathcal{D}$ and a loss function $f(\mathbf{x},d)$ where $\mathbf{x}\in\mathcal{C}$ is a model in the constraint set $\mathcal{C}$ and $d$ is an example. We assume for any $d\in supp(\mathcal{D})$, $f$ is convex, $L$-Lipschitz, and $M$-smooth over $\mathcal{C}$. We receive a sequence of $n$ i.i.d. examples $D=\{d_{1},d_{2},\ldots,d_{n}\}\sim\mathcal{D}$. Using these examples, our goal is to choose a model that minimizes the population loss function $F(\mathbf{x})=\mathbb{E}_{d\sim\mathcal{D}}\left[f(\mathbf{x},d)\right]$. We will operate in the single-pass streaming setting: we must process examples in the order they appear, and each example appears only once in this order. Finally, the model we output should satisfy $(\varepsilon,\delta)$-DP with respect to the examples. For simplicity, we use the zero-out notion of DP [46]: if $P$ is the distribution of models our algorithm outputs given $\{d_{1},d_{2},\ldots,d_{n}\}$ and $Q$ is the distribution given the same sequence, except one example $d_{i}$ instead has $\nabla f(\mathbf{x},d_{i})=0$ for all $\mathbf{x}$, then we require $\max_{S}[P(S)-e^{\varepsilon}Q(S)],\max_{S}[Q(S)-e^{\varepsilon}P(S)]\leq\delta$.
Adaptive oracle complexity (a.k.a. batch gradient steps): We enforce our algorithm to be single epoch (2.1). Furthermore, we quantify the computational complexity of the algorithms in terms of two quantities: a) Number of adaptive gradient computations ($T$) the algorithm has to perform, i.e., the current computation depends on the output of the gradient computations in the previous state [50, 53], and b) number of gradient oracle calls made to each data point overall. Our algorithm has $T=O(\sqrt{n})$, and makes exactly two gradient calls per data point. In comparison, [27, 55] has $T=O(n)$[27] makes one gradient call/data point (but relies on convexity for privacy guarantee), [55] makes two gradient calls/data point (same as ours). In practice, adaptive calls are much more expensive [6, 37], hence we focus on reducing adaptive computation cost as opposed to focusing on the gap between one vs two gradient oracle calls per data point.
Designing private learning algorithms that minimize adaptive computations is a well-established problem in the literature [50, 53, 21]. Even in practice, adaptive gradient evaluations are more expensive than gradient calls on multiple data points at the same state, especially in distributed learning settings like federated learning [6, 37]. To the best of our knowledge, this is the first paper to explicitly optimize for the adaptive oracle complexity for the DP-SCO problem. [53] show that in the non-private setting, the adaptive oracle complexity for optimal SCO guarantee is $T=\Omega(n^{1/4})$ (with a batch size of $B=O(n^{3/4})$). This is achieved by Nesterov’s accelerated SGD. It is worth pointing out that Accelerated-DP-SRGD does not match this non-private bound even for constant dimensions and $\varepsilon\to\infty$. The main reason being that we use stochastic recursive gradients (SRGs) [45, 55] (instead of minibatch gradients) for privacy purposes, and the variance of the SRGs overwhelm the variance due to DP noise at large $\varepsilon$’s. We leave it as an open problem to match the rates in [53] under DP.
Design of Accelerated-DP-SRGD (DP stochastic recursive gradient descent): In terms of design,
Accelerated-DP-SRGD is based on a careful adaptation of Nesterov’s accelerated stochastic gradient descent [44, 9], combined with the ideas of stochastic recursive gradients (SRG) [45, 55], and matrix factorization (MF) based private continual counting mechanisms [24, 19, 23, 42, 26, 33, 34, 5].
Recall, Nesterov’s acceleration can be expressed as the coupling of two gradient descent based algorithms with different learning rates. More formally, the following are the update rules [^6] Here, $\Pi_{\mathcal{C}}(\cdot)$ is the $\ell_{2}$-projection onto to the constraint set, $\{\eta_{t}\}$ and $\beta$ are scaling parameters, and $\{\tau_{t}\}$ are the coupling parameters.: a) $\mathbf{z}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{z}_{t}-\frac{\eta_{t}}{\beta}\nabla_{t}\right)$, b) $\mathbf{y}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}\right)$, and c) $\mathbf{x}_{t+1}\leftarrow\left(1-\tau_{t+1}\right)\mathbf{y}_{t+1}+\tau_{t+1}\mathbf{z}_{t+1}$. In our implementation of the Nesterov’s method, we will use a DP method to approximate the gradient computed at $\mathbf{x}_{t}$ computed on a minibatch $\mathcal{B}_{t}$ of size $B$: $\nabla_{t}(\mathbf{x}_{t})=\frac{1}{B}\sum\limits_{i\in\mathcal{B}_{t}}\nabla f(\mathbf{x}_{t};d_{i})$. We use the idea of stochastic recursive gradients (SRGs) from [45, 55] to first approximate $\nabla_{t}$ as $\eta_{t}\nabla_{t}\approx\Delta_{t}+\Delta_{t-1}+\ldots$, where each $\Delta_{t}=\eta_{t}\nabla_{t}(\mathbf{x}_{t})-\eta_{t-1}\nabla_{t}(\mathbf{x}_{t-1})$. We set $\eta_{t}=t$, $\beta\approx M\cdot\max\left\{\frac{n^{3/2}}{B^{2}},\frac{n}{B}\right\}$, and $\tau_{t}=\frac{\eta_{t}}{\sum\limits_{\tau\leq t}\eta_{\tau}}$. Following are the few crucial differences from [55]: a) We use a truly accelerated method to achieve the improved batch gradient steps, as opposed to tail-averaged online convex optimization methods (which does not include the term $\mathbf{y}_{t}$), b) because of acceleration, we cannot tolerate higher values of $\eta_{t}$. In particular the results of [55] hold with $\eta_{t}=t^{k}$ (for any $k\geq 1$). To bound the sensitivity of $\Delta_{t}$ for privacy, we need to slow down the learning by choosing $\beta\approx M\sqrt{n}$ (corresponding to the iterations $T=n/B\approx\sqrt{n}$), as opposed $\beta=M$ in the non-private case.
With this choice of parameters, we first show that $\Delta_{t}$ has sensitivity $O(1/B)$. Hence, we can use the standard MF based continual counting mechanisms (Section 2.2) to compute DP variants of each $\eta_{t}\nabla_{t}$ with low error. We finally get an SCO error of the form: $\widetilde{O}\left(\frac{B^{2}}{n^{2}}+\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\right)$, with $B$ being the minibatch size. The first two terms match the non-private optimal SCO error from [41].
Comparison to most related prior work [27, 55]: Two of the prior works which are most related to our work are of [27, 55]. Here, we position the ideas in this paper in regard to these works. For $d,\varepsilon=\Theta(1)$ both these algorithms use $T=O(n)$ batch gradient steps, with batch size $\approx O(1)$. Our main algorithm Accelerated-DP-SRGD , uses Nesterov’s accelerated method [44, 9] as the baseline algorithm to improve on the SoTA to $T=B=O(\sqrt{n})$. One natural question is whether the algorithms in the prior work with small modifications can achieve our adaptive oracle complexity.
[27] either relies on privacy amplification by iteration [28] or localization based methods (which requires adding adaptive $\ell_{2}$-regularization). In either case choosing a batch size larger than $O(\max\{\sqrt{d/n},1\})$ in majority of the batch gradient steps, will result in a sub-optimal DP-SCO error. In comparison, for $M=\Theta(1)$ our algorithm does not have any dependence of $d$ on the batch gradient complexity. Furthermore, [27] heavily uses convexity for their privacy analysis. In contrast, Accelerated-DP-SRGD
can be used seamlessly in non-convex settings without compromising on the DP guarantees (via clipping of $\Delta_{t}$).
The algorithm of [55] and their analysis relies on the access to a low-regret online convex optimization algorithm (OCO) [31]. We know that the regret guarantees of OCO methods cannot be improved beyond $\Omega(1/\sqrt{T})$ even under smoothness [31]. Hence, it is not obvious how the online-to-batch framework [55] can be used to achieve the number of batch gradient steps of our algorithm. One might think of replacing the OCO algorithm in [55] with a SGD method tailored to smooth losses [12, Theorem 6.3] to get around the reduced convergence rate of $\Omega(1/\sqrt{T})$. In Section 4.3 we show that unfortunately, because of the variance introduced by the SRGs, in the absence of an accelerated SGD method, the optimal batch size is still $O(1)$.
Other related work: We discuss some of the prior work [17, 10, 2, 30]. Although the settings/problems in these papers are different than the one we study, all are important for the broader question of adaptive oracle complexity for optimal DP-SCO.
[17] provides the best known gradient oracle complexity for optimal DP-SCO, for non-smooth convex losses under $\ell_{2}$-geometry. While their algorithm is not single-epoch, their overall gradient complexity (i.e., the number of gradient oracle calls) is $\widetilde{O}\left(\min\left\{n,\frac{n^{2}}{d}\right\}+\min\left\{\frac{nd^{2}}{\varepsilon},n^{4/3}\varepsilon^{1/3}\right\}\right)$. This result is incomparable to our result, since we rely on smoothness in our analysis, but our overall gradient oracle complexity is an improved $\widetilde{O}\left(\sqrt{n}\right)$. Under constant $\varepsilon$, under all regimes of the dimensionality where the DP-SCO guarantee is meaningful i.e., $d\leq n^{2}$, our gradient oracle complexity is superior. We leave the possibility of removing the smoothness assumption in our paper while obtaining the same (or better) gradient oracle complexity as an open problem. It is also worth mentioning that the results in [17] heavily rely on privacy amplification by sampling [36, 11], and convexity of the loss function for privacy. Our work does not rely on either. At the heart of [17], is a new variance bounded stochastic gradient estimator ReSQue. It is an important open question if this estimator can be used to improve the variance properties of our stochastic recursive gradients (SRGs).
[10, 2, 30] have looked at various extensions to the $\ell_{2}$-DP-SCO problem. For example, DP-SCO guarantees under the loss functions being $\ell_{1}$-Lipschitz, or SCO properties of second-order-stationary points. These results, while being important for the DP-SCO literature, are not comparable to our line of work.
Practical enhancements, and empirical evaluation: Along with demonstrating the theoretical benefits of Accelerated-DP-SRGD , we demonstrate that the core ideas in our algorithm can be used to improve well-optimized benchmark results for DP learning [38, 23, 15]. While we chose $\eta_{t}=t$ in Accelerated-DP-SRGD to get optimal DP-SCO bounds, in practice we found setting $\eta_{t}=e^{\lambda t}$ and optimizing for $\lambda$ provides the best results. With this reparameterization, we make the crucial observation that if $\lambda$ is very high and $\tau_{t}=1$, then the Nesterov’s update simplifies to standard SGD update. Hence, there exists a parameterization under which Accelerated-DP-SRGD is equivalent to DP-FTRL  (differentially private follow-the-regularized-leader) [38, 23]. Hence, with more optimized parameters we can only get a better algorithm. The exact algorithm we use for empirical evaluation is given in Algorithm 3 (Algorithm DP-SRG-MEMF ).
In our empirical evaluation, we provide results for multi-class logistic regression on MNIST and CIFAR-10 as a proof-of-concept. We show that Accelerated-DP-SRGD provides a utility benefit over the prior SoTA [15] for the no-amplification [^7] Prior works [38, 43] mentioned that in distributed learning environments like federated learning, privacy amplification by sampling [36] is undesirable, and even when it is applicable, the true privacy guarantee depends on the precise implementation [16]. Hence, we focus on the evaluation of our algorithms when it does not have access to any form of privacy amplification. (i.e., streaming) setting that has been optimized for over a series of work [38, 23, 15, 14, 13]. A detailed empirical exploration of Accelerated-DP-SRGD and its practical variant on large-scale practical tasks is left for future exploration.
1.1 Our Contributions
  1. 1.
    (Accelerated) DP stochastic recursive gradient descent (Accelerated-DP-SRGD ) (Section 3): In this work we provide a differentially private variant of Nesterov’s accelerated stochastic gradient descent [41, 9]. We use the idea of stochastic recursive gradients (SRGs) [45, 55] to approximate the gradient update at each stage, and use DP matrix factorization based continual counting machinery [23, 24, 19] to approximate the SRGs.
  2. 2.
    Obtaining optimal DP-SCO errors with improved number of batch gradient steps (Theorem 4.6): We show that Accelerated-DP-SRGD can achieve optimal DP-SCO errors for smooth convex losses, in a single epoch with batch gradient steps of $T\approx\sqrt{n}$ (batch size $B=n/T)$, and exactly two gradient calls per data point. We required two technical ideas to achieve this: a) reducing the learning rate of non-private Nesterov’s method to account for the noise added due to privacy, and b) show that there exists appropriate choice of parameters for the algorithm under which the SRGs are low-sensitive (i.e., does not change too much by changing one data sample in the data set $D$).
  3. 3.
    Surpassing SoTA on benchmark tasks (Section 5): Inspired by our Accelerated-DP-SRGD , we create a heuristic (via exponential re-weighting of gradient differences) that empirically outperforms the highly optimized class of DP-FTRL algorithms [38, 23, 15] on multi-class logistic regression on MNIST and CIFAR-10.
2 Background on (Private) Learning
2.1 Definition for Single Epoch Training
Definition 2.1 (Single epoch).
A gradient based algorithm $\mathcal{A}(D)\to\{s_{1},\ldots,s_{T}\}$ is called single epoch, if it computes the gradients on any data point $d\in D$ at exactly one state $s_{t}$. Without additional restrictions, a state is allowed to encode arbitrary amount of history, i.e., $s_{t}\supseteq\{s_{1},\ldots,s_{t-1}\}$.
Notice that 2.1 does not include any restriction on the memory of the algorithm. While stating our results in the single epoch setting, we additionally incorporate the batch size, and the total number gradients evaluated per data point.
2.2 DP-Matrix Factorization (DP-MF ) and Continual Counting
Continual counting under DP [19, 24] is one of the most important primitives in the DP literature. Given a sequence of adaptively chosen data vectors $D=\{\mathbf{v}_{1},\ldots,\mathbf{v}_{T}\}$ (with each $\left\|\mathbf{v}_{i}\right\|_{2}\leq 1$), the objective is to output $\mathbf{z}=\left[\sum\limits_{\tau\leq t}\mathbf{v}_{\tau}\right]_{t\in[T]}$ while preserving DP. It is well-known that one can output $\mathbf{z}_{\texttt{priv}}$ s.t. $\mathbb{E}[\|\mathbf{z}-\mathbf{z}_{\texttt{priv}}\|_{\infty}]=\text{polylog}(T)$. This can be achieved via the classic binary tree mechanism [19, 24, 32], see Lemma 4.3. We use the binary tree primitive for our theoretical guarantees. A recent line of work [23, 42, 26, 33, 34, 5] optimizes the error up to additive constant via factorization of a lower-triangular matrix of all ones $\mathbf{A}=\mathbf{B}\cdot\mathbf{C}$. For our empirical results, we appeal to the computational approach of [23] for finding the best factorizations ($\mathbf{B}$ and $\mathbf{C}$) w.r.t. the error in continual counting under DP.
2.3 DP-SGD : DP stochastic gradient descent
DP-SGD  [47, 11, 1] is a standard procedure for optimizing a loss function $\mathcal{L}(\mathbf{x};D)=\frac{1}{n}\sum\limits_{i\in[n]}\nabla\ell(\mathbf{x};d_{i})$ under DP. The algorithm is as follows: $\mathbf{x}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{x}_{t}-\eta\cdot\left(\nabla_{t}+\mathbf{b}_{t}\right)\right)$, where $\nabla_{t}=\frac{1}{|\mathcal{B}|}\sum\limits_{d\in\mathcal{B}}\texttt{clip}\left(\nabla\ell(\mathbf{x}_{t};d),L\right)$ is the gradient computed on a minibatch of samples in $\mathcal{B}\subseteq D$ at $\mathbf{x}_{t}$, $\mathbf{b}_{t}$ is a well-calibrated (independent) spherical Gaussian noise to ensure DP, and $\texttt{clip}\left(\mathbf{v},L\right)=\mathbf{v}\cdot\min\left\{\frac{L}{\left\|\mathbf{v}\right\|_{2}},1\right\}$. One can show that under appropriate choice of parameters DP-SGD achieves the optimal DP-SCO error [8, 7].
2.4 DP-FTRL : DP follow the regularized leader
DP-FTRL  [49, 38, 23] is another standard approach for optimization, where in its simplest form, the state update is exactly that of DP-SGD when $\mathcal{C}=\mathbb{R}^{d}$, except the noise in two different rounds $\mathbf{b}_{t},\mathbf{b}_{t^{\prime}}$ are not independent. Instead (in the one-dimensional setting) $(\mathbf{b}_{1},\mathbf{b}_{2},\ldots\mathbf{b}_{T})\sim\mathcal{N}(0,\left(\mathbf{C}^{\top}\mathbf{C}\right)^{-1})$, where $\mathbf{C}$ is computed via DP-MF . While DP-FTRL has strong empirical properties comparable to DP-SGD , it is unknown whether it can achieve optimal DP-SCO error unless under additional restrictions (e.g., realizability [3]).
3 Our Algorithm
In the following we provide our main algorithm (Algorithm 1, Accelerated-DP-SRGD ) treating the addition of noise for DP as a black box, and giving a utility analysis as a function of the noise. In Section 4.2 we will specify the noise mechanism and its privacy guarantees, and show the resulting utility analysis gives the desired optimal rates.
Our algorithm is a DP variant of Nesterov accelerated SGD [44, 9] using stochastic recursive gradients (SRGs): similarly to [45, 55] we maintain $\nabla_{t}$, an unbiased estimate of $\nabla F(\mathbf{x}_{t})$, and use $\nabla_{t}$ and batch $t+1$ to obtain $\nabla_{t+1}$. Our algorithm incorporates noise $\mathbf{b}_{t}$ for DP. We first give a utility analysis for a fixed value of the noise $\mathbf{b}_{t}$, and then in Section 4.2 we will specify how to set the noise to simultaneously achieve good privacy and utility. As mentioned in Section 1, we will use the DP-MF  machinery (Section 2.2) to set the $\mathbf{b}_{t}$’s.
It is worth mentioning that all three aspects of our algorithm are crucial: a) SRGs allow ensuring that the $\ell_{2}$-sensitivity of the gradient estimate at time $t$ is a factor of $1/t$, lower than that of the direct minibatch gradient ($\frac{1}{B}\sum\limits_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x}_{t};d)$), b) Noise addition via DP-MF  allows the standard deviation of the noise $\mathbf{b}_{t}$ to grow only polylogarithmically in $T$, as opposed to $\sqrt{T}$ via simple composition analysis [25], and c) Accelerated SGD allows appropriately balancing the variance introduced by the SRG estimates, and that due to the i.i.d. sampling of the minibatches from the population. In fact in Section 4.3 we argue that it is necessary to have an accelerated SGD algorithm to achieve a single epoch optimal DP-SCO algorithm with batch size $B=\omega(1)$, if we were to use the SRGs as the gradient approximators.
Algorithm 1 Accelerated-DP-SRGD
Inputs: Steps: $T$, batches: $\{\mathcal{B}_{t}\}_{0\leq t<T}$ of size $B$, learning rates: $\{\eta_{t}\}_{0\leq t<T}$, scaling $\beta$, interpolation params: $\{\tau_{t}\}_{t\in[T]}$, constraint set: $\mathcal{C}$ with $0\in\mathcal{C}$.
1:$\mathbf{x}_{0},\mathbf{z}_{0}\leftarrow 0$.
2:for $t\in\{0,\ldots,T-1\}$ do
3:     $\Delta_{t}\leftarrow\frac{1}{B}\sum_{d\in\mathcal{B}_{t}}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)]$
4:     $\nabla_{t}\leftarrow\frac{\sum_{i\leq t}\Delta_{i}}{\eta_{t}}$ (equivalently, $\nabla_{t}\leftarrow\frac{\Delta_{t}}{\eta_{t}}+\frac{\eta_{t-1}}{\eta_{t}}\cdot\nabla_{t-1}$)
5:     $\mathbf{z}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{z}_{t}-\frac{\eta_{t}}{\beta}\nabla_{t}+\mathbf{b}_{t}\right)$, where $\mathbf{b}_{t}$ is the noise added to ensure privacy.
6:     $\mathbf{y}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}+\frac{1}{\eta_{t}}\cdot\mathbf{b}_{t}\right)$.
7:     $\mathbf{x}_{t+1}\leftarrow\left(1-\tau_{t+1}\right)\mathbf{y}_{t+1}+\tau_{t+1}\mathbf{z}_{t+1}$.
8:end for
4 Utility and Privacy Analysis
In this section, we provide the detailed utility analysis and privacy analysis of Algorithm 1.
4.1 Utility Analysis
Here, we provide the excess population risk guarantee for any choice of the privacy noise random variables $\{\mathbf{b}_{t}\}$’s. We only assume that the variance (over the sampling of the data samples $d\sim\mathcal{D}$) in any individual term in the SRGs (i.e., $\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$) is bounded by $S^{2}$.
Eventually in Theorem 4.6 we will combine 4.1 below with the privacy analysis of Algorithm 1 to arrive at the final DP-SCO guarantee, with the appropriate choice of the batch gradient steps, and the batch size. In particular, there we instantiate the noise random variables $\{\mathbf{b}_{t}\}$’s with the (by now standard) binary tree mechanism [19, 24].
Theorem 4.1 (Excess population risk; Simplification of Theorem 4.7).
Fix some setting of the $\mathbf{b}_{t}$ such that $\left\|\mathbf{b}_{t}\right\|_{2}\leq b_{max}$ for all $t$. For an appropriate setting of $\eta_{t},\tau_{t}$ and $\beta\geq M$, suppose $\mathbb{E}[\left\|\Delta_{t}-\mathbb{E}[\Delta_{t}]\right\|_{2}^{2}]\leq S^{2}$. Then, over the randomness of the batches $\mathcal{B}$, the following is true for Algorithm 1:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=$
$O\left(\frac{SL+S^{2}}{\beta\sqrt{T}}+\frac{S\left\|\mathcal{C}\right\|_{2}}{T^{1.5}}+\frac{(S+L)b_{max}}{T}+\frac{\beta b_{max}\left\|\mathcal{C}\right\|_{2}}{T^{2}}+\frac{\beta}{T^{2}}\left\|\mathcal{C}\right\|_{2}^{2}\right).$
We defer the proof to Section 4.4.
Intrpretation of 4.1: To understand the bound in Theorem 4.1, we can view it as three groups of terms. The first group is $\frac{SL+S^{2}}{\beta\sqrt{T}}+\frac{S\left\|\mathcal{C}\right\|_{2}}{T^{1.5}}$, which is the error due to stochasticity of the gradients. The second is $\frac{(S+L)b_{max}}{T}+\frac{\beta b_{max}\left\|\mathcal{C}\right\|_{2}}{T^{2}}$, which is the error due to noise added for privacy. The last is $\frac{\beta}{T^{2}}\left\|\mathcal{C}\right\|_{2}^{2}$, which is the optimization error, i.e. the error of the algorithm without stochasticity or DP noise. Notice that the optimization error matches that of the non-stochastic part of non-private Nesterov’s acceleration [41]. Without privacy, the stochastic part of Nesterov’s acceleration depends on $\frac{V}{\beta\sqrt{T}}$ [41], where $V$ bounds stochasticity in evaluating the population gradient: $\mathbb{E}_{\mathcal{B}_{t}}\left[\left\|\frac{1}{|\mathcal{B}_{t}|}\sum\limits_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x};d)-\nabla F(\mathbf{x};d)\right\|_{2}^{2}\right]\leq V^{2}$. Morally, the parameter $S$ in 4.1 can be compared to $V$. However, $S$ encodes the additional variance due to stochastic recursive gradients. Modulo, the difference between $S$ and $V$4.1 cleanly decomposes the non-private stochastic Nesterov’s utility and the additional error due to the explicit noise introduced due to DP.
Proof overview: The proof is deferred to Section 4.4. We give a high-level overview here. We define the following potential. Let $\eta_{0:t}=\sum\limits_{i=0}^{t}\eta_{i}$, then
$\Phi_{t}=\eta_{0:t-1}\left(F(\mathbf{y}_{t})-F(\mathbf{x}^{*})\right)+2\beta\left\|\mathbf{z}_{t}-\mathbf{x}^{*}\right\|_{2}^{2}.$
The proof has three parts: a) We bound the potential drop in each iteration with terms arising from the stochasticity of sampling the $\{f_{t}\}$’s and the privacy noise $\{\mathbf{b}_{t}\}$’s, b) Use telescopic summation of the potential drops to bound the final potential $\Phi_{T}$ as a function of stochastic error and DP noise across all steps, and c) convert the potential bound to a loss bound to obtain the final error guarantee. Our proof closely follows the proof structure of [9] and draws inspiration from [54] in terms of parameter choices.
Independent batch gradients: In Theorem 4.1 we analyzed Algorithm 1 where the $\nabla_{t}$ are computed using SRG. The analysis can easily be extended to the standard setting where each $\nabla_{t}$ is an independent batch gradient. We give this analysis in Section 4.4.3. To the best of our knowledge, a direct potential-based analysis of stochastic accelerated gradient descent has not appeared in the literature before besides [51]. Our analysis is simpler than that of [51], so it may be of independent interest.
4.2 Privacy analysis
Given Theorem 4.1, our goal now is to bound $S$, which is equivalent to bounding the sensitivity of the $\Delta_{t}$, and then to show that there is a noise distribution for which $b_{max}$ is sufficiently small (with high probability) and we achieve $(\varepsilon,\delta)$-DP. We can then set $\beta$ to minimize the error bound, and determine the largest batch size that doesn’t cause the error bound to increase.
We first prove the sensitivity bound on each $\Delta_{t}$. The sensitivity bound requires a bound on the max noise we add in any iteration. The proof of sensitivity follows a similar structure to that of [55, Lemma 15]. However, because the gradients $\nabla f_{t}(\mathbf{x}_{t})$ are evaluated at a different point than that in [55], the exact argument is different.
Claim 4.2 (Bounding $\ell_{2}$-sensitivity).
In Algorithm 1, suppose for all $t$, $\left\|\mathbf{b}_{t}\right\|_{2}\leq b_{\max}$. Then for $\eta_{t}=t+1$, $\beta\geq 2MT$ for any $d\in\mathcal{D}$, $t\in\{0,1,\ldots,T-1\}$
$\left\|\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)\right\|_{2}\leq 2(Mb_{\max}+2M\left\|\mathcal{C}\right\|_{2}+L).$
Proof.
For convenience let $\Delta_{t}(d)=\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$. The bound holds by Lipschitzness for $t=0$. We prove the bound for $t>0$ by induction. We have:
$\Delta_{t}(d)$ $=\eta_{t}\left(\nabla f(\mathbf{x}_{t},d)-\nabla f(\mathbf{x}_{t-1},d)\right)+\left(\eta_{t}-\eta_{t-1}\right)\nabla f(\mathbf{x}_{t},d)$
$\Rightarrow\left\|\Delta_{t}(d)\right\|_{2}$ $\leq\eta_{t}\cdot M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}+\left|\eta_{t}-\eta_{t-1}\right|\cdot L.$ (1)
We now bound each of the terms in the R.H.S. of Equation 1 independently. Since $\eta_{t}=t+1$, $\eta_{t}-\eta_{t-1}=1$. In the following, we bound $\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$. We have,
$\eta_{0:t}\cdot\mathbf{x}_{t}$ $=\eta_{0:t-1}\cdot\mathbf{y}_{t}+\eta_{t}\cdot\mathbf{z}_{t}$
$\Leftrightarrow\eta_{0:t}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $=\left\|\eta_{0:t-1}(\mathbf{y}_{t}-\mathbf{x}_{t-1})+\eta_{t}(\mathbf{z}_{t}-\mathbf{x}_{t-1})\right\|_{2}$
$\leq\eta_{0:t-1}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|_{2}+\eta_{t}\left\|\mathbf{z}_{t}-\mathbf{x}_{t-1}\right\|_{2}$
$\leq\eta_{0:t-1}\left\|\frac{\nabla_{t-1}}{\beta}+\frac{\mathbf{b}_{t-1}}{\eta_{t}}\right\|_{2}+\eta_{t}\left\|\mathcal{C}\right\|_{2}$
$\Rightarrow\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{\eta_{0:t-1}}{\eta_{0:t}}\cdot\frac{\left\|\nabla_{t-1}\right\|_{2}}{\beta}+\frac{\eta_{0:t-1}\left\|\mathbf{b}_{t-1}\right\|_{2}}{\eta_{t}\cdot\eta_{0:t}}+\frac{\eta_{t}}{\eta_{0:t}}\left\|\mathcal{C}\right\|_{2}$
$\leq\frac{\sum_{i\leq t-1}\sum_{d\in\mathcal{B}_{i}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta(t+1)B}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}\left\|\mathcal{C}\right\|_{2}$
$\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}\left\|\mathcal{C}\right\|_{2}$
$\Rightarrow\eta_{t}\cdot M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{\eta_{t}M\cdot\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta}+Mb_{\max}+2M\left\|\mathcal{C}\right\|_{2}$
$\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{2}+Mb_{\max}+2M\left\|\mathcal{C}\right\|_{2}.$ (2)
Plugging Section 4.2 and the fact that $\eta_{t}-\eta_{t-1}=1$ into Equation 1, we have the following:
$\left\|\Delta_{t}(d)\right\|_{2}\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{2}+Mb_{max}+2M\left\|\mathcal{C}\right\|_{2}+L.$ (3)
By induction, this completes the proof of 4.2. ∎
A crucial aspect of 4.2 is that conditioned on $b_{max}$, the sensitivity is independent of $t$, even if the norm of the quantity that is being approximated (i.e., $\eta_{t}\nabla f(\mathbf{x}_{t};d)$) scales with $t$.
Deciding the noise scale for $(\varepsilon,\delta)$-DP: We now show how to sample $\mathbf{b}_{t}$ such that Algorithm 1 satisfies $(\varepsilon,\delta)$-DP while achieving strong utility guarantees. Each $\nabla_{t}$ is a (rescaling of a) prefix sum of the $\Delta_{t}$ values. So, choosing a noise mechanism that does not add too much noise to the gradients in Algorithm 1 is equivalent to privatizing the list of prefix sums $\{\Delta_{1},\Delta_{1}+\Delta_{2},\ldots,\Delta_{1}+\ldots+\Delta_{T}\}$, where by Claim 4.2, we know the sensitivity of each $\Delta_{t}$ is constant with respect to $t$. This is a well studied problem known as private continual counting, and a standard solution to this problem is the binary tree mechanism [24].
Lemma 4.3.
Let $\mathbf{g}_{i}=\sum_{j\leq i}\Delta_{i}$ for $i\in[T]$ and suppose the $\Delta_{i}$ are computed such that between any two adjacent datasets, only one $\Delta_{i}$ changes, and it changes by $\ell_{2}$-norm at most $C$. Then for any $T\geq 2$ there is a $\mu$-GDP algorithm (i.e., an algorithm which satisfies any DP guarantee satisfied by the Gaussian mechanism with sensitivity $\mu$ and variance 1) that gives an unbiased estimate of each $\mathbf{g}_{i}$ such that with probability $1-\delta$ all $\mathbf{g}_{i}$’s estimates have $\ell_{2}$-error at most $\frac{4C\log^{1.5}T\sqrt{d\ln(2T/\delta)}}{\mu}$.
Proof.
For completeness, we define the binary tree mechanism and state its guarantees here. For $k=0,1,2,\ldots,\lceil\log T\rceil$, $j=1,3,5,\ldots T/2^{k}-1$, the binary tree mechanism computes $s_{j,k}=\sum_{i=(j-1)\cdot 2^{k}+1}^{j\cdot 2^{k}}\Delta_{i}+\xi_{j,k},\xi_{j,k}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,\sigma^{2}\mathbb{I})$. An unbiased estimate of any prefix sum can be formed by adding at most $\lceil\log T\rceil$ of the $s_{j,k}$. For example, the sum of $\Delta_{1}$ to $\Delta_{7}$ has an unbiased estimator using at most $\lceil\log 7\rceil=3$ of the $s_{j,k}$, $s_{1,2}+s_{3,1}+s_{7,0}$.
Each $\Delta_{i}$ contributes to at most one $s_{j,k}$ for each value of $k$, i.e. at most a total of $1+\lceil\log T\rceil$ values of $s_{j,k}$. So, if $\left\|\Delta_{i}\right\|_{2}\leq C$ for all $i$, then with $\sigma=C\sqrt{1+\lceil\log T\rceil}/\mu$ we can release all $s_{j,k}$ with $\mu$-GDP. Furthermore, by a standard Gaussian tail bound with probability $21-\delta/$ 2 we have that $\left\|\xi_{i}\right\|_{2}\leq\sigma\sqrt{2d\ln(2T/\delta)}$ for all $i$. Since each $\mathbf{g}_{i}$ has an unbiased estimator formed by adding at most $\lceil\log T\rceil$ of the $s_{j,k}$, the maximum additive error is then at most
$\lceil\log T\rceil\max_{i}\left\|\xi_{i}\right\|_{2}\leq 2C\log T\sqrt{2d(1+\lceil\log T\rceil)\ln(2T/\delta)}/\mu\leq 4C\log^{1.5}T\sqrt{d\ln(2T/\delta)}/\mu.$
4.3 gives us a high probability bound on the max noise, which we can then plug into 4.2 to show that if the dimension (and thus the norm of the noise) is not too large, the noise term in 4.2 is not dominating:
Corollary 4.4.
Suppose we compute $\mathbf{b}_{i}$ in Algorithm 1 using the $\mu$-GDP binary tree mechanism (Lemma 4.3) with $C=\frac{4L+8M\left\|\mathcal{C}\right\|_{2}}{B\beta}$. Then for $d\leq\frac{B^{2}\beta^{2}\mu^{2}}{64M^{2}\log^{3}(T)\ln(4T/\delta)}$ and the parameter choices in Claim 4.2, with probability $21-\delta/$ 2, each $\Delta_{t}$ has $\ell_{2}$-sensitivity at most $\leq 4L+8M\left\|\mathcal{C}\right\|_{2}$ for all $i$.
Our eventual choice of $\beta$ will be proportional to $n^{3/2}/B^{2}$, i.e. the upper bound on $d$ is proportional to $n^{3}\varepsilon^{2}/B^{2}$. We can rewrite this as an upper bound on the batch size: $B\leq n^{1.5}\varepsilon/\sqrt{d}$. Ignoring dependence on logarithmic terms and $\delta$, when $d\geq n^{2}\varepsilon^{2}$, no $(\varepsilon,\delta)$-DP algorithm can achieve expected excess loss, $o(L\left\|\mathcal{C}\right\|_{2})$, which is achieved by any point in $\mathcal{C}$. So this upper bound is vacuous in the setting where $d\leq n$ and $\varepsilon=\Omega(1)$, and never worse than $\sqrt{n}$ for interesting settings of $\delta$.
We can now convert the sensitivity bound and the privacy guarantees of the binary tree mechanism into a privacy guarantee for Algorithm 1:
Corollary 4.5.
Suppose we compute $\mathbf{b}_{i}$ in Algorithm 1 using the binary tree mechanism (Lemma 4.3) with $\sigma=\frac{(8\sqrt{2}L+16\sqrt{2}M\left\|\mathcal{C}\right\|_{2})\sqrt{\log(T)\ln(2.5/\delta)}}{\varepsilon B\beta}$ on $\Delta_{i}$. Then for $d\leq\frac{B^{2}\beta^{2}\varepsilon^{2}}{128M^{2}\log^{3}(T)\ln(4T/\delta)\ln(2.5/\delta)}$ and the parameter choices in Claim 4.2, Algorithm 1 satisfies $(\varepsilon,\delta)$-DP.
Proof.
If we run Algorithm 1 but instead clip $\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$ for each $t$ and $d$ in $\mathcal{B}_{t}$ to norm at most $4L+8M\left\|\mathcal{C}\right\|_{2}$ if its norm exceeds $4L+8M\left\|\mathcal{C}\right\|_{2}$, still using the binary tree mechanism to generate the $\mathbf{b}_{i}$, then Algorithm 1 satisfies $\varepsilon/\sqrt{2\ln(2.5/\delta)}$-GDP which implies it satisfies $(\varepsilon,\delta/2)$-DP. Algorithm 1 without clipping is within total variation distance $\delta/2$ of the version with clipping by Corollary 4.4, i.e. Algorithm 1 satisfies $(\varepsilon,\delta$)-DP. ∎
Finally, we plug the error guarantees of the binary tree mechanism into Theorem 4.7 to get our main result:
Theorem 4.6.
[Simplification of Theorem 4.10 in Section 4.4] Algorithm 1 with batch size $B=\sqrt{n}\cdot\min\{1,M\}$ and $T=n/B$ (i.e., a single epoch over a dataset of size $n$), and an appropriate choice of $\eta_{t},\beta,\sigma$ is $(\varepsilon,\delta)$-DP and has the following excess population loss guarantee:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left((L\left\|\mathcal{C}\right\|_{2}+M\left\|\mathcal{C}\right\|_{2}^{2})\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right)\right).$
4.3 Unaccelerated Variant
A natural question is whether an unaccelerated variant of Accelerated-DP-SRGD can also achieve optimal DP-SCO rates with batch size $\omega(1)$. We give a sketch that the unaccelerated version requires batch size $O(1)$. Suppose we use $\Delta_{t}=c_{t}\nabla f(\mathbf{x}_{t},\mathcal{B}_{t})-c_{t-1}\nabla f(\mathbf{x}_{t-1},\mathcal{B}_{t})$ as our gradient differences and $\nabla_{t}=\frac{\sum_{i\leq t}\Delta_{i}}{c_{t}}$ as our stochastic recursive gradients. For simplicity we will consider using SGD with a fixed learning rate $\eta$ as the underlying optimization algorithm, assume $L=\left\|\mathcal{C}\right\|_{2}=M=1$ and assume $c_{t}$ grows at most polynomially, which implies $\sum_{i\leq t}c_{i}^{2}\approx tc_{t}^{2}$. Similarly to Claim 4.2, we can show the sensitivity of each $\Delta_{t}$ is roughly $c_{t}\eta$. Suppose we use $\eta\approx\sqrt{B/T}=B/\sqrt{n}$, which is standard for interpolation between full-batch GD and SGD with $B=1$ (see e.g. Section 6 of [12]). We compute the vector variance $\mathbf{Var}\left(\mathbf{v}\right):=\mathbb{E}[\left\|\mathbf{v}-\mathbb{E}[\mathbf{v}]\right\|_{2}^{2}]$ of $\nabla_{t}$:
$\mathbf{Var}\left(\nabla_{t}\right)=\mathbf{Var}\left(\frac{\sum_{i\leq t}\Delta_{t}}{c_{t}}\right)=\frac{\sum_{i\leq t}\mathbf{Var}\left(\Delta_{t}\right)}{c_{t}^{2}}=\frac{\sum_{i\leq t}c_{i}^{2}\eta^{2}}{c_{t}^{2}}=\frac{t\eta^{2}}{B}=\frac{tB}{n}.$
For $t=\Omega(T)=\Omega(n/B)$, the variance is $\Omega(1)$. In other words, regardless of the batch size, the variance of the gradients for the majority of iterations is comparable to that of vanilla SGD with batch size 1. SGD with batch size 1, i.e. with variance $O(1)$ rather than $O(1/B)$, requires $T=n$ iterations to converge, hence in the single-epoch setting we need $B=O(1)$.
4.4 Deferred Proofs
4.4.1 Proof of Theorem 4.1
Theorem 4.1 is given by plugging $\eta_{t}={t+1}$ into the following more general theorem, dropping lower-order terms, setting $S=Q/\sqrt{B}$, and taking the asymptotic bound.
Theorem 4.7 (Excess population risk).
Fix some setting of the $\mathbf{b}_{t}$. Let $\eta_{t},\beta$ be such that $\eta_{t}$ is non-decreasing, $\eta_{t}^{2}\leq 4\eta_{0:t}$, and $\beta\geq M$. For any $\mathbf{x}_{t},\mathbf{x}_{t-1}$, and for all $d\in\mathcal{D}$, assume
$\left\|(\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d))-\mathbb{E}_{d}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)]\right\|_{2}\leq Q.$
Then, over the randomness of the i.i.d. sampling of the loss functions $\{f(\cdot;d_{t})\}$ from $\mathcal{D}$, the following is true for Algorithm 1:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}\eta_{0:t}}+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)$
$+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
Proof.
Potential drop: We define
$\Phi_{t}=\eta_{0:t-1}\left(F(\mathbf{y}_{t})-F(\mathbf{x}^{*})\right)+2\beta\left\|\mathbf{z}_{t}-\mathbf{x}^{*}\right\|_{2}^{2}.$ (4)
We first show that the growth in the potential $\Phi_{t}$ at each iteration of Algorithm 1 is bounded. In particular, if the stochasticity of the algorithm is removed, i.e., if we set $Q=0$ and $\{\mathbf{b}_{t}\}$’s to zero, then the potential growth is non-positive. (This is consistent with the full-gradient version of Nesterov’s acceleration [9].)
Lemma 4.8 (Potential drop).
Let $\mathbf{q}_{t}=\nabla_{t}-\nabla F(\mathbf{x}_{t})$. Assuming $\eta_{t}^{2}\leq 4\eta_{0:t}$, the potential drop is bounded by the following (where $\Delta\Phi_{t}=\Phi_{t+1}-\Phi_{t}$):
$\mathbb{E}\left[\Delta\Phi_{t}\right]\leq\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{t+1}}{\sqrt{B}\eta_{0:t}}\right.$
$\left.+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right).$
Telescopic summation and averaging: By telescoping the potential differences $\Delta\Phi_{t}$’s and linearity of expectation we have the following over the sampling of the functions $\{f_{t}\}$’s:
$\mathbb{E}[\Phi_{T}-\Phi_{0}]\leq$ $\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{t+1}}{\sqrt{B}\eta_{0:t}}\right.$
$\left.+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right).$ (5)
Therefore, from Equation 5 and definition of $\Phi_{t}$, as desired we have:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{t+1}}{\sqrt{B}\eta_{0:t}}\right.$
$\left.+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
This completes the proof of Theorem 4.7. ∎
Proof of 4.8.
We have
$\Delta\Phi_{t}$ $=\underbrace{\eta_{0:{t}}(F(\mathbf{y}_{t+1})-F(\mathbf{x}_{t}))}_{A}-\underbrace{\eta_{0:t-1}(F(\mathbf{y}_{t})-F(\mathbf{x}_{t}))+\eta_{t}(F(\mathbf{x}_{t})-F(\mathbf{x}^{*}))}_{B}$
$+\underbrace{2\beta\left(\left\|\mathbf{z}_{t+1}-\mathbf{x}^{*}\right\|_{2}^{2}-\left\|\mathbf{z}_{t}-\mathbf{x}^{*}\right\|_{2}^{2}\right)}_{C}.$ (6)
By convexity, we have the following linearization bound on term $B$ in Equation 6:
$B$ $\leq-\eta_{0:t-1}\left\langle\nabla F(\mathbf{x}_{t}),\mathbf{y}_{t}-\mathbf{x}_{t}\right\rangle+\eta_{t}\left\langle\nabla F(\mathbf{x}_{t}),\mathbf{x}_{t}-\mathbf{x}^{*}\right\rangle$ (7)
Now, $(1-\tau_{t})(\mathbf{y}_{t}-\mathbf{x}_{t})=\tau_{t}(\mathbf{x}_{t}-\mathbf{z}_{t})$. Setting, $\tau_{t}=\frac{\eta_{t}}{\eta_{0:t}}$ in Equation 7 (and using the definition of $\mathbf{q}_{t}$), we have the following:
$B\leq\eta_{t}\left\langle\nabla F(\mathbf{x}_{t}),\mathbf{z}_{t}-\mathbf{x}^{*}\right\rangle=\eta_{t}\left\langle\nabla_{t},\mathbf{z}_{t}-\mathbf{x}^{*}\right\rangle-\eta_{t}\left\langle\mathbf{q}_{t},\mathbf{z}_{t}-\mathbf{x}^{*}\right\rangle.$ (8)
Now, we bound the term $C$ in Equation 6 as follows. Let $\mathbf{z}^{\prime}_{t+1}=\mathbf{z}_{t}-\frac{\eta_{t}}{2\beta}\nabla_{t}+\mathbf{b}_{t}$.
$C$ $=2\beta\left(\left\|\mathbf{z}_{t+1}-\mathbf{x}^{*}\right\|_{2}^{2}-\left\|\mathbf{z}_{t}-\mathbf{x}^{*}\right\|_{2}^{2}\right)$
$=2\beta\left(2\left\langle\mathbf{z}_{t+1}-\mathbf{z}_{t},\mathbf{z}_{t+1}-\mathbf{x}^{*}\right\rangle-\left\|\mathbf{z}_{t+1}-\mathbf{z}_{t}\right\|_{2}^{2}\right)$
$\leq 2\beta\left(\left\langle-\frac{\eta_{t}}{\beta}\nabla_{t}+\mathbf{b}_{t},\mathbf{z}_{t+1}-\mathbf{x}^{*}\right\rangle-\left\|\mathbf{z}_{t+1}-\mathbf{z}_{t}\right\|_{2}^{2}\right)$ (9)
$=-2\beta\left(\left\langle\frac{\eta_{t}}{\beta}\nabla_{t},\mathbf{z}_{t+1}-\mathbf{x}^{*}\right\rangle+\left\|\mathbf{z}_{t+1}-\mathbf{z}_{t}\right\|_{2}^{2}\right)+2\beta\left\langle\mathbf{b}_{t},\mathbf{z}_{t+1}-\mathbf{x}^{*}\right\rangle.$ (10)
In Equation 9, we used the Pythagorean property of convex projection. Combining Equation 10 and Equation 8, we have the following:
$B+C$ $\leq\eta_{t}\left\langle\nabla_{t},\mathbf{z}_{t}-\mathbf{z}_{t+1}\right\rangle-2\beta\left\|\mathbf{z}_{t+1}-\mathbf{z}_{t}\right\|_{2}^{2}+2\beta\left\langle\mathbf{b}_{t},\mathbf{z}_{t+1}-\mathbf{x}^{*}\right\rangle-\eta_{t}\left\langle\mathbf{q}_{t},\mathbf{z}_{t}-\mathbf{x}^{*}\right\rangle$ (11)
Using 4.9 below, and the fact that $\tau_{t}\leq 1$, we have the potential drop in 4.8 bound by the following.
$\Delta\Phi_{t}\leq-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}-\eta_{t}\left\langle\mathbf{q}_{t},\mathbf{z}_{t}-\mathbf{x}^{*}\right\rangle$ (12)
Recall, by definition $\mathbf{y}_{t+1}=\operatorname*{arg\,min}\limits_{\mathbf{y}\in\mathcal{C}}\left\langle\mathbf{q}_{t}+\nabla F(\mathbf{x}_{t})-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{2}\left\|\mathbf{y}-\mathbf{x}_{t}\right\|_{2}^{2}$. Let $\mathbf{y}_{t+1}^{\prime}=\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}+\frac{1}{\eta_{t}}\cdot\mathbf{b}_{t}$ be the unconstrained minimizer, and let $J(\mathbf{y})=\left\langle\mathbf{q}_{t}+\nabla F(\mathbf{x}_{t})-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}\right\rangle+\frac{\beta}{2}\left\|\mathbf{y}\right\|_{2}^{2}$. By definition, $J(\mathbf{y}_{t+1})\geq J(\mathbf{y}^{\prime}_{t+1})+\frac{\beta}{2}\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\prime}_{t+1}\right\|_{2}^{2}$. Therefore,
$\left\langle\mathbf{q}_{t}+\nabla F(\mathbf{x}_{t})-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}_{t+1}-\mathbf{y}^{\prime}_{t+1}\right\rangle$ $\geq\frac{\beta}{2}\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\prime}_{t+1}\right\|_{2}^{2}$
$\Rightarrow\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\prime}_{t+1}\right\|_{2}$ $\leq\frac{2\left\|\nabla_{t}\right\|_{2}}{\beta}+\frac{2\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}.$ (13)
We now apply (13) and Cauchy-Schwarz inequality to (12):
$\Delta\Phi_{t}\leq$ $-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}-\eta_{t}\left\langle\mathbf{q}_{t},\mathbf{z}_{t}-\mathbf{x}^{*}\right\rangle$
$=$ $-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{y}_{t+1}^{\prime}\right\rangle-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}^{\prime}-\mathbf{x}_{t}\right\rangle+\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}$
$+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}-\eta_{t}\left\langle\mathbf{q}_{t},\mathbf{z}_{t}-\mathbf{x}^{*}\right\rangle$
$=$ $-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{y}_{t+1}^{\prime}\right\rangle-\eta_{0:t}\left\langle\mathbf{q}_{t},-\frac{1}{\beta}\nabla_{t}+\frac{1}{\eta_{t}}\mathbf{b}_{t}\right\rangle+\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}$
$+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}-\eta_{t}\left\langle\mathbf{q}_{t},\mathbf{z}_{t}-\mathbf{x}^{*}\right\rangle$
$\leq$ $\eta_{0:t}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathbf{y}_{t+1}-\mathbf{y}_{t+1}^{\prime}\right\|_{2}+\frac{\eta_{0:t}}{\beta}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\nabla_{t}\right\|_{2}+\frac{\eta_{0:t}}{\eta_{t}}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathbf{b}_{t}\right\|_{2}+\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}$
$+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}+\eta_{t}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathbf{z}_{t}-\mathbf{x}^{*}\right\|_{2}$
$\leq$ $\eta_{0:t}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathbf{y}_{t+1}-\mathbf{y}_{t+1}^{\prime}\right\|_{2}+\frac{\eta_{0:t}}{\beta}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\nabla_{t}\right\|_{2}+\frac{\eta_{0:t}}{\eta_{t}}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathbf{b}_{t}\right\|_{2}+\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}$
$+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}+\eta_{t}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}.$ (14)
Now, we observe that:
$\mathbf{q}_{t}=\nabla_{t}-\nabla F(\mathbf{x}_{t})=\frac{\sum_{i\leq t}\sum_{d\in\mathcal{B}_{i}}((\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d))-\mathbb{E}_{d}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)])}{B\eta_{t}}.$
The sum $\sum_{i\leq t}\sum_{d\in\mathcal{B}_{i}}((\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d))-\mathbb{E}_{d}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)])$ is a mean-zero vector martingale with $B(t+1)$ terms each with norm at most $Q$ with probability 1 by the assumptions of the theorem. So we have
$\mathbb{E}\left[\left\|\sum_{i\leq t}((\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d))-\mathbb{E}_{d}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)])\right\|_{2}^{2}\right]\leq Q^{2}B(t+1)$
and thus $\mathbb{E}[\left\|\mathbf{q}_{t}\right\|_{2}^{2}]\leq\frac{Q^{2}(t+1)}{B\eta_{t}^{2}}$, $\mathbb{E}[\left\|\mathbf{q}_{t}\right\|_{2}]\leq\frac{Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}$ by Jensen’s inequality. We also unconditionally have $\left\|\mathbf{q}_{t}\right\|_{2}\leq Q(t+1)/\eta_{t}$ and thus $\left\|\nabla_{t}\right\|_{2}\leq L+Q(t+1)/\eta_{t}$ by triangle inequality. Plugging these bounds into (14) we have:
$\mathbb{E}[\Delta\Phi_{t}]\leq\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{t+1}}{\sqrt{B}\eta_{0:t}}\right.$
$\left.+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right).$
Claim 4.9.
Assuming $\eta_{t}^{2}\leq 4\eta_{0:t}$. We have the following:
$\eta_{0:t}(F(\mathbf{y}_{t+1})-F(\mathbf{x}_{t}))+\eta_{t}\left\langle\nabla_{t},\mathbf{z}_{t}-\mathbf{z}_{t+1}\right\rangle-2\beta\left\|\mathbf{z}_{t+1}-\mathbf{z}_{t}\right\|_{2}^{2}\leq$
$-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\left(\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}+\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}\right).$
Proof of 4.9.
By smoothness we have,
$F(\mathbf{y}_{t+1})-F(\mathbf{x}_{t})\leq\left\langle\nabla_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle-\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{M}{2}\left\|\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\|_{2}^{2}$ (15)
By definition $\mathbf{y}_{t+1}=\operatorname*{arg\,min}\limits_{\mathbf{y}\in\mathcal{C}}\left(\left\langle\nabla_{t}-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{2}\left\|\mathbf{y}-\mathbf{x}_{t}\right\|_{2}^{2}\right)$. Therefore, for any $\mathbf{v}\in\mathcal{C}$, the following is true:
$\left\langle\nabla_{t}-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{2}\left\|\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\|_{2}^{2}\leq\left\langle\nabla_{t}-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{v}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{2}\left\|\bf\mathbf{v}-\mathbf{x}_{t}\right\|_{2}^{2}$
$\Leftrightarrow\left\langle\nabla_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle\leq\left\langle\nabla_{t},\mathbf{v}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{\eta_{t}}\cdot\left\langle\mathbf{b}_{t},\mathbf{y}_{t+1}-\mathbf{v}\right\rangle+\frac{\beta}{2}\left\|\bf\mathbf{v}-\mathbf{x}_{t}\right\|_{2}^{2}-\frac{\beta}{2}\left\|\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\|_{2}^{2}.$ (16)
Plugging Equation 16 into Equation 15 and using $\beta\geq M$ we have the following:
$F(\mathbf{y}_{t+1})-F(\mathbf{x}_{t})$ $\leq\left\langle\nabla_{t},\mathbf{v}-\mathbf{x}_{t}\right\rangle-\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{\eta_{t}}\cdot\left\langle\mathbf{b}_{t},\mathbf{y}_{t+1}-\mathbf{v}\right\rangle$
$-\frac{\beta}{2}\left\|\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\|_{2}^{2}+\frac{\beta}{2}\left\|\mathbf{v}-\mathbf{x}_{t}\right\|_{2}^{2}.$ (17)
Define $\mathbf{v}=(1-\tau_{t})\mathbf{y}_{t}+\tau_{t}\mathbf{z}_{t+1}$. Which implies $\mathbf{v}-\mathbf{x}_{t}=\tau_{t}(\mathbf{z}_{t+1}-\mathbf{z}_{t})$. Substituting $\mathbf{v}$ in Equation 17, we have the following:
$F(\mathbf{y}_{t+1})-F(\mathbf{x}_{t})$ $\leq\tau_{t}\left\langle\nabla_{t},\mathbf{z}_{t+1}-\mathbf{z}_{t}\right\rangle-\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{\eta_{t}}\cdot\left\langle\mathbf{b}_{t},\mathbf{y}_{t+1}-\mathbf{v}\right\rangle$
$-\frac{\beta}{2}\left\|\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\|_{2}^{2}+\frac{\beta\tau_{t}^{2}}{2}\left\|\mathbf{z}_{t+1}-\mathbf{z}_{t}\right\|_{2}^{2}.$ (18)
Assume, $\frac{\eta_{t}^{2}}{\eta_{0:t}}\leq 4$. Combining Equation 18 and Equation 15, and setting $\tau_{t}=\frac{\eta_{t}}{\eta_{0:t}}$, we have the following:
$\eta_{0:t}(F(\mathbf{y}_{t+1})-F(\mathbf{x}_{t}))+\eta_{t}\left\langle\nabla_{t},\mathbf{z}_{t}-\mathbf{z}_{t+1}\right\rangle-2\beta\left\|\mathbf{z}_{t+1}-\mathbf{z}_{t}\right\|_{2}^{2}\leq$ (19)
$-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{\tau_{t}}\left\langle\mathbf{b}_{t},\mathbf{y}_{t+1}-\mathbf{v}\right\rangle+\frac{\beta}{2}{\left({\eta_{0:t}\cdot\tau_{t}^{2}}-4\right)\left\|\mathbf{z}_{t+1}-\mathbf{z}_{t}\right\|_{2}^{2}}-\frac{\beta\eta_{0:t}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\|_{2}^{2}$
$\leq-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{\tau_{t}}\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathbf{y}_{t+1}-\mathbf{x}_{t}-\tau_{t}(\mathbf{z}_{t+1}-\mathbf{z}_{t})\right\|_{2}$
$\leq-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{\tau_{t}}\left\|\mathbf{b}_{t}\right\|_{2}\left(\frac{L}{\beta}+\frac{\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\tau_{t}\left\|\mathcal{C}\right\|_{2}\right)$
$\leq-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\left(\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}+\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}\right).$ (20)
This completes the proof of 4.9. ∎
4.4.2 Proof of Theorem 4.6
We give the more formal statement of Theorem 4.6 and its proof here:
Theorem 4.10.
Algorithm 1 with batch size
$B\leq\min\left\{\sqrt{n},\frac{(L+2M\left\|\mathcal{C}\right\|_{2})n^{1.5}\varepsilon}{4\sqrt{2}\left\|\mathcal{C}\right\|_{2}\sqrt{d}\log^{1.5}(T)\sqrt{\ln(4T/\delta)\ln(2.5)/\delta}}\right\}$
and $T=n/B$ (i.e., a single pass over a dataset of size $n$), $\eta_{t}=t+1$, $\beta=M+\frac{(8L+16M\left\|\mathcal{C}\right\|_{2})n^{3/2}}{\left\|\mathcal{C}\right\|_{2}B^{2}}$, $\sigma=\frac{(8\sqrt{2}L+16\sqrt{2}M\left\|\mathcal{C}\right\|_{2})\sqrt{\log(T)\ln(2.5/\delta)}}{\varepsilon B\beta}$ is $(\varepsilon,\delta)$-DP and has the following excess population loss guarantee:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=$
$O\left(\frac{M\left\|\mathcal{C}\right\|_{2}^{2}B^{2}}{n^{2}}+\frac{(L+M\left\|\mathcal{C}\right\|_{2})\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}+\frac{(L+\frac{M\left\|\mathcal{C}\right\|_{2}}{\sqrt{B}})\left\|\mathcal{C}\right\|_{2}\sqrt{d}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right).$
In particular, if we choose $B=\min\{\sqrt{n},Mn^{1.5}\varepsilon/(\sqrt{d}\cdot\text{polylog}(n/B,1/\delta))\}$ then Algorithm 1 requires
$\max\{\sqrt{n},\sqrt{d}\cdot\text{polylog}(n/B,1/\delta)/M\varepsilon\sqrt{n}\}\}$
batch gradients and:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left((L\left\|\mathcal{C}\right\|_{2}+M\left\|\mathcal{C}\right\|_{2}^{2})\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right)\right).$
Proof.
Note that the upper bound on $B$ gives $\beta\geq 16Mn/B\geq 2MT$, satisfying the assumption in Claim 4.2, and also gives $d$ satisfying the assumption in Corollary 4.5. Let $Q=8L+16M\left\|\mathcal{C}\right\|_{2}=O(L+M\left\|\mathcal{C}\right\|_{2})$. We plug the tail bound of Lemma 4.3 into the error bound of Theorem 4.7, use a failure probability of $1/T^{3}$ in that bound (the increase in the expected error due to this low-probability event is $O(L\left\|\mathcal{C}\right\|_{2}/T^{3})$, which is a lower order term in our error bound) and choose $\eta_{t}=t+1,\beta=\frac{QT^{3/2}}{\left\|\mathcal{C}\right\|_{2}\sqrt{B}}$, let:
$c_{DP}=\frac{\sqrt{d\log^{3}(T)\log(T/\delta)\log(1/\delta)}}{\varepsilon}$
We can assume that $d=O(n^{2}\varepsilon^{2})$, i.e. $c_{DP}=O(n\cdot{\rm polylog}\left(T,1/\delta\right)$, as if $d=\omega(n^{2}\varepsilon^{2})$ then any point in $\mathcal{C}$ achieves the desired excess loss bound. Now we split the bound in Theorem 4.7 into six terms and bound them individually. Specifically, from Theorem 4.7 we have:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\underbrace{\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}}_{E_{1}(t)}+\underbrace{\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}}_{E_{2}(t)}+\underbrace{\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}\eta_{0:t}}}_{E_{3}(t)}+\underbrace{\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}}_{E_{4}(t)}+\underbrace{\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}}_{E_{5}(t)}\right)$
$+\underbrace{\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}}_{E_{6}}.$
Now, using (i) $Q/\left\|C\right\|_{2}=O(M)$, (ii) $n=BT$, (iii) $B\leq\sqrt{n}$ (and thus $T\geq\sqrt{n}$) and (iv) $\left\|\mathbf{b}_{t}\right\|_{2}=O\left(\frac{Qc_{DP}}{B\beta}\right)$ by Lemma 4.3 and our choice of $\sigma$, we can bound the contribution of each of $E_{1},E_{2},E_{3},E_{4},E_{5},E_{6}$ individually:
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{1}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right)=O\left(\frac{Q\sqrt{T}(L+\frac{Q}{\sqrt{B}})}{\sqrt{B}\beta}\right)$
$=O\left(\frac{\left\|\mathcal{C}\right\|_{2}(L+\frac{Q}{\sqrt{B}})}{T}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{T}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}\right),$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{2}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}\right)=O\left(Q\left\|\mathbf{b}_{t}\right\|_{2}\right)=O\left(\frac{Q^{2}c_{DP}}{B\beta}\right)$
$=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{\sqrt{B}T^{3/2}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{T\sqrt{n}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{n}\right),$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{3}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}\eta_{0:t}}\right)=\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}\right),$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{4}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}\right)=O\left(\frac{1}{T^{2}}\sum\limits_{t=0}^{T-1}\frac{Qc_{DP}\left\|\mathcal{C}\right\|_{2}}{B}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{n}\right),$
$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{5}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)=O\left(\frac{1}{T^{2}}\sum\limits_{t=0}^{T-1}\beta\left(\frac{Qc_{DP}}{B\beta}\right)^{2}\right)=O\left(\frac{Q^{2}c_{DP}^{2}}{B^{2}T\beta}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}^{2}}{B^{3/2}T^{5/2}}\right)$
$=$ $O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}B}{n^{3/2}}{\rm polylog}\left(T,1/\delta\right)\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{n}{\rm polylog}\left(n/B,1/\delta\right)\right),$
$E_{6}=$ $\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}=O\left(\frac{QT^{3/2}}{\left\|\mathcal{C}\right\|_{2}\sqrt{B}}\cdot\frac{1}{T^{2}}\cdot\left\|\mathcal{C}\right\|_{2}^{2}\right)=O\left(\frac{M\left\|\mathcal{C}\right\|_{2}^{2}}{\sqrt{n}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}\right).$
Combining the asymptotic bounds, we get our final bound:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left(Q\left\|\mathcal{C}\right\|_{2}\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\text{polylog}(n/B,1/\delta)\right)\right).$
4.4.3 Analysis of Stochastic Accelerated Gradient Descent using Independent Gradients
Theorem 4.11.
If we use $\nabla_{t}=\frac{1}{B}\sum_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x}_{t},d)$ instead in Accelerated-DP-SRGD , if $\eta_{t}^{2}\leq 4\eta_{0:t}$ and $\beta\geq M$ then we have:
$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq\frac{1}{\eta_{0:T}}\sum_{i=0}^{T-1}\left[\eta_{0:t}\left(\frac{8L^{2}}{\sqrt{B}\beta}+\frac{5L\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)\right]+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
Proof.
Each $\mathbf{q}_{t}$ is now the average of $B$ random vectors which have norm at most $2L$. So, we have $\mathbb{E}_{\mathbf{q}_{t}}\left[\left\|\mathbf{q}_{t}\right\|_{2}^{2}|\mathbf{x}_{t}\right]\leq\frac{4L^{2}}{B}$. By Jensen’s inequality, this gives $\mathbb{E}_{\mathbf{q}_{t}}\left[\left\|\mathbf{q}_{t}\right\|_{2}|\mathbf{x}_{t}\right]\leq\frac{2L}{\sqrt{B}}$. Now, using the fact that $\mathbf{q}_{t}$ is mean zero conditioned on $\mathbf{x}_{t}$, we have (using $\mathbf{y}^{\prime}_{t+1}$ as defined in the proof of Theorem 4.7):
$\mathbb{E}_{\mathbf{q}_{t}}\left[-\left\langle\mathbf{q}_{t},\mathbf{y}^{\prime}_{t+1}-\mathbf{x}_{t}\right\rangle|\mathbf{x}_{t}\right]=\frac{1}{\beta}\mathbb{E}_{\mathbf{q}_{t}}\left[\left\|\mathbf{q}_{t}\right\|_{2}^{2}|\mathbf{x}_{t}\right]\leq\frac{4L^{2}}{B\beta}.$ (21)
Combining Equation 13Equation 21, and the fact that now $\left\|\nabla_{t}\right\|_{2}\leq L$ unconditionally, we have the following:
$-\eta_{0:t}\mathbb{E}_{\mathbf{q}_{t}}\left[\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle|\mathbf{x}_{t}\right]\leq\eta_{0:t}\left(\frac{8L^{2}}{\sqrt{B}\beta}+\frac{4L\left\|\mathbf{b}_{t}\right\|_{2}}{\sqrt{B}\eta_{t}}\right).$ (22)
Using the fact that $\mathbf{q}_{t}$ is mean-zero and $\mathbb{E}_{\mathbf{q}_{t}}\left[\left\|\mathbf{q}_{t}\right\|_{2}|\mathbf{x}_{t}\right]\leq\frac{2L}{\sqrt{B}}$ we then have the following bound on the potential change by plugging (22) into (12):
$\mathbb{E}_{\mathbf{q}_{t}}\left[\Delta\Phi_{t}|\mathbf{x}_{t}\right]\leq\eta_{0:t}\left(\frac{8L^{2}}{\sqrt{B}\beta}+\frac{5L\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)$ (23)
We now proceed similarly to Theorem 4.7. ∎
5 Empirical Evaluation
To evaluate the practicality of DP-SRG algorithms, we perform empirical evaluations comparing DP-SRG algorithms to their analogs using batch gradients. We consider the multi-pass streaming setting where we can take multiple passes over the dataset in a fixed order, but this order is treated as public information, i.e. we cannot randomize the order and assume privacy amplification. This aligns with many practical settings for DP optimization, where the standard amplification techniques of Poisson sampling / uniformly random shuffling are difficult to implement [38, 16].
In all settings, we compare to the baseline DP-MEMF algorithm (Algorithm 2) of [15] which is the SoTA in the absence of privacy amplification by sampling. To combine SRG with DP-MEMF , we give an algorithm DP-SRG-MEMF as Algorithm 3. In DP-SRG-MEMF we take the underlying first-order optimization algorithm as a black-box. For simplicity of presentation we state DP-SRG-MEMF for a one-dimensional optimization problem, but it is easy to extend to a multi-dimensional problem. As long as $\max_{j\in[b]}\left\|\sum_{i=0}^{k-1}\mathbf{C}_{:,ib+j}\right\|_{2}\leq 1$, DP-SRG-MEMF satisfies $\rho$-zCDP, i.e. $(\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP [15]. [15] provides guidance for choosing a matrix $\mathbf{C}$. When clipping and noising are no-ops, DP-SRG-MEMF computes gradients using the SRG framework and passes them to the first-order optimization method. DP-SRG-MEMF privatizes the gradients by using correlated noise like DP-MEMF , but we apply the correlated noise to $\Delta_{t}$, the “update” term in SRG, rather than the gradients themselves (as was done in DP-MEMF ).
Algorithm 2 Multi-Epoch Matrix Factorization (DP-MEMF )
Inputs: Epochs $k$, batches $\mathcal{B}_{1},\mathcal{B}_{2},\ldots,\mathcal{B}_{b}$ of size $B$, lower-triangular matrix $\mathbf{C}\in\mathbb{R}_{\geq 0}^{kb\times kb}$ such that $\max_{j\in[b]}\left\|\sum_{i=0}^{k-1}\mathbf{C}_{:,ib+j}\right\|_{2}\leq 1$, target zCDP $\rho$, clip norm $C$.
1:$\mathbf{Z}\sim N(0,\frac{1}{2\rho}\mathbb{I}_{kb\times kb})$
2:for $i$ in $\{0,1,\ldots,k-1\}$ do
3:     for $j$ in $[b]$ do
4:         $\nabla_{ik+j}\leftarrow\frac{1}{B}\sum_{d\in\mathcal{B}_{j}}\text{clip}(\nabla\ell(\mathbf{x}_{ik+j};d),C)+(\mathbf{C}^{-1}\mathbf{Z})_{ik+j}$
5:         Use $\nabla_{ik+j}$ as gradient in first-order optimization algorithm to update $\mathbf{x}_{ik+j}$ to $\mathbf{x}_{ik+j+1}$.
6:     end for
7:end for
Algorithm 3 SRG-Multi-Epoch Matrix Factorization (DP-SRG-MEMF )
Inputs: Epochs $k$, batches $\mathcal{B}_{1},\mathcal{B}_{2},\ldots,\mathcal{B}_{b}$ of size $B$, lower-triangular matrix $\mathbf{C}\in\mathbb{R}_{\geq 0}^{kb\times kb}$ such that $\max_{j\in[b]}\left\|\sum_{i=0}^{k-1}\mathbf{C}_{:,ib+j}\right\|_{2}\leq 1$, target zCDP $\rho$, clip norm $C$, decay sequence $\{c_{t}\}_{t\in[kb]}$.
1:$\mathbf{Z}\sim N(0,\frac{1}{2\rho}\mathbb{I}_{kb\times kb})$
2:for $i$ in $\{0,1,\ldots,k-1\}$ do
3:     for $j$ in $[b]$ do
4:         $\Delta_{ik+j}\leftarrow\frac{1}{B}\sum_{d\in\mathcal{B}_{j}}\text{clip}(\nabla\ell(\mathbf{x}_{ik+j};d)-c_{ik+j}\nabla\ell(\mathbf{x}_{ik+j-1};d),C)$
5:         $\widetilde{\Delta}_{ik+j}\leftarrow\Delta_{ik+j}+(\mathbf{C}^{-1}\mathbf{Z})_{ik+j}$
6:         $\nabla_{ik+j}=c_{ik+j}\nabla_{ik+j-1}+\widetilde{\Delta}_{ik+j}$
7:         Use $\nabla_{t}+(\mathbf{C}^{-1}\mathbf{Z})_{ik+j}$ as gradient in first-order optimization algorithm to update $\mathbf{x}_{ik+j}$ to $\mathbf{x}_{ik+j+1}$.
8:     end for
9:end for
5.1 Logistic Regression
We compare DP-MEMF and DP-SRG-MEMF on multiclass logistic regression on the MNIST [20] and CIFAR10 [39] datasets. For both we load from tensorflow_datasets and use the pre-defined train/test split. We fix a batch size of 500 for each dataset, and train for $k=1$ and $6$ epochs respectively, the number of epochs we observed were needed for convergence in the non-private setting. We target $(0.1,10^{-6})$ and $(2,10^{-6})$-DP respectively. We use SGD with a standard momentum of $0.9$ as the underlying first-order algorithm for both DP-MEMF and DP-SRG-MEMF . For both datasets and both algorithms, we sweep the learning rate and clip norm, and pick the best performing parameters per algorithm. For the choice of $c_{t}$ in DP-SRG-MEMF , we fix $c_{t}=c$ for a constant $c$ less than 1. This is analogous to choosing $\eta_{t}=c^{t}$ in Accelerated-DP-SRGD . We sweep over the values of $c$ and find that $c=e^{-5/2}\approx 1/12$ works well.
For the choice of $\mathbf{C}$, in [15], $\mathbf{C}$ is chosen to minimize $\|\mathbf{A}\mathbf{C}^{-1}\|_{F}$ under the constraint
$\max_{j\in[b]}\left\|\sum_{i=0}^{k-1}\mathbf{C}_{:,ib+j}\right\|_{2}\leq 1$, where $\mathbf{A}$ is the lower-triangular all ones matrix. The motivation for this choice of $\mathbf{A}$ is that in SGD, $\mathbf{x}_{i}$ are equal to $\mathbf{x}_{0}$ plus the $i$th row of $\mathbf{A}\mathbf{G}$ where $\mathbf{G}$ is the matrix whose rows are the gradients. Since we add noise $\mathbf{C}^{-1}\mathbf{Z}$ to the gradients, $\|\mathbf{A}\mathbf{C}^{-1}\|_{F}$ gives some measure of the total error of the algorithm due to noise. However, in SGD with momentum, we instead have that $\mathbf{x}_{i}$ is equal to $\mathbf{x}_{0}$ plus the $i$th row of $\mathbf{M}\mathbf{A}\mathbf{G}$, where $\mathbf{M}$ is some function of the momentum parameter. Taking this one step further, if $\mathbf{D}$ is the matrix whose rows are the $\Delta_{i}$ in DP-SRG-MEMF , the $i$th row of $\mathbf{x}_{i}$ is now equal to $\mathbf{x}_{0}$ plus the $i$th row of $\mathbf{M}\mathbf{A}\mathbf{L}\mathbf{D}$, where $\mathbf{L}$ is some function of the decay parameter $c_{t}$. While [15] observed that using $\mathbf{C}$ in DP-MEMF optimized for $\mathbf{A}$ is usually as good as or better than optimizing for $\mathbf{M}\mathbf{A}$ in practice, it is unclear whether the same is true for DP-SRG-MEMF . So, we try two different choices of $\mathbf{C}$ for both algorithms: $\mathbf{C}$ optimized to minimize $\|\mathbf{A}\mathbf{C}^{-1}\|_{F}$the “ones” workload), and $\mathbf{C}$ optimized to minimize $\|\mathbf{M}\mathbf{A}\mathbf{C}^{-1}\|_{F}$ for DP-MEMF and $\|\mathbf{M}\mathbf{A}\mathbf{L}\mathbf{C}^{-1}\|_{F}$ for DP-SRG-MEMF ( the “true” workload).
Algorithm, Workload MNIST Acc. @ $(0.1,10^{-6})$-DP CIFAR Acc. @ $(2.0,10^{-6})$-DP
DP-MEMF , ones 82.555 $\pm$ .144 35.680 $\pm$ .166
DP-MEMF , true 83.593 $\pm$ .123 35.272 $\pm$ .246
DP-SRG-MEMF , ones 82.456 $\pm$ .143 36.854 $\pm$ .123
DP-SRG-MEMF , true 83.753 $\pm$ .121 36.816 $\pm$ .113
Non-private 91.530 $\pm$ .042 37.971 $\pm$ .076
Figure 1: Average test accuracy on MNIST/CIFAR for the different approaches with 96% CI.
In Figure 1 we compare the test accuracy of DP-MEMF and DP-SRG-MEMF , using the two different choices of $\mathbf{C}$, on the two tasks. Using the code base of [15], we are able to generate the $\mathbf{C}$ matrices in under 2 minutes per matrix on CPU. For each task, one training run takes less than 10 minutes to run on CPU. Each reported accuracy is the average over 100 training runs. We see that DP-SRG-MEMF always outperforms DP-MEMF . For CIFAR the absolute improvement is reasonably large for the given task, despite the modest choice of $c$. Furthermore, depending on the task, the choice of $\mathbf{C}$ can heavily affect the performance of the algorithm. Interestingly, even though we look at logistic regression tasks for both datasets, for MNIST choosing $\mathbf{C}$ optimized for the “true” workload is much better for both algorithms, whereas for CIFAR choosing $\mathbf{C}$ optimized for the “ones” workload is slightly better for both algorithms. It is an interesting question to better understand how the dataset affects the best choice of $\mathbf{C}$ in DP-MEMF and its variants.
Note on not comparing to methods specific to convex losses: The task we consider is technically a convex problem, and some of the methods specific to convex losses are (potentially) applicable to this problem [40, 52, 28]. We do not compare to them because of the following reason: Because we do not assume any apriori bound on the norm of the gradients, we clip the gradients to a fixed norm. Unfortunately, once clipping is involved, multi-class logistic loss does not any remain a well-defined function [48, Section 5.3], let alone a convex function.
6 Conclusion and Future Work
In this work, we demonstrated an optimal DP-SCO algorithm (Accelerated-DP-SRGD ); a single epoch first-order method that can operate on much larger batch sizes ($B\approx\sqrt{n}$ or larger) with $T=n/B$ batch gradient steps. It is interesting to note that our algorithm does not match the optimal non-private batch gradient complexity for SCO, achieved via accelerated SGD method [53, 41], even when $\varepsilon\to\infty$. In Accelerated-DP-SRGD , we use stochastic recursive gradients (SRGs) instead of the standard minibatch gradient to reduce the $\ell_{2}$-sensitivity by factor of $\approx 1/t$ for privacy purposes. The variance introduced by SRGs become the dominant effect even when $\varepsilon\to\infty$, and lower bounds the number of batch gradient steps the algorithm needs to make. An important question that evolve from our work is whether we can match the optimal non-private batch gradient steps, and still achieve optimal DP-SCO error? A related question is whether our rates can be matched or beaten by an algorithm that uses standard gradients, and in particular one gradient query per example. Such an algorithm is more likely to match the non-private number of gradient steps as $\varepsilon\rightarrow\infty$.
In the current paper, we did not explore the applicability of the stochastic recursive gradients approach beyond standard (accelerated) SGD based methods. We leave the exploration of the possible gains in adaptive optimization settings like AdaGrad [22, 4] as an open problem.
As mentioned earlier, our algorithm is applicable to non-convex settings without compromising on privacy guarantees. In this work, we experimented with (clipped) logistic regression as a proof-of-concept. We leave a detailed evaluation along the lines of [15] for future exploration.

### v2

@@
-Optimal Rates for DP-SCO with a Single Epoch and Large Batches
+Optimal Rates for $O(1)$-Smooth DP-SCO with a Single Epoch and Large Batches
 Abstract
-The most common algorithms for differentially private (DP) machine learning (ML) are all based on stochastic gradient descent, for example, DP-SGD .
-These algorithms achieve DP by treating each gradient as an independent private query.
-However, this independence can cause us to overpay in privacy loss because we don’t analyze the entire gradient trajectory.
-In this work, we propose a new DP algorithm, which we call Accelerated-DP-SRGD (DP stochastic recursive gradient descent), that enables us to break this independence and only pay for privacy in the gradient difference, i.e., in the new information at the current step.
-Our algorithm achieves the optimal DP-stochastic convex optimization (DP-SCO) error (up to polylog factors) using only a single epoch over the dataset, and converges at the Nesterov’s accelerated rate.
-Our algorithm can be run in at most $\sqrt{n}$ batch gradient steps with batch size at least $\sqrt{n}$, unlike prior work which required $O(n)$ queries with mostly constant batch sizes.
+In this paper we revisit the DP stochastic convex optimization (SCO) problem.
+For convex smooth losses, it is well-known that the canonical DP-SGD (stochastic gradient descent) achieves the optimal rate of $O\left(\frac{LR}{\sqrt{n}}+\frac{LR\sqrt{p\log(1/\delta)}}{\varepsilon n}\right)$ under $(\varepsilon,\delta)$-DP [11], and also well-known that variants of DP-SGD can achieve the optimal rate in a single epoch [23].
+However, the batch gradient complexity (i.e., number of adaptive optimization steps), which is important in applications like federated learning, is less well-understood.
+In particular, all prior work on DP-SCO requires $\Omega(n)$ batch gradient steps, multiple epochs, or convexity for privacy.
+We propose an algorithm, Accelerated-DP-SRGD (stochastic recursive gradient descent), which bypasses the limitations of past work: it achieves the optimal rate for DP-SCO (up to polylog factors), in a single epoch using $\sqrt{n}$ batch gradient steps with batch size $\sqrt{n}$, and can be made private for arbitrary (non-convex) losses via clipping.
+If the global minimizer is in the constraint set, we can further improve this to $n^{1/4}$ batch gradient steps with batch size $n^{3/4}$.
 To achieve this, our algorithm combines three key ingredients, a variant of stochastic recursive gradients (SRG), accelerated gradient descent, and correlated noise generation from DP continual counting.
-Finally, we also show that our algorithm improves over existing SoTA on multi-class logistic regression on MNIST and CIFAR-10.
 1 Introduction
@@
 Optimal rates are known for nearly all classes of convex losses, e.g., Lipschitz (and smooth) [7, 23], and Lipschitz with strong convexity [24].
-However, there is comparatively little work in optimizing the number of gradient calls [12] required to achieve these rates [35, 27, 29].
-In this work, we show that it is possible to obtain optimal DP-SCO error (upto polylog $n,1/\delta$ factors) with a single epoch gradient based algorithm for $\ell_{2}$-Lipschitz and smooth losses.
-More precisely, we show the following:
-There exists a gradient based DP algorithm that makes a single pass over the data set (a.k.a.
-single epoch  [^4] Formally, a gradient based algorithm $\mathcal{A}(D)\to\{s_{1},\ldots,s_{T}\}$ is called single epoch, if it computes the gradients on any data point $d\in D$ at exactly one state $s_{t}$.
-Without additional restrictions, a state is allowed to encode arbitrary amount of history, i.e., $s_{t}\supseteq\{s_{1},\ldots,s_{t-1}\}$.) and achieves optimal DP-SCO error (within polylog factors), with $T=n/B$ batch gradient steps and batch size $B=O(\sqrt{n})$ The algorithm uses exactly two gradients per training example [^5] Here, $n$ is the number of training examples..
-Our work significantly improves the results of [27, 55] that required at least $T=\Omega(n)$ batch gradient steps (a.k.a.
-adaptive gradient oracle calls) with batch size $B=n/T$ to reach the optimal DP-SCO error.
-Further, unlike [27] the privacy guarantee of our algorithm does not need to assume convexity, which makes it empirically applicable in non-convex settings too without violating DP guarantees.
-We demonstrate this with an empirical comparison on multi-class logistic regression problems (for which the clipped loss is non-convex [48]), where we show improvements over the strong baseline algorithm of DP-FTRL (DP follow the regularized leader) [38, 23, 15, 14, 13] on standard benchmarks.
-(See Section 2 for more details on DP-SGD and DP-FTRL .)
-Problem definition: There is a distribution of examples $\mathcal{D}$ and a loss function $f(\mathbf{x},d)$ where $\mathbf{x}\in\mathcal{C}$ is a model in the constraint set $\mathcal{C}$ and $d$ is an example.
+There also have been extensive work in the recent past [31, 23, 24, 26, 51] optimizing for both the number of gradient oracle calls [12] and rounds of adaptive interactions [46, 48] while achieving optimal DP-SCO error.
+In this work, we revisit the problem of reducing both the gradient complexity, and the number of adaptive interactions (a.k.a.
+batch gradient steps) for smooth [^4] Unless mentioned otherwise, we refer to $O(L/\left\|\mathcal{C}\right\|_{2})$ $\ell_{2}$-smoothness.
+For brevity, we will sometimes refer to this as $O(1)$-smoothness as this gives smoothness parameter $M=O(1)$ if $L,\left\|\mathcal{C}\right\|_{2}=\Theta(1)$, and $M=O(L/\left\|\mathcal{C}\right\|_{2})$ is within a constant of the lower bound $M\geq L/\left\|\mathcal{C}\right\|_{2}$ that holds if $\mathcal{C}$ contains the global minimizer.
+and convex losses.
+We show that it is possible to obtain optimal DP-SCO error (upto $\text{polylog}(n,1/\delta)$ factors), in a single epoch algorithm with sublinear batch gradient steps.
+Problem definition: There is a distribution of examples $\mathcal{D}$ and a loss $f(\mathbf{x},d)$ where $\mathbf{x}\in\mathcal{C}\subseteq\mathbb{R}^{p}$ is a model in the constraint set $\mathcal{C}$, $\left\|\mathcal{C}\right\|_{2}\leq R$, and $d$ is an example.
 We assume for any $d\in supp(\mathcal{D})$, $f$ is convex, $L$-Lipschitz, and $M$-smooth over $\mathcal{C}$.
+We assume for some $c_{M}=O(1)$, we have $M\leq c_{M}L/\left\|\mathcal{C}\right\|_{2}$.
 We receive a sequence of $n$ i.i.d.
 examples $D=\{d_{1},d_{2},\ldots,d_{n}\}\sim\mathcal{D}$.
-Using these examples, our goal is to choose a model that minimizes the population loss function $F(\mathbf{x})=\mathbb{E}_{d\sim\mathcal{D}}\left[f(\mathbf{x},d)\right]$.
+Using these examples, our goal is to choose a model in $\mathcal{C}$ that minimizes the population risk function $F(\mathbf{x})=\mathbb{E}_{d\sim\mathcal{D}}\left[f(\mathbf{x},d)\right]$.
 We will operate in the single-pass streaming setting: we must process examples in the order they appear, and each example appears only once in this order.
-Finally, the model we output should satisfy $(\varepsilon,\delta)$-DP with respect to the examples.
+Finally, the model we output should satisfy $(\varepsilon,\delta)$-DP w.r.t.
+the examples.
 For simplicity, we use the zero-out notion of DP [42]: if $P$ is the distribution of models our algorithm outputs given $\{d_{1},d_{2},\ldots,d_{n}\}$ and $Q$ is the distribution given the same sequence, except one example $d_{i}$ instead has $\nabla f(\mathbf{x},d_{i})=0$ for all $\mathbf{x}$, then we require $\max_{S}[P(S)-e^{\varepsilon}Q(S)],\max_{S}[Q(S)-e^{\varepsilon}P(S)]\leq\delta$.
-Adaptive oracle complexity (a.k.a.
-batch gradient steps): We enforce our algorithm to be single epoch (2.1).
-Furthermore, we quantify the computational complexity of the algorithms in terms of two quantities: a) Number of adaptive gradient computations ($T$) the algorithm has to perform, i.e., the current computation depends on the output of the gradient computations in the previous state [50, 53], and b) number of gradient oracle calls made to each data point overall.
-Our algorithm has $T=O(\sqrt{n})$, and makes exactly two gradient calls per data point.
-In comparison, [27, 55] has $T=O(n)$[27] makes one gradient call/data point (but relies on convexity for privacy guarantee), [55] makes two gradient calls/data point (same as ours).
-In practice, adaptive calls are much more expensive [6, 37], hence we focus on reducing adaptive computation cost as opposed to focusing on the gap between one vs two gradient oracle calls per data point.
-Designing private learning algorithms that minimize adaptive computations is a well-established problem in the literature [50, 53, 21].
-Even in practice, adaptive gradient evaluations are more expensive than gradient calls on multiple data points at the same state, especially in distributed learning settings like federated learning [6, 37].
-To the best of our knowledge, this is the first paper to explicitly optimize for the adaptive oracle complexity for the DP-SCO problem.
-[53] show that in the non-private setting, the adaptive oracle complexity for optimal SCO guarantee is $T=\Omega(n^{1/4})$ (with a batch size of $B=O(n^{3/4})$).
-This is achieved by Nesterov’s accelerated SGD.
-It is worth pointing out that Accelerated-DP-SRGD does not match this non-private bound even for constant dimensions and $\varepsilon\to\infty$.
-The main reason being that we use stochastic recursive gradients (SRGs) [45, 55] (instead of minibatch gradients) for privacy purposes, and the variance of the SRGs overwhelm the variance due to DP noise at large $\varepsilon$’s.
-We leave it as an open problem to match the rates in [53] under DP.
-Design of Accelerated-DP-SRGD (DP stochastic recursive gradient descent): In terms of design,
-Accelerated-DP-SRGD is based on a careful adaptation of Nesterov’s accelerated stochastic gradient descent [44, 9], combined with the ideas of stochastic recursive gradients (SRG) [45, 55], and matrix factorization (MF) based private continual counting mechanisms [24, 19, 23, 42, 26, 33, 34, 5].
+Our result: We show the following:
+  1. 1.
+    Unconstrained optimization: Assuming the unconstrained population risk minimizer is in the interior of the constraint set (i.e., $\mathbf{x}^{\dagger}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{p}}F(\mathbf{x})$, and $\mathbf{x}^{\dagger}\in\mathcal{C}$) [^5] In the unconstrained optimization setting, a standard assumption is a known bound $R$ on the norm of $\mathbf{x}^{\dagger}$ and we can set $\mathcal{C}=B(\boldsymbol{0},R)$ to satisfy our assumption.
+Hence, our assumption can be viewed as implying an unconstrained optimization of the population risk $F(\mathbf{x})$.
+A similar assumption was made in [49, Theorem 8] while designing accelerated version of dual averaging methods., there exists a gradient based DP algorithm that makes a single pass over the data set (a.k.a.
+single epoch) and achieves optimal DP-SCO error (within polylog factors) for smooth and convex losses, with $T=\widetilde{\Theta}(n^{1/4})$ batch gradient steps [^6] Here, $n$ is the number of training examples., batch size $B=n/T$, and using exactly two gradients per training example.
+  2. 2.
+    Constrained optimization: We also show that even without assuming $\mathbf{x}^{\dagger}\in\mathcal{C}$our algorithm obtains the optimal DP-SCO bounds with $T=\sqrt{n}$ batch gradient steps, batch size $B=\sqrt{n}$ and in a single epoch.
+Our work uniformly improves over all prior [23, 50, 51] and contemporary [26] works under the assumption that $\mathbf{x}^{\dagger}\in\mathcal{C}$.
+Under smoothness assumption, prior SoTA (in terms of batch gradient complexity) required gradient complexity of $\tilde{\Theta}(n)$, and with batch gradient complexity of $T=\tilde{\Theta}(\sqrt{n})$.
+Under a single-epoch constraint, to the best of our knowledge no prior work achieves $T=o(n)$.
+In comparison, our work improves the gradient complexity to $2n$, and batch gradient steps to $T=\widetilde{\Theta}(n^{1/4}$).
+A contemporary work [26] matches our batch gradient steps ($T=\tilde{\Theta}(n^{1/4})$).
+However, they have worse gradient complexity of $\tilde{\Theta}(n^{9/8})$.
+Our guarantees, even without assuming $\mathbf{x}^{\dagger}\in\mathcal{C}$, improves over all prior work (at least by $\text{polylog}(n)$), and is incomparable to the contemporary work of [26].
+(See Table 1.)
+Works of [23, 14, 26] also study DP-SCO with higher smoothness conditions of e.g.
+$M=O(\sqrt{n})$, which by Nesterov smoothing [40] corresponds to the non-smooth setting.
+However, these works require $n^{\Omega(1)}$ passes over the dataset; to the best of our knowledge there is no result in the literature giving optimal rates for DP-SCO in $n^{o(1)}$ epochs in the non-smooth setting (for all settings of the dimension $p$; [14] uses $O(1)$ epochs if $p\leq\varepsilon^{3/2}$).
+Since we are interested in the single-epoch setting, we hence only focus on $O(1)$-smoothness, and higher smoothness is beyond scope.
+As an aside, in the $M=O(\sqrt{n})$ regime, [14] achieves worst-case gradient complexity of $\widetilde{\Theta}(n^{4/3})$, as opposed to $\widetilde{\Theta}(n^{11/8})$ by [26].
+Gradient complexity and adaptive interaction complexity (a.k.a.
+batch gradient steps): We measure the computational complexity of our algorithms under standard complexity measures in the convex optimization literature: a) Gradient complexity [12]: It measures the total number of gradient evaluations of the loss function $f(\mathbf{x},\cdot)$, and b) Batch gradient steps (a.k.a.
+adaptive interaction complexity)  [^7] Henceforth, we will refer to it as batch gradient steps for consistency.
+[46, 48]: It corresponds to the number of number of adaptive gradient computations ($T$) the algorithm has to perform, i.e., the current computation depends on the output of the gradient computations in the previous state.
+Understanding the complexity of an algorithm in terms of number of batch gradient steps have significant attention in the recent past [46, 48, 17, 14].
+It is especially important in the context of distributed/federated optimization, where adaptive interactions are much more expensive compared to parallel gradient evaluations at a fixed model state.
+For example, in cross-device federated learning [33], the number of rounds adaptive server/device communication dominate the overall optimization complexity.
+Furthermore, our algorithm has additional desirable properties, which are crucial in any practical deployment:
+  1. (a)
+    Single pass/single epoch: A gradient based algorithm $\mathcal{A}(D)\to\{s_{1},\ldots,s_{T}\}$ is called single epoch, if it computes the gradients on any data point $d\in D$ at exactly one state $s_{t}$.
+Without additional restrictions, a state is allowed to encode arbitrary amount of history, i.e., $s_{t}\supseteq\{s_{1},\ldots,s_{t-1}\}$.
+Single epoch  [^8] Notice that the definition does not include any restriction on the memory of the algorithm.
+While stating our results in the single epoch setting, we additionally incorporate the batch size, and the total number gradients evaluated per data point.
+algorithms are crucial in learning over data streams (where you can process any data point at any single model state), or in cross-device federated learning (FL) where a device holding a specific data point appears only once during the training [10].
+    It is worth emphasizing that there is a crucial difference between an algorithm using a single epoch and an algorithm that performs even two epochs over the data; one may be implementation wise infeasible because of the reasons stated above.
+So, in Table 1 even if the gradient complexity matches asymptotically between two algorithms, still one may be infeasible to implement due to the fact of not being single epoch.
+  2. (b)
+    Assuming convexity for privacy: In DP-SCO literature (see Table 1), prior works have assumed that the loss function $f(\mathbf{x},d)$ is convex to prove both the privacy and utility guarantee of their algorithms.
+Assuming convexity for differential privacy guarantees extremely restricts the applicability of the algorithms in practical scenarios where provable convexity property may not hold (e.g., training non-convex models with DP-SGD [1], or simply clipping the gradients to a fixed $\ell_{2}$-norm [45].)
+Surprisingly, our algorithms, even being single epoch and not assuming convexity for privacy, uniformly beat the prior SoTA in terms of gradient complexity and batch gradient steps.
+Algorithms Gradient Batch gradient Single Convexity
+complexity steps epoch for privacy
+Ours (assuming $\mathbf{x}^{\dagger}\in\mathcal{C}$) $2n$ $\widetilde{\Theta}(n^{1/4})$ Yes No
+Ours (no assumption) $2n$ $\sqrt{n}$ Yes No
+[23] $n$ $\widetilde{\Theta}(n)$ Yes Yes
+[50] $2n$ $n$ Yes No
+[51] $\widetilde{\Theta}(n)$ $\widetilde{\Theta}(\sqrt{n})$ No Yes
+[26] $\widetilde{\Theta}(n^{9/8})$ $\widetilde{\Theta}(n^{1/4})$ No No
+Table 1: Comparison of gradient complexity and batch gradient steps.
+Here, $\mathbf{x}^{\dagger}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{p}}F(\mathbf{x})$ is the unconstrained population risk minimizer, and the Lipschitz constant ($L$) and the smoothness required for all results is $O(1)$.
+$\widetilde{\Theta}(\cdot)$ hides $\text{polylog}(n)$ factors.
+Our algorithm [Accelerated-DP-SRGD (DP stochastic recursive gradient descent)]: In terms of design, our algorithm is based on a careful adaptation of Nesterov’s accelerated stochastic gradient descent [38, 8], combined with the ideas of stochastic recursive gradients (SRG) [41, 50], and private continual counting mechanisms [19, 16, 18, 37, 22, 29, 30, 5].
+In the hindsight, Accelerated-DP-SRGD is arguably more natural than prior/contemporary adaptations of accelerated methods [26, 51] towards reducing the computational complexity of obtaining optimal DP-SCO guarantees.
 Recall, Nesterov’s acceleration can be expressed as the coupling of two gradient descent based algorithms with different learning rates.
@@
 We use the idea of stochastic recursive gradients (SRGs) from [41, 50] to first approximate $\nabla_{t}$ as $\eta_{t}\nabla_{t}\approx\Delta_{t}+\Delta_{t-1}+\ldots$, where each $\Delta_{t}=\eta_{t}\nabla_{t}(\mathbf{x}_{t})-\eta_{t-1}\nabla_{t}(\mathbf{x}_{t-1})$.
-We set $\eta_{t}=t$, $\beta\approx M\cdot\max\left\{\frac{n^{3/2}}{B^{2}},\frac{n}{B}\right\}$, and $\tau_{t}=\frac{\eta_{t}}{\sum\limits_{\tau\leq t}\eta_{\tau}}$.
-Following are the few crucial differences from [55]: a) We use a truly accelerated method to achieve the improved batch gradient steps, as opposed to tail-averaged online convex optimization methods (which does not include the term $\mathbf{y}_{t}$), b) because of acceleration, we cannot tolerate higher values of $\eta_{t}$.
+We set $\eta_{t}=t$, $\beta\approx M$ if we assume $\mathbf{x}^{\dagger}\in\mathcal{C}$ and $\beta\approx M\cdot\max\left\{\frac{n^{3/2}}{B^{2}},\frac{n}{B}\right\}$ otherwise, and $\tau_{t}=\frac{\eta_{t}}{\sum\limits_{\tau\leq t}\eta_{\tau}}$.
+Following are the few crucial differences from [50]: a) We use a truly accelerated method to achieve the improved batch gradient steps, as opposed to tail-averaged online convex optimization methods (which does not include the term $\mathbf{y}_{t}$), and b) because of acceleration, we cannot tolerate higher values of $\eta_{t}$.
 In particular the results of [50] hold with $\eta_{t}=t^{k}$ (for any $k\geq 1$).
-To bound the sensitivity of $\Delta_{t}$ for privacy, we need to slow down the learning by choosing $\beta\approx M\sqrt{n}$ (corresponding to the iterations $T=n/B\approx\sqrt{n}$), as opposed $\beta=M$ in the non-private case.
-With this choice of parameters, we first show that $\Delta_{t}$ has sensitivity $O(1/B)$.
-Hence, we can use the standard MF based continual counting mechanisms (Section 2.2) to compute DP variants of each $\eta_{t}\nabla_{t}$ with low error.
-We finally get an SCO error of the form: $\widetilde{O}\left(\frac{B^{2}}{n^{2}}+\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\right)$, with $B$ being the minibatch size.
-The first two terms match the non-private optimal SCO error from [41].
-Comparison to most related prior work [27, 55]: Two of the prior works which are most related to our work are of [27, 55].
-Here, we position the ideas in this paper in regard to these works.
-For $d,\varepsilon=\Theta(1)$ both these algorithms use $T=O(n)$ batch gradient steps, with batch size $\approx O(1)$.
-Our main algorithm Accelerated-DP-SRGD , uses Nesterov’s accelerated method [44, 9] as the baseline algorithm to improve on the SoTA to $T=B=O(\sqrt{n})$.
-One natural question is whether the algorithms in the prior work with small modifications can achieve our adaptive oracle complexity.
-[27] either relies on privacy amplification by iteration [28] or localization based methods (which requires adding adaptive $\ell_{2}$-regularization).
-In either case choosing a batch size larger than $O(\max\{\sqrt{d/n},1\})$ in majority of the batch gradient steps, will result in a sub-optimal DP-SCO error.
-In comparison, for $M=\Theta(1)$ our algorithm does not have any dependence of $d$ on the batch gradient complexity.
-Furthermore, [27] heavily uses convexity for their privacy analysis.
-In contrast, Accelerated-DP-SRGD
-can be used seamlessly in non-convex settings without compromising on the DP guarantees (via clipping of $\Delta_{t}$).
-The algorithm of [55] and their analysis relies on the access to a low-regret online convex optimization algorithm (OCO) [31].
+To bound the sensitivity of $\Delta_{t}$ for privacy, we need to either (i) assume $\mathbf{x}^{\dagger}\in\mathcal{C}$, which allows us to show the gradient steps are decreasing over time and hence the gradient difference between $\mathbf{x}_{t}$ and $\mathbf{x}_{t-1}$ is small by smoothness, or (ii) slow down the learning by choosing $\beta\approx M\sqrt{n}$ (corresponding to the iterations $T=n/B\approx\sqrt{n}$), as opposed to $\beta=M$ in the non-private case.
+Our analysis: While we argue above that our algorithm is a natural extension of Nesterov’s accelerated SGD, the privacy and the utility analysis of the algorithm is novel, and far from immediate.
+We give a brief sketch of our analysis below.
+At a high-level our analysis requires bounding the $\ell_{2}$-sensitivity of the gradient difference $\Delta_{t}$ to $O(\frac{1}{B\cdot t})$, and then use the binary tree mechanism [19] to compute DP variants of each $\eta_{t}\nabla_{t}$ with low error.
+We finally get an SCO error of the form: $\frac{LR}{\sqrt{n}}+\frac{LR\sqrt{p}}{\varepsilon n}+\frac{LRB}{n}+\frac{LRB\sqrt{p}}{\varepsilon n^{2}}$, with $B$ being the minibatch size.
+The first two terms match the optimal DP-SCO error.
+Setting $B\leq\sqrt{n}$ gives that the third and fourth term also are at most the optimal DP-SCO error.
+The main challenge is bounding the sensitivity of term $\Delta_{t}$ (stated above).
+Bounding it naively to $O\left(\frac{tM}{B\beta}+L\right)$ (using the analysis of [50]), may result in a sub-optimal choice of the batch size for the final DP-SCO guarantee.
+Instead we couple the privacy analysis and the utility analysis together.
+Recall, $\mathbf{x}^{\dagger}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{p}}F(\mathbf{x})$ is the unconstrained minimizer of the population risk $F(\mathbf{x})$, and we assume $\mathbf{x}^{\dagger}\in\mathcal{C}$.
+We observe that $\left\|\nabla F(\mathbf{x})\right\|_{2}^{2}\leq 2M(F(\mathbf{x})-F(\mathbf{x}^{\dagger}))$ for any $\mathbf{x}\in\mathcal{C}$.
+This allows us to get a tighter control over the norm of the difference in mini-batch gradients: $\frac{1}{B}\cdot\left\|\sum\limits_{d\in\mathcal{B}}\nabla f(\mathbf{x}_{t};d)-\sum\limits_{d\in\mathcal{B}}\nabla f(\mathbf{x}_{t-1};d)\right\|_{2}$, rather than bounding it bluntly by $\frac{M}{B\beta}$.
+As the algorithm proceeds and the excess population risk $F(\mathbf{x})-F(\mathbf{x}^{*})$ goes down, $\left\|\frac{1}{B}\sum\limits_{d\in\mathcal{B}}\nabla f(\mathbf{x}_{t};d)\right\|_{2}$ becomes smaller.
+This in-turn can be used to obtain a stronger bound of $O\left(\frac{M}{B\cdot\beta t}\right)$, on the norm of the difference in mini-batch gradients.
+We use this to reduce the sensitivity of $\Delta_{t}$ to $O\left(\frac{M}{B\beta}+L\right)$.
+This is formalized via Lemma 5.5.
+If we do not assume that $\mathbf{x}^{\dagger}\in\mathcal{C}$, then we proceed with the naive bound on the gradient difference, and compensate by increasing $\beta$, which in turn requires more iterations for the non-private algorithm to converge and hence a smaller batch size of $\Omega(\sqrt{n})$ (as opposed to $B=\widetilde{\Omega}(n^{3/4})$).
+This still improves on prior work at least by logarithmic factors.
+More importantly, we still keep the algorithm single epoch and do not require convexity, which is not possible with the comparable prior work of [51].
+Comparison to relevant prior work: While DP-SCO has been studied extensively in the literature, we first focus on the works that are most relevant for this paper [23, 50, 26, 51].
+As summarized in Table 1, our bounds are uniformly better than prior and contemporary works under the assumption on the unconstrained population risk minimizer being in the interior of $\mathcal{C}$ (i.e., $\mathbf{x}^{\dagger}\in\mathcal{C}$).
+Without this assumption, our algorithm is still uniformly better than prior works, and incomparable to the contemporary work of [26].
+Unlike many of the prior works, our algorithm is single epoch, and does not require convexity assumption to ensure differential privacy.
+In terms of techniqiue, our algorithm heavily relies on the idea of coupling DP-continual observation [16, 19, 13, 29] with stochastcic recursive gradients [41] from [50].
+However, there are a few crucial differences which allows us to handle much larger batch sizes while still being a single epoch algorithm.
+[50] and their analysis relies on the access to a low-regret online convex optimization algorithm (OCO) [28].
 We know that the regret guarantees of OCO methods cannot be improved beyond $\Omega(1/\sqrt{T})$ even under smoothness [28].
@@
 In Section 4.3 we show that unfortunately, because of the variance introduced by the SRGs, in the absence of an accelerated SGD method, the optimal batch size is still $O(1)$.
+The fact that we use Nesterov’s acceleration allows us to handle much larger batch sizes.
 Other related work: We discuss some of the prior work [14, 9, 2, 25].
@@
 [14] provides the best known gradient oracle complexity for optimal DP-SCO, for non-smooth convex losses under $\ell_{2}$-geometry.
-While their algorithm is not single-epoch, their overall gradient complexity (i.e., the number of gradient oracle calls) is $\widetilde{O}\left(\min\left\{n,\frac{n^{2}}{d}\right\}+\min\left\{\frac{nd^{2}}{\varepsilon},n^{4/3}\varepsilon^{1/3}\right\}\right)$.
+While their algorithm is not single-epoch, their overall gradient complexity (i.e., the number of gradient oracle calls) is $\widetilde{O}\left(\min\left\{n,\frac{n^{2}}{p}\right\}+\min\left\{\frac{np^{2}}{\varepsilon},n^{4/3}\varepsilon^{1/3}\right\}\right)$.
 This result is incomparable to our result, since we rely on smoothness in our analysis, but our overall gradient oracle complexity is an improved $\widetilde{O}\left(\sqrt{n}\right)$.
-Under constant $\varepsilon$, under all regimes of the dimensionality where the DP-SCO guarantee is meaningful i.e., $d\leq n^{2}$, our gradient oracle complexity is superior.
+Under constant $\varepsilon$, under all regimes of the dimensionality where the DP-SCO guarantee is meaningful i.e., $p\leq n^{2}$, our gradient oracle complexity is superior.
 We leave the possibility of removing the smoothness assumption in our paper while obtaining the same (or better) gradient oracle complexity as an open problem.
@@
 For example, DP-SCO guarantees under the loss functions being $\ell_{1}$-Lipschitz, or SCO properties of second-order-stationary points.
-These results, while being important for the DP-SCO literature, are not comparable to our line of work.
-Practical enhancements, and empirical evaluation: Along with demonstrating the theoretical benefits of Accelerated-DP-SRGD , we demonstrate that the core ideas in our algorithm can be used to improve well-optimized benchmark results for DP learning [38, 23, 15].
-While we chose $\eta_{t}=t$ in Accelerated-DP-SRGD to get optimal DP-SCO bounds, in practice we found setting $\eta_{t}=e^{\lambda t}$ and optimizing for $\lambda$ provides the best results.
-With this reparameterization, we make the crucial observation that if $\lambda$ is very high and $\tau_{t}=1$, then the Nesterov’s update simplifies to standard SGD update.
-Hence, there exists a parameterization under which Accelerated-DP-SRGD is equivalent to DP-FTRL  (differentially private follow-the-regularized-leader) [38, 23].
-Hence, with more optimized parameters we can only get a better algorithm.
-The exact algorithm we use for empirical evaluation is given in Algorithm 3 (Algorithm DP-SRG-MEMF ).
-In our empirical evaluation, we provide results for multi-class logistic regression on MNIST and CIFAR-10 as a proof-of-concept.
-We show that Accelerated-DP-SRGD provides a utility benefit over the prior SoTA [15] for the no-amplification [^7] Prior works [38, 43] mentioned that in distributed learning environments like federated learning, privacy amplification by sampling [36] is undesirable, and even when it is applicable, the true privacy guarantee depends on the precise implementation [16].
-Hence, we focus on the evaluation of our algorithms when it does not have access to any form of privacy amplification.
-(i.e., streaming) setting that has been optimized for over a series of work [38, 23, 15, 14, 13].
-A detailed empirical exploration of Accelerated-DP-SRGD and its practical variant on large-scale practical tasks is left for future exploration.
-1.1 Our Contributions
+These results, while being important for the DP-SCO literature, are not comparable to our work.
+Future directions: While our work significantly improves on the SoTA in terms of gradient complexity and batch gradient steps in DP-SCO for smooth convex losses, it raises a lot of interesting open questions:
+  1. (a)
+    Can we obtain the same tighter batch gradient steps guarantee in Table 1, without assuming the unconstrained minimizer to be in the set $\mathcal{C}$?
+Informally, we only use this assumption to show the gradient norm is decreasing, i.e.
+the distance between $\mathbf{x}_{t}-\mathbf{x}_{t-1}$ is decreasing in $t$, and so the gradient differences $\Delta_{t}$ are bounded by smoothness.
+However, if gradients remain large during training, we expect training to reach the boundary of $\mathcal{C}$ quickly, in which case $\mathbf{x}_{t}-\mathbf{x}_{t-1}$ should also be small (because the gradient step is mostly reversed by the projection into $\mathcal{C}$).
+Hence we believe this assumption is not required for Accelerated-DP-SRGD to achieve $T=n^{1/4}$, but we do not know how to formally show this.
+  2. (b)
+    In our paper we only focused on $O(1)$-smooth convex losses.
+To the best of our knowledge, there is no single-epoch algorithm for the non-smooth setting.
+The best result we are aware of in this setting is that of [14], which uses $O(n)$ gradient oracle calls when e.g.
+$p,\varepsilon$ are constant, but can require up to $O(n^{4/3})$ gradient oracle calls for higher dimensions.
+It remains an open question what the smallest dimension-independent gradient complexity and batch gradient complexity required for optimal DP-SCO bounds in the non-smooth setting is.
+Our approach seems to require smoothness as the main primitive, stochastic recursive gradients, cares about differences of gradients at different points.
+So it is unlikely our techniques can be used to achieve similar improvements in the non-smooth setting.
+  3. (c)
+    Finally, while we believe the batch gradient steps bound of $n^{1/4}$ is tight for DP-SCO, we do not know how to prove a lower bound.
+Standard lower bounding techniques from the non-private literature [48] for batch gradient steps break down under DP setting as they operate in the regime where the dimensionality $p=\Omega(n^{2})$.
+In such high-dimensional regime, DP-SCO guarantees are vacuous.
+A well known result of [39] states that that no first-order method (i.e., algorithm that is only allowed to make gradient queries to points in the span of all gradients received so far) can achieve an empirical loss bound of $o(1/T^{2})$ in $T$ gradient queries for convex smooth losses, which even ignoring privacy suggests $T=n^{1/4}$ iterations are needed for any algorithm’s empirical loss to match the optimal stochastic error $1/\sqrt{n}$.
+It would be surprising if a DP algorithm could achieve optimal rates in $o(n^{1/4})$ iterations, as in the limit when $\varepsilon\rightarrow\infty$ it would beat this non-private lower bound.
+However, it is not obvious how to extend this lower bound to DP algorithms, as the noise addition violates the constraint that only points in the span of gradients are queried.
+1.1 Our Contribution
   1. 1.
@@
   2. 2.
-    Obtaining optimal DP-SCO errors with improved number of batch gradient steps (Theorem 4.6): We show that Accelerated-DP-SRGD can achieve optimal DP-SCO errors for smooth convex losses, in a single epoch with batch gradient steps of $T\approx\sqrt{n}$ (batch size $B=n/T)$, and exactly two gradient calls per data point.
-We required two technical ideas to achieve this: a) reducing the learning rate of non-private Nesterov’s method to account for the noise added due to privacy, and b) show that there exists appropriate choice of parameters for the algorithm under which the SRGs are low-sensitive (i.e., does not change too much by changing one data sample in the data set $D$).
-  3. 3.
-    Surpassing SoTA on benchmark tasks (Section 5): Inspired by our Accelerated-DP-SRGD , we create a heuristic (via exponential re-weighting of gradient differences) that empirically outperforms the highly optimized class of DP-FTRL algorithms [38, 23, 15] on multi-class logistic regression on MNIST and CIFAR-10.
+    Obtaining optimal DP-SCO errors with improved number of batch gradient steps (Theorem 4.2): We show that Accelerated-DP-SRGD can achieve optimal DP-SCO errors for $O(1)$-smooth convex losses, in a single epoch with batch gradient steps of $T\approx n^{1/4}$ (batch size $B=n/T$), if the unconstrained population risk minimizer $\mathbf{x}^{\dagger}$ is within the constraint set $\mathcal{C}$.
+Furthermore, with $T\approx\sqrt{n}$ (batch size $B=n/T$) when $\mathbf{x}^{\dagger}\not\in\mathcal{C}$.
+Our algorithm exactly two gradient calls per data point, and is single epoch, and does not make any convexity assumption for privacy guarantees.
 2 Background on (Private) Learning
-2.1 Definition for Single Epoch Training
-Definition 2.1 (Single epoch).
-A gradient based algorithm $\mathcal{A}(D)\to\{s_{1},\ldots,s_{T}\}$ is called single epoch, if it computes the gradients on any data point $d\in D$ at exactly one state $s_{t}$.
-Without additional restrictions, a state is allowed to encode arbitrary amount of history, i.e., $s_{t}\supseteq\{s_{1},\ldots,s_{t-1}\}$.
-Notice that 2.1 does not include any restriction on the memory of the algorithm.
-While stating our results in the single epoch setting, we additionally incorporate the batch size, and the total number gradients evaluated per data point.
-2.2 DP-Matrix Factorization (DP-MF ) and Continual Counting
-Continual counting under DP [19, 24] is one of the most important primitives in the DP literature.
-Given a sequence of adaptively chosen data vectors $D=\{\mathbf{v}_{1},\ldots,\mathbf{v}_{T}\}$ (with each $\left\|\mathbf{v}_{i}\right\|_{2}\leq 1$), the objective is to output $\mathbf{z}=\left[\sum\limits_{\tau\leq t}\mathbf{v}_{\tau}\right]_{t\in[T]}$ while preserving DP.
-It is well-known that one can output $\mathbf{z}_{\texttt{priv}}$ s.t.
-$\mathbb{E}[\|\mathbf{z}-\mathbf{z}_{\texttt{priv}}\|_{\infty}]=\text{polylog}(T)$.
-This can be achieved via the classic binary tree mechanism [19, 24, 32], see Lemma 4.3.
-We use the binary tree primitive for our theoretical guarantees.
-A recent line of work [23, 42, 26, 33, 34, 5] optimizes the error up to additive constant via factorization of a lower-triangular matrix of all ones $\mathbf{A}=\mathbf{B}\cdot\mathbf{C}$.
-For our empirical results, we appeal to the computational approach of [23] for finding the best factorizations ($\mathbf{B}$ and $\mathbf{C}$) w.r.t.
-the error in continual counting under DP.
 2.1 DP-SGD : DP stochastic gradient descent
@@
 2.2 DP-FTRL : DP follow the regularized leader
-DP-FTRL  [49, 38, 23] is another standard approach for optimization, where in its simplest form, the state update is exactly that of DP-SGD when $\mathcal{C}=\mathbb{R}^{d}$, except the noise in two different rounds $\mathbf{b}_{t},\mathbf{b}_{t^{\prime}}$ are not independent.
-Instead (in the one-dimensional setting) $(\mathbf{b}_{1},\mathbf{b}_{2},\ldots\mathbf{b}_{T})\sim\mathcal{N}(0,\left(\mathbf{C}^{\top}\mathbf{C}\right)^{-1})$, where $\mathbf{C}$ is computed via DP-MF .
+DP-FTRL  [44, 34, 18] is another standard approach for optimization, where in its simplest form, the state update is exactly that of DP-SGD when $\mathcal{C}=\mathbb{R}^{p}$, except the noise in two different rounds $\mathbf{b}_{t},\mathbf{b}_{t^{\prime}}$ are not independent.
+Instead (in the one-dimensional setting) $(\mathbf{b}_{1},\mathbf{b}_{2},\ldots\mathbf{b}_{T})\sim\mathcal{N}(0,\left(\mathbf{C}^{\top}\mathbf{C}\right)^{-1})$, where $\mathbf{C}$ is chosen to minimize the error on prefix sums of $\mathbf{x}_{t}$.
+A canonical and asymptotically optimal implementation of DP-FTRL  is the binary tree mechanism of [19, 16].
 While DP-FTRL has strong empirical properties comparable to DP-SGD , it is unknown whether it can achieve optimal DP-SCO error unless under additional restrictions (e.g., realizability [3]).
+A line of work (e.g.
+[13, 29, 30]) has looked at optimizing the choice of $\mathbf{C}$.
+These works offer constant-factor improvements over the binary tree mechanism.
+While these improvements are meaningful practically, we aim for asymptotic guarantees in this paper, and so we will restrict our choice of correlated noise to the binary tree mechanism for simplicity.
 3 Our Algorithm
@@
 We first give a utility analysis for a fixed value of the noise $\mathbf{b}_{t}$, and then in Section 5.2 we will specify how to set the noise to simultaneously achieve good privacy and utility.
-As mentioned in Section 1, we will use the DP-MF  machinery (Section 2.2) to set the $\mathbf{b}_{t}$’s.
-It is worth mentioning that all three aspects of our algorithm are crucial: a) SRGs allow ensuring that the $\ell_{2}$-sensitivity of the gradient estimate at time $t$ is a factor of $1/t$, lower than that of the direct minibatch gradient ($\frac{1}{B}\sum\limits_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x}_{t};d)$), b) Noise addition via DP-MF  allows the standard deviation of the noise $\mathbf{b}_{t}$ to grow only polylogarithmically in $T$, as opposed to $\sqrt{T}$ via simple composition analysis [25], and c) Accelerated SGD allows appropriately balancing the variance introduced by the SRG estimates, and that due to the i.i.d.
+As mentioned in Section 1, we will use the binary tree mechanism to set the $\mathbf{b}_{t}$’s.
+It is worth mentioning that all three aspects of our algorithm are crucial: a) SRGs allow ensuring that the $\ell_{2}$-sensitivity of the gradient estimate at time $t$ is a factor of $1/t$, lower than that of the direct minibatch gradient ($\frac{1}{B}\sum\limits_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x}_{t};d)$), b) Noise addition via binary tree mechanism allows the standard deviation of the noise $\mathbf{b}_{t}$ to grow only polylogarithmically in $T$, as opposed to $\sqrt{T}$ via simple composition analysis [20], and c) Accelerated SGD allows appropriately balancing the variance introduced by the SRG estimates, and that due to the i.i.d.
 sampling of the minibatches from the population.
 In fact in Section 4.3 we argue that it is necessary to have an accelerated SGD algorithm to achieve a single epoch optimal DP-SCO algorithm with batch size $B=\omega(1)$, if we were to use the SRGs as the gradient approximators.
+1:Input: Steps: $T$, batches: $\{\mathcal{B}_{t}\}_{0\leq t<T}$ of size $B$, learning rates: $\{\eta_{t}\}_{0\leq t<T}$, scaling $\beta$, interpolation params: $\{\tau_{t}\}_{t\in[T]}$, $\mathcal{C}:\left\|\mathcal{C}\right\|_{2}\leq R$.
+2:$\mathbf{x}_{0},\mathbf{z}_{0}\leftarrow 0$.
+3:for $t\in\{0,\ldots,T-1\}$ do
+4:     $\Delta_{t}\leftarrow\frac{1}{B}\sum_{d\in\mathcal{B}_{t}}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)]$
+5:     $\nabla_{t}\leftarrow\frac{\sum_{i\leq t}\Delta_{i}}{\eta_{t}}$ (equivalently, $\nabla_{t}\leftarrow\frac{\Delta_{t}}{\eta_{t}}+\frac{\eta_{t-1}}{\eta_{t}}\cdot\nabla_{t-1}$)
+6:
+7:     $\mathbf{z}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{z}_{t}-\frac{\eta_{t}}{\beta}\nabla_{t}+\mathbf{b}_{t}\right)$, where $\mathbf{b}_{t}$ is the noise added to ensure privacy.
+8:     $\mathbf{y}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}+\frac{1}{\eta_{t}}\cdot\mathbf{b}_{t}\right)$.
+9:     $\mathbf{x}_{t+1}\leftarrow\left(1-\tau_{t+1}\right)\mathbf{y}_{t+1}+\tau_{t+1}\mathbf{z}_{t+1}$.
+10:end for
 Algorithm 1 Accelerated-DP-SRGD
-Inputs: Steps: $T$, batches: $\{\mathcal{B}_{t}\}_{0\leq t<T}$ of size $B$, learning rates: $\{\eta_{t}\}_{0\leq t<T}$, scaling $\beta$, interpolation params: $\{\tau_{t}\}_{t\in[T]}$, constraint set: $\mathcal{C}$ with $0\in\mathcal{C}$.
-1:$\mathbf{x}_{0},\mathbf{z}_{0}\leftarrow 0$.
-2:for $t\in\{0,\ldots,T-1\}$ do
-3:     $\Delta_{t}\leftarrow\frac{1}{B}\sum_{d\in\mathcal{B}_{t}}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)]$
-4:     $\nabla_{t}\leftarrow\frac{\sum_{i\leq t}\Delta_{i}}{\eta_{t}}$ (equivalently, $\nabla_{t}\leftarrow\frac{\Delta_{t}}{\eta_{t}}+\frac{\eta_{t-1}}{\eta_{t}}\cdot\nabla_{t-1}$)
-5:     $\mathbf{z}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{z}_{t}-\frac{\eta_{t}}{\beta}\nabla_{t}+\mathbf{b}_{t}\right)$, where $\mathbf{b}_{t}$ is the noise added to ensure privacy.
-6:     $\mathbf{y}_{t+1}\leftarrow\Pi_{\mathcal{C}}\left(\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}+\frac{1}{\eta_{t}}\cdot\mathbf{b}_{t}\right)$.
-7:     $\mathbf{x}_{t+1}\leftarrow\left(1-\tau_{t+1}\right)\mathbf{y}_{t+1}+\tau_{t+1}\mathbf{z}_{t+1}$.
-8:end for
-4 Utility and Privacy Analysis
-In this section, we provide the detailed utility analysis and privacy analysis of Algorithm 1.
+4 Overview of Analysis of Algorithm 1
+In this section we provide a high-level overview of the analysis of Algorithm 1.
 4.1 Utility Analysis
-Here, we provide the excess population risk guarantee for any choice of the privacy noise random variables $\{\mathbf{b}_{t}\}$’s.
+First, ignoring privacy concerns we provide a excess population risk guarantee of Algorithm 1 for any choice of the privacy noise random variables $\{\mathbf{b}_{t}\}$’s.
 We only assume that the variance (over the sampling of the data samples $d\sim\mathcal{D}$) in any individual term in the SRGs (i.e., $\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$) is bounded by $S^{2}$.
-Eventually in Theorem 4.6 we will combine 4.1 below with the privacy analysis of Algorithm 1 to arrive at the final DP-SCO guarantee, with the appropriate choice of the batch gradient steps, and the batch size.
+Eventually in 4.2 we will combine Theorem 4.1 below with the privacy analysis of Algorithm 1 to arrive at the final DP-SCO guarantee, with the appropriate choice of the batch gradient steps, and the batch size.
 In particular, there we instantiate the noise random variables $\{\mathbf{b}_{t}\}$’s with the (by now standard) binary tree mechanism [16, 19].
@@
 For an appropriate setting of $\eta_{t},\tau_{t}$ and $\beta\geq M$, suppose $\mathbb{E}[\left\|\Delta_{t}-\mathbb{E}[\Delta_{t}]\right\|_{2}^{2}]\leq S^{2}$.
-Then, over the randomness of the batches $\mathcal{B}$, the following is true for Algorithm 1:
+Then, with high probability over the randomness of the batches $\mathcal{B}$, the following is true for Algorithm 1:
 $\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=$
-$O\left(\frac{SL+S^{2}}{\beta\sqrt{T}}+\frac{S\left\|\mathcal{C}\right\|_{2}}{T^{1.5}}+\frac{(S+L)b_{max}}{T}+\frac{\beta b_{max}\left\|\mathcal{C}\right\|_{2}}{T^{2}}+\frac{\beta}{T^{2}}\left\|\mathcal{C}\right\|_{2}^{2}\right).$
-We defer the proof to Section 4.4.
-Intrpretation of 4.1: To understand the bound in Theorem 4.1, we can view it as three groups of terms.
-The first group is $\frac{SL+S^{2}}{\beta\sqrt{T}}+\frac{S\left\|\mathcal{C}\right\|_{2}}{T^{1.5}}$, which is the error due to stochasticity of the gradients.
-The second is $\frac{(S+L)b_{max}}{T}+\frac{\beta b_{max}\left\|\mathcal{C}\right\|_{2}}{T^{2}}$, which is the error due to noise added for privacy.
-The last is $\frac{\beta}{T^{2}}\left\|\mathcal{C}\right\|_{2}^{2}$, which is the optimization error, i.e.
+$\widetilde{O}\left(\frac{SL+S^{2}}{\beta\sqrt{T}}+\frac{SR}{T^{3/2}}+\frac{(S+L)b_{max}}{T}+\frac{\beta b_{max}R}{T^{2}}+\frac{\beta}{T^{2}}R^{2}\right).$
+Interpretation of Theorem 4.1: To understand the bound in Theorem 4.1, we can view it as three groups of terms.
+The first group is $\frac{SL+S^{2}}{\beta\sqrt{T}}+\frac{SR}{T^{3/2}}$, which is the error due to stochasticity of the gradients.
+The second is $\frac{(S+L)b_{max}}{T}+\frac{\beta b_{max}R}{T^{2}}$, which is the error due to noise added for privacy.
+The last is $\frac{\beta}{T^{2}}R^{2}$, which is the optimization error, i.e.
 the error of the algorithm without stochasticity or DP noise.
 Notice that the optimization error matches that of the non-stochastic part of non-private Nesterov’s acceleration [36].
-Without privacy, the stochastic part of Nesterov’s acceleration depends on $\frac{V}{\beta\sqrt{T}}$ [41], where $V$ bounds stochasticity in evaluating the population gradient: $\mathbb{E}_{\mathcal{B}_{t}}\left[\left\|\frac{1}{|\mathcal{B}_{t}|}\sum\limits_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x};d)-\nabla F(\mathbf{x};d)\right\|_{2}^{2}\right]\leq V^{2}$.
-Morally, the parameter $S$ in 4.1 can be compared to $V$.
+Without privacy, the stochastic part of Nesterov’s acceleration depends on $\frac{V}{\beta\sqrt{T}}$ [36], where $V$ bounds stochasticity in evaluating the population gradient:
+$\mathbb{E}_{\mathcal{B}_{t}}\left[\left\|\frac{1}{|\mathcal{B}_{t}|}\sum\limits_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x};d)-\nabla F(\mathbf{x};d)\right\|_{2}^{2}\right]\leq V^{2}.$
+Morally, the parameter $S$ in Theorem 4.1 can be compared to $V$.
 However, $S$ encodes the additional variance due to stochastic recursive gradients.
-Modulo, the difference between $S$ and $V$4.1 cleanly decomposes the non-private stochastic Nesterov’s utility and the additional error due to the explicit noise introduced due to DP.
-Proof overview: The proof is deferred to Section 4.4. We give a high-level overview here. We define the following potential. Let $\eta_{0:t}=\sum\limits_{i=0}^{t}\eta_{i}$, then
+Modulo the difference between $S$ and $V$, Theorem 4.1 cleanly decomposes the non-private stochastic Nesterov’s utility and the additional error due to the explicit noise introduced due to DP.
+Proof overview: We define the following potential. Let $\eta_{0:t}=\sum\limits_{i=0}^{t}\eta_{i}$, then
 $\Phi_{t}=\eta_{0:t-1}\left(F(\mathbf{y}_{t})-F(\mathbf{x}^{*})\right)+2\beta\left\|\mathbf{z}_{t}-\mathbf{x}^{*}\right\|_{2}^{2}.$
@@
 The analysis can easily be extended to the standard setting where each $\nabla_{t}$ is an independent batch gradient.
-We give this analysis in Section 4.4.3.
+We give this analysis in Section 5.5.
 To the best of our knowledge, a direct potential-based analysis of stochastic accelerated gradient descent has not appeared in the literature before besides [47].
 Our analysis is simpler than that of [47], so it may be of independent interest.
-4.2 Privacy analysis
-Given Theorem 4.1, our goal now is to bound $S$, which is equivalent to bounding the sensitivity of the $\Delta_{t}$, and then to show that there is a noise distribution for which $b_{max}$ is sufficiently small (with high probability) and we achieve $(\varepsilon,\delta)$-DP.
-We can then set $\beta$ to minimize the error bound, and determine the largest batch size that doesn’t cause the error bound to increase.
-We first prove the sensitivity bound on each $\Delta_{t}$.
-The sensitivity bound requires a bound on the max noise we add in any iteration.
-The proof of sensitivity follows a similar structure to that of [55, Lemma 15].
-However, because the gradients $\nabla f_{t}(\mathbf{x}_{t})$ are evaluated at a different point than that in [55], the exact argument is different.
-Claim 4.2 (Bounding $\ell_{2}$-sensitivity).
-In Algorithm 1, suppose for all $t$, $\left\|\mathbf{b}_{t}\right\|_{2}\leq b_{\max}$. Then for $\eta_{t}=t+1$, $\beta\geq 2MT$ for any $d\in\mathcal{D}$, $t\in\{0,1,\ldots,T-1\}$
-$\left\|\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)\right\|_{2}\leq 2(Mb_{\max}+2M\left\|\mathcal{C}\right\|_{2}+L).$
-Proof.
-For convenience let $\Delta_{t}(d)=\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$.
-The bound holds by Lipschitzness for $t=0$.
-We prove the bound for $t>0$ by induction.
-We have:
-$\Delta_{t}(d)$ $=\eta_{t}\left(\nabla f(\mathbf{x}_{t},d)-\nabla f(\mathbf{x}_{t-1},d)\right)+\left(\eta_{t}-\eta_{t-1}\right)\nabla f(\mathbf{x}_{t},d)$
-$\Rightarrow\left\|\Delta_{t}(d)\right\|_{2}$ $\leq\eta_{t}\cdot M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}+\left|\eta_{t}-\eta_{t-1}\right|\cdot L.$ (1)
-We now bound each of the terms in the R.H.S.
-of Equation 1 independently.
-Since $\eta_{t}=t+1$, $\eta_{t}-\eta_{t-1}=1$.
-In the following, we bound $\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$.
-We have,
-$\eta_{0:t}\cdot\mathbf{x}_{t}$ $=\eta_{0:t-1}\cdot\mathbf{y}_{t}+\eta_{t}\cdot\mathbf{z}_{t}$
-$\Leftrightarrow\eta_{0:t}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $=\left\|\eta_{0:t-1}(\mathbf{y}_{t}-\mathbf{x}_{t-1})+\eta_{t}(\mathbf{z}_{t}-\mathbf{x}_{t-1})\right\|_{2}$
-$\leq\eta_{0:t-1}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|_{2}+\eta_{t}\left\|\mathbf{z}_{t}-\mathbf{x}_{t-1}\right\|_{2}$
-$\leq\eta_{0:t-1}\left\|\frac{\nabla_{t-1}}{\beta}+\frac{\mathbf{b}_{t-1}}{\eta_{t}}\right\|_{2}+\eta_{t}\left\|\mathcal{C}\right\|_{2}$
-$\Rightarrow\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{\eta_{0:t-1}}{\eta_{0:t}}\cdot\frac{\left\|\nabla_{t-1}\right\|_{2}}{\beta}+\frac{\eta_{0:t-1}\left\|\mathbf{b}_{t-1}\right\|_{2}}{\eta_{t}\cdot\eta_{0:t}}+\frac{\eta_{t}}{\eta_{0:t}}\left\|\mathcal{C}\right\|_{2}$
-$\leq\frac{\sum_{i\leq t-1}\sum_{d\in\mathcal{B}_{i}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta(t+1)B}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}\left\|\mathcal{C}\right\|_{2}$
-$\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}\left\|\mathcal{C}\right\|_{2}$
-$\Rightarrow\eta_{t}\cdot M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{\eta_{t}M\cdot\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta}+Mb_{\max}+2M\left\|\mathcal{C}\right\|_{2}$
-$\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{2}+Mb_{\max}+2M\left\|\mathcal{C}\right\|_{2}.$ (2)
-Plugging Section 4.2 and the fact that $\eta_{t}-\eta_{t-1}=1$ into Equation 1, we have the following:
-$\left\|\Delta_{t}(d)\right\|_{2}\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{2}+Mb_{max}+2M\left\|\mathcal{C}\right\|_{2}+L.$ (3)
-By induction, this completes the proof of 4.2. ∎
-A crucial aspect of 4.2 is that conditioned on $b_{max}$, the sensitivity is independent of $t$, even if the norm of the quantity that is being approximated (i.e., $\eta_{t}\nabla f(\mathbf{x}_{t};d)$) scales with $t$.
-Deciding the noise scale for $(\varepsilon,\delta)$-DP: We now show how to sample $\mathbf{b}_{t}$ such that Algorithm 1 satisfies $(\varepsilon,\delta)$-DP while achieving strong utility guarantees.
-Each $\nabla_{t}$ is a (rescaling of a) prefix sum of the $\Delta_{t}$ values.
-So, choosing a noise mechanism that does not add too much noise to the gradients in Algorithm 1 is equivalent to privatizing the list of prefix sums $\{\Delta_{1},\Delta_{1}+\Delta_{2},\ldots,\Delta_{1}+\ldots+\Delta_{T}\}$, where by Claim 4.2, we know the sensitivity of each $\Delta_{t}$ is constant with respect to $t$.
-This is a well studied problem known as private continual counting, and a standard solution to this problem is the binary tree mechanism [24].
-Lemma 4.3.
-Let $\mathbf{g}_{i}=\sum_{j\leq i}\Delta_{i}$ for $i\in[T]$ and suppose the $\Delta_{i}$ are computed such that between any two adjacent datasets, only one $\Delta_{i}$ changes, and it changes by $\ell_{2}$-norm at most $C$.
-Then for any $T\geq 2$ there is a $\mu$-GDP algorithm (i.e., an algorithm which satisfies any DP guarantee satisfied by the Gaussian mechanism with sensitivity $\mu$ and variance 1) that gives an unbiased estimate of each $\mathbf{g}_{i}$ such that with probability $1-\delta$ all $\mathbf{g}_{i}$’s estimates have $\ell_{2}$-error at most $\frac{4C\log^{1.5}T\sqrt{d\ln(2T/\delta)}}{\mu}$.
-Proof.
-For completeness, we define the binary tree mechanism and state its guarantees here.
-For $k=0,1,2,\ldots,\lceil\log T\rceil$, $j=1,3,5,\ldots T/2^{k}-1$, the binary tree mechanism computes $s_{j,k}=\sum_{i=(j-1)\cdot 2^{k}+1}^{j\cdot 2^{k}}\Delta_{i}+\xi_{j,k},\xi_{j,k}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,\sigma^{2}\mathbb{I})$.
-An unbiased estimate of any prefix sum can be formed by adding at most $\lceil\log T\rceil$ of the $s_{j,k}$.
-For example, the sum of $\Delta_{1}$ to $\Delta_{7}$ has an unbiased estimator using at most $\lceil\log 7\rceil=3$ of the $s_{j,k}$, $s_{1,2}+s_{3,1}+s_{7,0}$.
-Each $\Delta_{i}$ contributes to at most one $s_{j,k}$ for each value of $k$, i.e.
-at most a total of $1+\lceil\log T\rceil$ values of $s_{j,k}$.
-So, if $\left\|\Delta_{i}\right\|_{2}\leq C$ for all $i$, then with $\sigma=C\sqrt{1+\lceil\log T\rceil}/\mu$ we can release all $s_{j,k}$ with $\mu$-GDP.
-Furthermore, by a standard Gaussian tail bound with probability $21-\delta/$ 2 we have that $\left\|\xi_{i}\right\|_{2}\leq\sigma\sqrt{2d\ln(2T/\delta)}$ for all $i$.
-Since each $\mathbf{g}_{i}$ has an unbiased estimator formed by adding at most $\lceil\log T\rceil$ of the $s_{j,k}$, the maximum additive error is then at most
-$\lceil\log T\rceil\max_{i}\left\|\xi_{i}\right\|_{2}\leq 2C\log T\sqrt{2d(1+\lceil\log T\rceil)\ln(2T/\delta)}/\mu\leq 4C\log^{1.5}T\sqrt{d\ln(2T/\delta)}/\mu.$
-4.3 gives us a high probability bound on the max noise, which we can then plug into 4.2 to show that if the dimension (and thus the norm of the noise) is not too large, the noise term in 4.2 is not dominating:
-Corollary 4.4.
-Suppose we compute $\mathbf{b}_{i}$ in Algorithm 1 using the $\mu$-GDP binary tree mechanism (Lemma 4.3) with $C=\frac{4L+8M\left\|\mathcal{C}\right\|_{2}}{B\beta}$.
-Then for $d\leq\frac{B^{2}\beta^{2}\mu^{2}}{64M^{2}\log^{3}(T)\ln(4T/\delta)}$ and the parameter choices in Claim 4.2, with probability $21-\delta/$ 2, each $\Delta_{t}$ has $\ell_{2}$-sensitivity at most $\leq 4L+8M\left\|\mathcal{C}\right\|_{2}$ for all $i$.
-Our eventual choice of $\beta$ will be proportional to $n^{3/2}/B^{2}$, i.e.
-the upper bound on $d$ is proportional to $n^{3}\varepsilon^{2}/B^{2}$.
-We can rewrite this as an upper bound on the batch size: $B\leq n^{1.5}\varepsilon/\sqrt{d}$.
-Ignoring dependence on logarithmic terms and $\delta$, when $d\geq n^{2}\varepsilon^{2}$, no $(\varepsilon,\delta)$-DP algorithm can achieve expected excess loss, $o(L\left\|\mathcal{C}\right\|_{2})$, which is achieved by any point in $\mathcal{C}$.
-So this upper bound is vacuous in the setting where $d\leq n$ and $\varepsilon=\Omega(1)$, and never worse than $\sqrt{n}$ for interesting settings of $\delta$.
-We can now convert the sensitivity bound and the privacy guarantees of the binary tree mechanism into a privacy guarantee for Algorithm 1:
-Corollary 4.5.
-Suppose we compute $\mathbf{b}_{i}$ in Algorithm 1 using the binary tree mechanism (Lemma 4.3) with $\sigma=\frac{(8\sqrt{2}L+16\sqrt{2}M\left\|\mathcal{C}\right\|_{2})\sqrt{\log(T)\ln(2.5/\delta)}}{\varepsilon B\beta}$ on $\Delta_{i}$.
-Then for $d\leq\frac{B^{2}\beta^{2}\varepsilon^{2}}{128M^{2}\log^{3}(T)\ln(4T/\delta)\ln(2.5/\delta)}$ and the parameter choices in Claim 4.2, Algorithm 1 satisfies $(\varepsilon,\delta)$-DP.
-Proof.
-If we run Algorithm 1 but instead clip $\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$ for each $t$ and $d$ in $\mathcal{B}_{t}$ to norm at most $4L+8M\left\|\mathcal{C}\right\|_{2}$ if its norm exceeds $4L+8M\left\|\mathcal{C}\right\|_{2}$, still using the binary tree mechanism to generate the $\mathbf{b}_{i}$, then Algorithm 1 satisfies $\varepsilon/\sqrt{2\ln(2.5/\delta)}$-GDP which implies it satisfies $(\varepsilon,\delta/2)$-DP.
-Algorithm 1 without clipping is within total variation distance $\delta/2$ of the version with clipping by Corollary 4.4, i.e.
-Algorithm 1 satisfies $(\varepsilon,\delta$)-DP.
-∎
-Finally, we plug the error guarantees of the binary tree mechanism into Theorem 4.7 to get our main result:
-Theorem 4.6.
-[Simplification of Theorem 4.10 in Section 4.4] Algorithm 1 with batch size $B=\sqrt{n}\cdot\min\{1,M\}$ and $T=n/B$ (i.e., a single epoch over a dataset of size $n$), and an appropriate choice of $\eta_{t},\beta,\sigma$ is $(\varepsilon,\delta)$-DP and has the following excess population loss guarantee:
-$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left((L\left\|\mathcal{C}\right\|_{2}+M\left\|\mathcal{C}\right\|_{2}^{2})\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right)\right).$
+4.2 Privacy analysis and DP-SCO bounds
+Bounding gradient differences: We next need to bound the norms of the gradient differences for a single example $\Delta_{t}(d):=\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$.
+This gives us a bound on the variance $S^{2}$ in the loss bound of Theorem 4.1, and a bound on the sensitivity of the $\Delta_{t}$, which will let us bound the noise $\mathbf{b}_{t}$ needed for privacy.
+By smoothness and Lipschitzness, for any $d$ we have:
+$\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$ $=\eta_{t}(\nabla f(\mathbf{x}_{t},d)-\nabla f(\mathbf{x}_{t-1},d))+(\eta_{t}-\eta_{t-1})\nabla f(\mathbf{x}_{t-1},d)$
+$\leq\eta_{t}M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}+(\eta_{t}-\eta_{t-1})L.$
+We will choose $\eta_{t}=t+1$, so the second term is just $L$. The main challenge is then to bound the first term. We will choose $\tau_{t}\approx 1/t$ to be vanishing, so
+$\mathbf{x}_{t}=(1-\tau_{t})\mathbf{y}_{t}+\tau_{t}\mathbf{z}_{t}\approx\mathbf{y}_{t},$
+and in turn
+$\eta_{t}M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}\approx\eta_{t}M\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|_{2}\leq\left\|\frac{\eta_{t}M}{\beta}\cdot\nabla_{t}+\frac{M}{\beta}\mathbf{b}_{t}\right\|_{2}.$
+The term $\frac{M}{\beta}\mathbf{b}_{t}$ ends up having norm at most $L$ under the right parameter choices, so the main challenge is to bound $\frac{\eta_{t}M}{\beta}\cdot\nabla_{t}$.
+Our approach for bounding this term differs based on whether or not we assume $\mathbf{x}^{\dagger}\in\mathcal{C}$.
+If $\mathbf{x}^{\dagger}\in\mathcal{C}$, $\mathbf{x}^{*}=\mathbf{x}^{\dagger}$ and by smoothness we have
+$\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}\leq\sqrt{2M(F(\mathbf{x}_{t})-F(\mathbf{x}^{*}))}.$
+Then since $\nabla_{t}$ is a stochastic estimate of $\nabla F(\mathbf{x}_{t})$, by Theorem 4.1 and a concentration inequality on $\nabla_{t}$, the fact that $F(\mathbf{x}_{t})-F(\mathbf{x}^{*})$ implies a bound on $\left\|\nabla_{t}\right\|_{2}$ that is also decreasing.
+In turn, while $\eta_{t}$ is increasing, we are able to show $\frac{\eta_{t}M}{\beta}\cdot\nabla_{t}$ remains constant even with $\beta=O(M)$ (Lemma 5.5).
+A subtle challenge is that in order to obtain a bound on $\left\|\nabla_{t}\right\|_{2}$ that is decreasing, we actually employ a tighter version of Theorem 4.1 that uses the exact norms $\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}$ rather than the Lipschitz constant $L$ in the final bound.
+In turn, there is a circular dependence between the bound of Theorem 4.1 and the norm of the gradient differences.
+Because of the circular dependence, bounding both simultaneously requires a proof by induction with careful choice of parameters (Lemma 5.4).
+Without assuming $\mathbf{x}^{\dagger}\in\mathcal{C}$, it’s no longer the case that $\nabla_{t}$ is necessarily decreasing over time (e.g.
+in the case of linear losses, $\nabla_{t}$ will remain constant).
+In this case, we more directly bound $\frac{\eta_{t}M}{\beta}\cdot\nabla_{t}$ by choosing $\beta\geq TM\geq\eta_{t}M$, which ensures the gradient differences still have constant sensitivity (Lemma 5.9).
+The fact that we have to increase $\beta$ (i.e., slow down learning) means we need to increase $T$ to compensate in utility, leading to the increased iteration complexity $T=\sqrt{n}$ instead of $T\approx n^{1/4}$.
+Privatizing the gradient differences: Once we have a bound on $\Delta_{t}(d)$, we now need to bound $\max_{t}\left\|\mathbf{b}_{t}\right\|_{2}$, i.e.
+the maximum noise added to any gradient estimate $\nabla_{t}$.
+Since each $\nabla_{t}$ is a (rescaling of) prefix sums of $\Delta_{t}$, we do this using the binary tree mechanism of [19], which adds noise to the prefix sums with variance scaling as ${\rm polylog}\left(T\right)$ rather than $T$, what we would get by independently noising each $\Delta_{t}$.
+We can use existing results on the binary tree mechanism (for convenience, reproven as Lemma 5.6) to obtain bounds on the maximum noise $\max_{t}\left\|\mathbf{b}_{t}\right\|_{2}$ needed to achieve $(\varepsilon,\delta)$-DP.
+Finally, it is straightforward to get our main results by plugging in the bounds on $\max_{t}\left\|\mathbf{b}_{t}\right\|_{2}$ and $\Delta_{t}(d)$ into Theorem 4.1:
+Theorem 4.2 (Simplified version of Theorems 5.8 and 5.12).
+For appropriate settings of $\beta,T$, Algorithm 1 is $(\varepsilon,\delta)$-DP and satisfies:
+$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})=O\left(LR\cdot{\rm polylog}\left(n,1/\delta\right)\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{p}}{\varepsilon n}\right)\right).$
+If we assume $\mathbf{x}^{\dagger}\in\mathcal{C}$, then the parameter settings are such that $T=\widetilde{\Theta}(n^{1/4})$ and $BT=n$, i.e.
+only a single epoch is required.
+Without this assumption, the parameter settings are such that we can allow any $T\geq\sqrt{n}$ and $BT=n$, i.e.
+still only a single epoch is required and the batch size can be as large as $\sqrt{n}$.
 4.3 Unaccelerated Variant
@@
 We give a sketch that the unaccelerated version requires batch size $O(1)$.
-Suppose we use $\Delta_{t}=c_{t}\nabla f(\mathbf{x}_{t},\mathcal{B}_{t})-c_{t-1}\nabla f(\mathbf{x}_{t-1},\mathcal{B}_{t})$ as our gradient differences and $\nabla_{t}=\frac{\sum_{i\leq t}\Delta_{i}}{c_{t}}$ as our stochastic recursive gradients.
-For simplicity we will consider using SGD with a fixed learning rate $\eta$ as the underlying optimization algorithm, assume $L=\left\|\mathcal{C}\right\|_{2}=M=1$ and assume $c_{t}$ grows at most polynomially, which implies $\sum_{i\leq t}c_{i}^{2}\approx tc_{t}^{2}$.
-Similarly to Claim 4.2, we can show the sensitivity of each $\Delta_{t}$ is roughly $c_{t}\eta$.
-Suppose we use $\eta\approx\sqrt{B/T}=B/\sqrt{n}$, which is standard for interpolation between full-batch GD and SGD with $B=1$ (see e.g.
+Suppose we use $\Delta_{t}=\eta_{t}\nabla f(\mathbf{x}_{t},\mathcal{B}_{t})-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},\mathcal{B}_{t})$ as our gradient differences and $\nabla_{t}=\frac{\sum_{i\leq t}\Delta_{i}}{\eta_{t}}$ as our stochastic recursive gradients.
+For simplicity we will consider using SGD with a fixed learning rate $1/\beta$ as the underlying optimization algorithm, assume $L=R=M=1$ and assume $\eta_{t}$ grows at most polynomially, which implies $\sum_{i\leq t}\eta_{i}^{2}\approx t\eta_{t}^{2}$.
+Similarly to our sensitivity bound for the accelerated algorithm (Lemma 5.9), we can show the sensitivity of each $\Delta_{t}$ is roughly $\eta_{t}/\beta$.
+Suppose we use $1/\beta\approx\sqrt{B/T}=B/\sqrt{n}$, which is standard for interpolation between full-batch GD and SGD with $B=1$ (see e.g.
 Section 6 of [12]).
 We compute the vector variance $\mathbf{Var}\left(\mathbf{v}\right):=\mathbb{E}[\left\|\mathbf{v}-\mathbb{E}[\mathbf{v}]\right\|_{2}^{2}]$ of $\nabla_{t}$:
-$\mathbf{Var}\left(\nabla_{t}\right)=\mathbf{Var}\left(\frac{\sum_{i\leq t}\Delta_{t}}{c_{t}}\right)=\frac{\sum_{i\leq t}\mathbf{Var}\left(\Delta_{t}\right)}{c_{t}^{2}}=\frac{\sum_{i\leq t}c_{i}^{2}\eta^{2}}{c_{t}^{2}}=\frac{t\eta^{2}}{B}=\frac{tB}{n}.$
+$\mathbf{Var}\left(\nabla_{t}\right)=\mathbf{Var}\left(\frac{\sum_{i\leq t}\Delta_{t}}{\eta_{t}}\right)=\frac{\sum_{i\leq t}\mathbf{Var}\left(\Delta_{t}\right)}{\eta_{t}^{2}}=\frac{\sum_{i\leq t}\eta_{i}^{2}}{\eta_{t}^{2}\beta^{2}}=\frac{t}{B\beta^{2}}=\frac{tB}{n}.$
 For $t=\Omega(T)=\Omega(n/B)$, the variance is $\Omega(1)$.
@@
 with variance $O(1)$ rather than $O(1/B)$, requires $T=n$ iterations to converge, hence in the single-epoch setting we need $B=O(1)$.
-4.4 Deferred Proofs
-4.4.1 Proof of Theorem 4.1
-Theorem 4.1 is given by plugging $\eta_{t}={t+1}$ into the following more general theorem, dropping lower-order terms, setting $S=Q/\sqrt{B}$, and taking the asymptotic bound.
+5 Utility and Privacy Analysis
+In this section, we provide the detailed utility analysis and privacy analysis of Algorithm 1.
+5.1 Utility Analysis
+First, we provide the excess population risk guarantee for any choice of the privacy noise random variables $\{\mathbf{b}_{t}\}$’s.
+We only assume that the variance (over the sampling of the data samples $d\sim\mathcal{D}$) in any individual term in the SRGs (i.e., $\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$) is bounded.
 Theorem 5.1 (Excess population risk).
@@
 $\left\|(\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d))-\mathbb{E}_{d}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)]\right\|_{2}\leq Q.$
-Then, over the randomness of the i.i.d. sampling of the loss functions $\{f(\cdot;d_{t})\}$ from $\mathcal{D}$, the following is true for Algorithm 1:
-$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq$
-$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}\eta_{0:t}}+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)$
-$+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
+Then, w.p. $1-\gamma$ over the randomness of the i.i.d. sampling of the batches $\{\mathcal{B}_{t}\}$, the following is true for Algorithm 1:
+$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})\leq$
+$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{3(t+1)\log(6T/\gamma)}(\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}+\frac{Q\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right.$
+$\left.+\frac{3Q\sqrt{3(t+1)\log(6T/\gamma)}\left\|\mathbf{b}_{t}\right\|_{2}}{\sqrt{B}\eta_{t}^{2}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{0:t}}+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}\right.$
+$\left.+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
 Proof.
@@
 Lemma 5.2 (Potential drop).
-Let $\mathbf{q}_{t}=\nabla_{t}-\nabla F(\mathbf{x}_{t})$. Assuming $\eta_{t}^{2}\leq 4\eta_{0:t}$, the potential drop is bounded by the following (where $\Delta\Phi_{t}=\Phi_{t+1}-\Phi_{t}$):
-$\mathbb{E}\left[\Delta\Phi_{t}\right]\leq\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{t+1}}{\sqrt{B}\eta_{0:t}}\right.$
-$\left.+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right).$
-Telescopic summation and averaging: By telescoping the potential differences $\Delta\Phi_{t}$’s and linearity of expectation we have the following over the sampling of the functions $\{f_{t}\}$’s:
-$\mathbb{E}[\Phi_{T}-\Phi_{0}]\leq$ $\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{t+1}}{\sqrt{B}\eta_{0:t}}\right.$
+Let $\mathbf{q}_{t}=\nabla_{t}-\nabla F(\mathbf{x}_{t})$.
+Assuming $\eta_{t}^{2}\leq 4\eta_{0:t}$, w.p.
+$1-\gamma$ over $\{\mathcal{B}_{t}\}$ the potential drop is bounded by the following (where $\Delta\Phi_{t}=\Phi_{t+1}-\Phi_{t}$):
+$\Delta\Phi_{t}$ $\leq\eta_{0:t}\left(\frac{2Q\sqrt{3(t+1)\log(6/\gamma)}(\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}+\frac{Q\sqrt{3(t+1)\log(6/\gamma)}}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right.$
+$+\frac{3Q\sqrt{3(t+1)\log(6/\gamma)}\left\|\mathbf{b}_{t}\right\|_{2}}{\sqrt{B}\eta_{t}^{2}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{3(t+1)\log(6/\gamma)}}{\sqrt{B}\eta_{0:t}}$
+$+\left.\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right).$
+Telescopic summation and averaging: By telescoping the potential differences $\Delta\Phi_{t}$’s and a union bound over failure probabilities we have the following w.p.
+$1-\gamma$ over the sampling of the functions $\{f_{t}\}$’s:
+$\Phi_{T}-\Phi_{0}\leq$ $\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{3(t+1)\log(6T/\gamma)}(\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}+\frac{Q\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right.$
+$+\frac{3Q\sqrt{3(t+1)\log(6T/\gamma)}\left\|\mathbf{b}_{t}\right\|_{2}}{\sqrt{B}\eta_{t}^{2}}\left.+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{0:t}}\right.$ (2)
 $\left.+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right).$ (3)
-Therefore, from Equation 5 and definition of $\Phi_{t}$, as desired we have:
-$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq$
-$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{t+1}}{\sqrt{B}\eta_{0:t}}\right.$
-$\left.+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
+Therefore, from Equation 3 and definition of $\Phi_{t}$, as desired we have w.p. $1-\gamma$:
+$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})\leq$
+$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{3(t+1)\log(6T/\gamma)}(\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}+\frac{Q\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right.$
+$+\frac{3Q\sqrt{3(t+1)\log(6T/\gamma)}\left\|\mathbf{b}_{t}\right\|_{2}}{\sqrt{B}\eta_{t}^{2}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{3(t+1)\log(6T/\gamma)}}{\sqrt{B}\eta_{0:t}}+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}$
+$\left.+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
 This completes the proof of Theorem 5.1. ∎
-Proof of 4.8.
+Proof of Lemma 5.2.
 We have
@@
 $B+C$ $\leq\eta_{t}\left\langle\nabla_{t},\mathbf{z}_{t}-\mathbf{z}_{t+1}\right\rangle-2\beta\left\|\mathbf{z}_{t+1}-\mathbf{z}_{t}\right\|_{2}^{2}+2\beta\left\langle\mathbf{b}_{t},\mathbf{z}_{t+1}-\mathbf{x}^{*}\right\rangle-\eta_{t}\left\langle\mathbf{q}_{t},\mathbf{z}_{t}-\mathbf{x}^{*}\right\rangle$ (9)
-Using 4.9 below, and the fact that $\tau_{t}\leq 1$, we have the potential drop in 4.8 bound by the following.
+Using Lemma 5.3 below, and the fact that $\tau_{t}\leq 1$, we have the potential drop in Lemma 5.2 bound by the following.
 $\Delta\Phi_{t}\leq-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}-\eta_{t}\left\langle\mathbf{q}_{t},\mathbf{z}_{t}-\mathbf{x}^{*}\right\rangle$ (10)
 Recall, by definition $\mathbf{y}_{t+1}=\operatorname*{arg\,min}\limits_{\mathbf{y}\in\mathcal{C}}\left\langle\mathbf{q}_{t}+\nabla F(\mathbf{x}_{t})-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}-\mathbf{x}_{t}\right\rangle+\frac{\beta}{2}\left\|\mathbf{y}-\mathbf{x}_{t}\right\|_{2}^{2}$.
-Let $\mathbf{y}_{t+1}^{\prime}=\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}+\frac{1}{\eta_{t}}\cdot\mathbf{b}_{t}$ be the unconstrained minimizer, and let $J(\mathbf{y})=\left\langle\mathbf{q}_{t}+\nabla F(\mathbf{x}_{t})-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}\right\rangle+\frac{\beta}{2}\left\|\mathbf{y}\right\|_{2}^{2}$.
-By definition, $J(\mathbf{y}_{t+1})\geq J(\mathbf{y}^{\prime}_{t+1})+\frac{\beta}{2}\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\prime}_{t+1}\right\|_{2}^{2}$.
-Therefore,
+Let $\mathbf{y}_{t+1}^{\prime}=\mathbf{x}_{t}-\frac{1}{\beta}\nabla_{t}+\frac{1}{\eta_{t}}\cdot\mathbf{b}_{t}$ be the unconstrained minimizer, and let
+$J(\mathbf{y})=\left\langle\mathbf{q}_{t}+\nabla F(\mathbf{x}_{t})-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}\right\rangle+\frac{\beta}{2}\left\|\mathbf{y}\right\|_{2}^{2}.$
+By definition, $J(\mathbf{y}_{t+1})\geq J(\mathbf{y}^{\prime}_{t+1})+\frac{\beta}{2}\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\prime}_{t+1}\right\|_{2}^{2}$. Therefore,
 $\left\langle\mathbf{q}_{t}+\nabla F(\mathbf{x}_{t})-\frac{\beta}{\eta_{t}}\cdot\mathbf{b}_{t},\mathbf{y}_{t+1}-\mathbf{y}^{\prime}_{t+1}\right\rangle$ $\geq\frac{\beta}{2}\left\|\mathbf{y}_{t+1}-\mathbf{y}^{\prime}_{t+1}\right\|_{2}^{2}$
@@
 $\leq$ $\eta_{0:t}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathbf{y}_{t+1}-\mathbf{y}_{t+1}^{\prime}\right\|_{2}+\frac{\eta_{0:t}}{\beta}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\nabla_{t}\right\|_{2}+\frac{\eta_{0:t}}{\eta_{t}}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathbf{b}_{t}\right\|_{2}+\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}$
+$+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}+\eta_{t}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}$
+$\leq$ $\frac{3\eta_{0:t}}{\beta}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\nabla_{t}\right\|_{2}+\frac{3\eta_{0:t}}{\eta_{t}}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathbf{b}_{t}\right\|_{2}+\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}$
 $+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}+\eta_{t}\left\|\mathbf{q}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}.$ (12)
 Now, we observe that:
-$\mathbf{q}_{t}=\nabla_{t}-\nabla F(\mathbf{x}_{t})=\frac{\sum_{i\leq t}\sum_{d\in\mathcal{B}_{i}}((\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d))-\mathbb{E}_{d}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)])}{B\eta_{t}}.$
+$\mathbf{q}_{t}=\nabla_{t}-\nabla F(\mathbf{x}_{t})=\frac{\sum_{i\leq t}\sum_{d\in\mathcal{B}_{i}}((\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d))\\ -\mathbb{E}_{d}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)])}{B\eta_{t}}.$
 The sum $\sum_{i\leq t}\sum_{d\in\mathcal{B}_{i}}((\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d))-\mathbb{E}_{d}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)])$ is a mean-zero vector martingale with $B(t+1)$ terms each with norm at most $Q$ with probability 1 by the assumptions of the theorem.
-So we have
-$\mathbb{E}\left[\left\|\sum_{i\leq t}((\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d))-\mathbb{E}_{d}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)])\right\|_{2}^{2}\right]\leq Q^{2}B(t+1)$
-and thus $\mathbb{E}[\left\|\mathbf{q}_{t}\right\|_{2}^{2}]\leq\frac{Q^{2}(t+1)}{B\eta_{t}^{2}}$, $\mathbb{E}[\left\|\mathbf{q}_{t}\right\|_{2}]\leq\frac{Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}$ by Jensen’s inequality.
-We also unconditionally have $\left\|\mathbf{q}_{t}\right\|_{2}\leq Q(t+1)/\eta_{t}$ and thus $\left\|\nabla_{t}\right\|_{2}\leq L+Q(t+1)/\eta_{t}$ by triangle inequality.
-Plugging these bounds into (14) we have:
-$\mathbb{E}[\Delta\Phi_{t}]\leq\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}+\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{t+1}}{\sqrt{B}\eta_{0:t}}\right.$
+By concentration of bounded vector martingales (e.g., [27]), w.p.
+$1-\gamma$ we have
+$\left\|\sum_{i\leq t}\sum_{d\in\mathcal{B}_{i}}((\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d))-\mathbb{E}_{d}[\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)])\right\|_{2}^{2}\leq 3Q^{2}B(t+1)\log(6/\gamma).$
+and thus $\left\|\mathbf{q}_{t}\right\|_{2}\leq\frac{Q\sqrt{3(t+1)\log(6/\gamma)}}{\sqrt{B}\eta_{t}}$ with the same probability.
+Under this event we also have $\left\|\nabla_{t}\right\|_{2}\leq\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}+\frac{Q\sqrt{3(t+1)\log(6/\gamma)}}{\sqrt{B}\eta_{t}}$ by triangle inequality.
+Plugging these bounds into (12) we have w.p.
+$1-\gamma$:
+$\Delta\Phi_{t}\leq$ $\eta_{0:t}\left(\frac{2Q\sqrt{3(t+1)\log(6/\gamma)}(\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}+\frac{Q\sqrt{3(t+1)\log(6/\gamma)}}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right.$
+$\left.+\frac{3Q\sqrt{3(t+1)\log(6/\gamma)}\left\|\mathbf{b}_{t}\right\|_{2}}{\sqrt{B}\eta_{t}^{2}}+\frac{Q\left\|\mathcal{C}\right\|_{2}\sqrt{3(t+1)\log(6/\gamma)}}{\sqrt{B}\eta_{0:t}}\right.$
 $\left.+\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right).$
-Claim 4.9.
+Lemma 5.3.
 Assuming $\eta_{t}^{2}\leq 4\eta_{0:t}$. We have the following:
@@
 $-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\left(\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}+\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}\right).$
-Proof of 4.9.
+Proof of Lemma 5.3.
 By smoothness we have,
@@
 $\leq-\eta_{0:t}\left\langle\mathbf{q}_{t},\mathbf{y}_{t+1}-\mathbf{x}_{t}\right\rangle+\left(\frac{L\left\|\mathbf{b}_{t}\right\|_{2}}{\tau_{t}}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}\tau_{t}}+\beta\left\|\mathbf{b}_{t}\right\|_{2}\left\|\mathcal{C}\right\|_{2}\right).$ (18)
-This completes the proof of 4.9. ∎
-4.4.2 Proof of Theorem 4.6
-We give the more formal statement of Theorem 4.6 and its proof here:
+This completes the proof of Lemma 5.3. ∎
+5.2 Privacy analysis
+Given Theorem 4.1, our goal now is to bound $S$, which is equivalent to bounding the sensitivity of the $\Delta_{t}$, and then to show that there is a noise distribution for which $b_{max}$ is sufficiently small (with high probability) and we achieve $(\varepsilon,\delta)$-DP.
+Before proceeding with the proofs, we will specify the setting of parameters we will use in advance:
+$\eta_{t}$ $=t+1,$
+$c_{DP}$ $=4\sqrt{2}\log^{3/2}(n)\sqrt{\ln(2n/\delta)}\ln(2.5/\delta),$
+$\beta$ $=\frac{c_{DP}^{2}L\max\{1,\sqrt{p}/\varepsilon\sqrt{n}\}}{R},$
+$T$ $=c_{DP}\cdot n^{1/4},$
+$B$ $=n/T=n^{3/4}/c_{DP}.$ (19)
+5.2.1 Bounding gradient norms
+We first need to bound the gradient norms in each round:
+Lemma 5.4.
+Using the parameter choices in Section 5.2, additionally assume:
+  • Letting $\Delta_{t}(d):=\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$, for all $t\leq T-1$ and $d$, $\left\|\Delta_{t}(d)\right\|_{2}\leq c_{DP}L$.
+  • For all $t$: $\left\|\mathbf{b}_{t}\right\|_{2}\leq b_{max}:=\frac{c_{DP}^{2}L\sqrt{p}}{\varepsilon B\beta}$ (which under Section 5.2, is at most $Rc_{DP}/n^{1/4}\leq R$ for all sufficiently large $n$ assuming $\delta=2^{-o(n^{1/6})}$).
+Then, for any $t$, all sufficiently large $n$ and $\delta<1/n$, and some constant $c_{1}$ we have under the probability $1-\gamma$ events of Theorem 5.1:
+$\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}\leq c_{1}\sqrt{M}\left(\sqrt{\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B(t+1)}}}+\sqrt{\frac{\beta b_{max}R}{t+1}}+\frac{2\sqrt{\beta}R}{t+1}\right).$
+Proof.
+We proceed by induction.
+For all sufficiently large $n$, we have $c_{DP}^{2}\geq c_{M}$ and so $\beta\geq c_{DP}^{2}L/R\geq M$.
+So for $t=0$ we have by smoothness and the assumption that $\mathbf{x}^{*}=\mathbf{x}^{\dagger}$, i.e.
+$\left\|\nabla F(\mathbf{x}^{*})\right\|_{2}=0$:
+$\left\|\nabla F(\mathbf{x}_{0})\right\|_{2}\leq M\left\|\mathbf{x}_{0}-\mathbf{x}^{*}\right\|_{2}\leq MR\leq\sqrt{M\beta}R.$
+Hence the lemma holds for $t=0$ as long as $c_{1}\geq 1/2$.
+Now inductively assume the lemma holds for rounds $10,1,\ldots,t-$ 1.
+We can plug in $Q=2c_{DP}L$ into Theorem 5.1, and then using Section 5.2 and the assumptions of the lemma we can combine and simplify terms to show that for some constant $c_{2}$:
+$F(\mathbf{y}_{t})-F(\mathbf{x}^{*})\leq$ $c_{2}\left(\frac{1}{t^{2}}\sum\limits_{i=0}^{t-1}\frac{c_{DP}Li^{3/2}\sqrt{\log(T/\gamma)}\left\|\nabla F(\mathbf{x}_{i})\right\|_{2}}{\sqrt{B}\beta}+\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B(t+1)}}\right.$
+$\left.+\frac{\beta b_{max}R}{t+1}+\frac{\beta}{(t+1)^{2}}R^{2}\right).$
+By smoothness we have:
+$\left\|\nabla F(\mathbf{y}_{t})\right\|_{2}$ $\leq\sqrt{2M(F(\mathbf{y}_{t})-F(\mathbf{x}^{*}))}$
+$=\sqrt{2M}c_{2}\left(\sqrt{\frac{1}{t^{2}}\sum\limits_{i=0}^{t-1}\frac{c_{DP}Li^{3/2}\sqrt{\log(T/\gamma)}\left\|\nabla F(\mathbf{x}_{i})\right\|_{2}}{\sqrt{B}\beta}+\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B(t+1)}}+\frac{\beta b_{max}R}{t+1}+\frac{\beta R^{2}}{(t+1)^{2}}}\right).$
+Then by smoothness and definition of $\mathbf{x}_{t},\mathbf{y}_{t}$ we have:
+$\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}\leq$ $M\left\|\mathbf{x}_{t}-\mathbf{y}_{t}\right\|_{2}+\left\|\nabla F(\mathbf{y}_{t})\right\|_{2}\leq 2MR/t+\left\|\nabla F(\mathbf{y}_{t})\right\|_{2}\leq 2\sqrt{M\beta}R/t+\left\|\nabla F(\mathbf{y}_{t})\right\|_{2}$
+$\leq$ $\sqrt{2M}c_{2}\left(\sqrt{\frac{1}{t^{2}}\sum\limits_{i=0}^{t-1}\frac{c_{DP}Li^{3/2}\sqrt{\log(T/\gamma)}\left\|\nabla F(\mathbf{x}_{i})\right\|_{2}}{\sqrt{B}\beta}+\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B(t+1)}}+\frac{\beta b_{max}R}{t+1}+\frac{\beta R^{2}}{(t+1)^{2}}}\right.$
+$+\left.\frac{\sqrt{2\beta}R}{c_{2}t}\right)$ (20)
+[…]
+$\frac{c_{1}c_{DP}L\sqrt{M\log(T/\gamma)}}{t^{2}\sqrt{B}\beta}\sum\limits_{i=0}^{t-1}i^{3/2}\left(\sqrt{\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B(i+1)}}}+\sqrt{\frac{\beta b_{max}R}{i+1}}+\frac{2\sqrt{\beta}R}{i+1}\right)$
+$\leq$ $\frac{c_{1}c_{DP}L\sqrt{M\log(T/\gamma)}}{\sqrt{B}\beta}\left(\frac{4t^{1/4}}{9}\sqrt{\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B}}}+\frac{\sqrt{\beta b_{max}R}}{2}+\frac{4\sqrt{\beta}R}{3\sqrt{t}}\right)$
+$\leq$ $\frac{4c_{1}\sqrt{M}}{3}\left(\frac{(c_{DP}L)^{3/2}\sqrt{R}\log(T/\gamma)t^{1/4}}{B^{3/4}\beta}+\frac{c_{DP}L\sqrt{\log(T/\gamma)b_{max}R}}{\sqrt{B\beta}}+\frac{c_{DP}LR\sqrt{\log(T/\gamma)}}{\sqrt{B\beta t}}\right)$ (21)
+We now use the assumptions in the lemma to simplify further:
+$\frac{4c_{1}\sqrt{M}}{3}\left(\frac{(c_{DP}L)^{3/2}\sqrt{R}\log(T/\gamma)t^{1/4}}{B^{3/4}\beta}+\frac{c_{DP}L\sqrt{\log(T/\gamma)b_{max}R}}{\sqrt{B\beta}}+\frac{c_{DP}LR\sqrt{\log(T/\gamma)}}{\sqrt{B\beta t}}\right)$
+$\leq$ $\frac{4c_{1}c_{DP}\sqrt{M}LR\log(T/\gamma)}{3}\left(\frac{t^{1/4}}{B^{3/4}\sqrt{c_{DP}\beta}}+\frac{p^{1/4}\sqrt{c_{DP}Lc_{DP}}}{B\beta\sqrt{\varepsilon R}}+\frac{1}{\sqrt{B\beta t}}\right)$
+$\leq$ $\frac{4c_{1}\sqrt{c_{DP}}LR\log(T/\gamma)}{3}\left(\frac{t^{1/4}}{B^{3/4}c_{DP}\max\{1,\sqrt{p}/\varepsilon\sqrt{n}\}}\right.$
+$\left.+\frac{p^{1/4}\sqrt{M}}{B\max\{\sqrt{\varepsilon},p^{1/4}/n^{1/4}\}\sqrt{\beta c_{DP}}}+\frac{1}{\sqrt{c_{DP}Bt}}\right)$
+$\leq$ $\frac{4c_{1}\sqrt{c_{DP}}LR\log(T/\gamma)}{3}\left(\frac{T}{n^{3/4}c_{DP}}+\frac{n^{1/4}}{B\sqrt{c_{DP}}}+\frac{1}{\sqrt{c_{DP}Bt}}\right)$
+$\leq$ $\frac{4c_{1}\sqrt{c_{DP}}LR\log(T/\gamma)}{3}\left(\frac{2\sqrt{c_{DP}}}{n^{1/2}}+\frac{1}{\sqrt{c_{DP}Bt}}\right)$
+$\leq$ $\frac{4c_{1}\sqrt{c_{DP}}LR\log(T/\gamma)}{3}\left(\frac{2\sqrt{c_{DP}}}{n^{1/2}}+\frac{1}{\sqrt{c_{DP}Bt}}\right).$ (22)
+We can now plug Eq. 21 and Eq. 22 into Eq. 20, use subadditivity of the square root, and use the fact that for all sufficiently large $n$, $c_{DP}\geq c_{M}$:
+$\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}$
+$\leq$ $\sqrt{2M}c_{2}\left(\sqrt{\frac{8c_{1}c_{DP}LR\log(T/\gamma)}{3}\cdot\frac{1}{n^{1/2}}+\frac{(4c_{1}/3c_{DP}+1)c_{DP}LR\log(T/\gamma)}{\sqrt{B(t+1)}}+\frac{\beta b_{max}R}{t+1}+\frac{\beta R^{2}}{(t+1)^{2}}}+\frac{\sqrt{2\beta}R}{c_{2}t}\right)$
+$\leq$ $\sqrt{2M}c_{2}\left(\sqrt{(8c_{1}\sqrt{c_{\beta}}+\frac{4c_{1}}{3c_{DP}}+1)\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B(t+1)}}+\frac{\beta b_{max}R}{t+1}+\frac{\beta R^{2}}{(t+1)^{2}}}+\frac{\sqrt{2\beta}R}{c_{2}t}\right)$
+$\leq$ $\sqrt{2M}c_{2}\left(\sqrt{(8c_{1}+\frac{4c_{1}}{3c_{DP}}+1)\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B(t+1)}}}+\sqrt{\frac{\beta b_{max}R}{t+1}}+\sqrt{\frac{\beta R^{2}}{(t+1)^{2}}}+\frac{\sqrt{2\beta}R}{c_{2}t}\right)$
+$\leq$ $\sqrt{2M}c_{2}\left(\sqrt{(8c_{1}+\frac{4c_{1}}{3c_{DP}}+1)\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B(t+1)}}}+\sqrt{\frac{\beta b_{max}R}{t+1}}+\left(1+\frac{\sqrt{2}}{c_{2}}\right)\frac{\sqrt{\beta}R}{t}\right)$
+$\leq$ $\sqrt{M}\left(\sqrt{2}c_{2}\max\left\{\sqrt{8c_{1}+\frac{4c_{1}}{3c_{DP}}+1},1+\frac{\sqrt{2}}{c_{2}}\right\}\right)\cdot\left(\sqrt{\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B(t+1)}}}+\sqrt{\frac{\beta b_{max}R}{t+1}}+\frac{\sqrt{\beta}R}{t}\right)$
+The inductive step is now proven if $c_{1}$ is such that:
+$\sqrt{2}c_{2}\max\left\{\sqrt{8c_{1}+\frac{4c_{1}}{3c_{DP}}+1},1+\frac{\sqrt{2}}{c_{2}}\right\}\leq c_{1}$
+The left hand side has a sublinear dependence on $c_{1}$, and all other terms on the left hand side are constant except $1/c_{DP}$ which is decreasing in $n,1/\delta$.
+So for a sufficiently large choice of $c_{1}$ and all sufficiently large $n,\delta\leq 1/n$, the inductive step holds.
+∎
+5.3 Bounding sensitivity
+We now prove a sensitivity bound on each $\Delta_{t}$.
+The proof of sensitivity follows a similar structure to that of [50, Lemma 15].
+However, because the gradients $\nabla f_{t}(\mathbf{x}_{t})$ are evaluated at a different point than that in [50], the exact argument is different.
+In particular, we need the tight bound on gradient norms of Lemma 5.4 to ensure the sensitivity is constant while keeping $\beta$ roughly constant.
+Lemma 5.5 (Bounding $\ell_{2}$-sensitivity).
+In Algorithm 1, assume for all $t$ and $c_{1}$ as defined in Lemma 5.4,
+$\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}\leq c_{1}\sqrt{M}\left(\sqrt{\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B(t+1)}}}+\sqrt{\frac{\beta b_{max}R}{t+1}}+\frac{2\sqrt{\beta}R}{t+1}\right).$
+Then, using the parameters (5.2), for all sufficiently large $n$, $\gamma=\Omega(\delta)$, $\delta=2^{-o(n^{1/3})}$, $p\leq\varepsilon^{2}n^{2}$, under a probability $1-2\gamma$ event contained in the probability $1-\gamma$ event of Theorem 5.1, letting $\Delta_{t}(d):=\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$,
+$\forall t\leq T-1,d\in supp(\mathcal{D}):\left\|\Delta_{t}(d)\right\|_{2}\leq c_{DP}L.$
+Proof.
+We have for any $d$:
+$\Delta_{t}(d)$ $=\eta_{t}\left(\nabla f(\mathbf{x}_{t},d)-\nabla f(\mathbf{x}_{t-1},d)\right)+\left(\eta_{t}-\eta_{t-1}\right)\nabla f(\mathbf{x}_{t},d)$
+$\Rightarrow\left\|\Delta_{t}(d)\right\|_{2}$ $\leq\eta_{t}\cdot M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}+\left|\eta_{t}-\eta_{t-1}\right|\cdot L.$ (23)
+For $t\leq 1+(c_{DP}-1)/c_{M}$:
+$\eta_{t}\cdot M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}+\left|\eta_{t}-\eta_{t-1}\right|\cdot L$
+$\leq$ $\eta_{t}\cdot MR+L$
+$\leq$ $(\eta_{t}\cdot c_{M}+1)+L$
+$\leq$ $c_{DP}L$
+We prove the bound for $t>1+(c_{DP}-1)/c_{M}$ by induction.
+Assuming the bound holds for $10,1,\ldots,t-$ 1, we bound each of the terms in the R.H.S.
+of Equation 23 independently.
+Since $\eta_{t}=t+1$, $\eta_{t}-\eta_{t-1}=1$.
+In the following, we bound $\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$.
+We use the fact that using 5.2, for all sufficiently large $n$ we have $\beta\geq c_{DP}M$ to get:
+$\eta_{0:t}\cdot\mathbf{x}_{t}$ $=\eta_{0:t-1}\cdot\mathbf{y}_{t}+\eta_{t}\cdot\mathbf{z}_{t}$
+$\Leftrightarrow\eta_{0:t}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $=\left\|\eta_{0:t-1}(\mathbf{y}_{t}-\mathbf{x}_{t-1})+\eta_{t}(\mathbf{z}_{t}-\mathbf{x}_{t-1})\right\|_{2}$
+$\leq\eta_{0:t-1}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|_{2}+\eta_{t}\left\|\mathbf{z}_{t}-\mathbf{x}_{t-1}\right\|_{2}$
+$\leq\eta_{0:t-1}\left\|\frac{\nabla_{t-1}}{\beta}+\frac{\mathbf{b}_{t-1}}{\eta_{t}}\right\|_{2}+\eta_{t}R$
+$\Rightarrow\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{\eta_{0:t-1}}{\eta_{0:t}}\cdot\frac{\left\|\nabla F(\mathbf{x}_{t-1})\right\|_{2}}{\beta}$
+$+\frac{\eta_{0:t-1}}{\eta_{0:t}}\cdot\frac{\left\|\nabla_{t-1}-\nabla F(\mathbf{x}_{t-1})\right\|_{2}}{\beta}+\frac{\eta_{0:t-1}\left\|\mathbf{b}_{t-1}\right\|_{2}}{\eta_{t}\cdot\eta_{0:t}}+\frac{\eta_{t}}{\eta_{0:t}}R$
+$\stackrel{{\scriptstyle(\ast_{1})}}{{\leq}}\frac{c_{1}\sqrt{M}\left(\sqrt{\frac{c_{DP}LR\log(T/\gamma)}{\sqrt{B(t+1)}}}+\sqrt{\frac{\beta b_{max}R}{t+1}}+\frac{2\sqrt{\beta}R}{t+1}\right)}{\beta}$
+$+\frac{\left\|\nabla_{t-1}-\nabla F(\mathbf{x}_{t-1})\right\|_{2}}{\beta}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}R$
+$\leq\frac{c_{1}\sqrt{LR\log(T/\gamma)}}{\sqrt{\beta c_{DP}}(B(t+1))^{1/4}}+\frac{c_{1}\sqrt{b_{max}R}}{\sqrt{c_{DP}(t+1)}}+\frac{2c_{1}R}{(t+1)}$
+$+\frac{\left\|\nabla_{t-1}-\nabla F(\mathbf{x}_{t-1})\right\|_{2}}{\beta}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}R$
+$\Rightarrow\eta_{t}\cdot M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{c_{1}M(t+1)^{3/4}\sqrt{LR\log(T/\gamma)}}{\sqrt{\beta c_{DP}}B^{1/4}}+\frac{c_{1}M\sqrt{b_{max}R(t+1)}}{\sqrt{c_{DP}}}$
+$+\eta_{t}\left\|\nabla_{t-1}-\nabla F(\mathbf{x}_{t-1})\right\|_{2}+Mb_{\max}+(2c_{1}+2)MR.$ (24)
+$(\ast_{1})$ is by Lemma 5.4. We bound the first three terms individually. For the first term:
+$\frac{c_{1}M(t+1)^{3/4}\sqrt{LR\log(T/\gamma)}}{\sqrt{\beta c_{DP}}B^{1/4}}$
+$\leq$ $\frac{c_{1}MT^{3/4}\sqrt{LR\log(T/\gamma)}}{\sqrt{\beta c_{DP}}B^{1/4}}$
+$=$ $\frac{c_{1}MT\sqrt{LR\log(T/\gamma)}}{\sqrt{\beta c_{DP}}n^{1/4}}$
+$=$ $\frac{c_{1}MR\sqrt{\log(T/\gamma)}}{\sqrt{c_{DP}}\max\{1,\sqrt{p}/\varepsilon\sqrt{n}\}}$
+$\leq$ $c_{1}MR\sqrt{\log(T/\gamma)}.$
+For the second term:
+$\frac{c_{1}M\sqrt{b_{max}R(t+1)}}{\sqrt{c_{DP}}}$
+$=$ $c_{1}M\sqrt{\frac{c_{DP}LR\sqrt{p}(t+1)}{\varepsilon B\beta}}$
+$=$ $c_{1}MR\sqrt{\frac{\sqrt{p}(t+1)}{c_{DP}B\max\{\varepsilon,\sqrt{p}/\sqrt{n}\}}}$
+$\leq$ $c_{1}MR\sqrt{\frac{\sqrt{n}(t+1)}{c_{DP}B}}$
+$\leq$ $c_{1}MR\sqrt{\frac{\sqrt{n}T}{c_{DP}B}}$
+$=$ $c_{1}MR\sqrt{c_{DP}}.$
+For the third term:
+$\nabla_{t-1}-\nabla F(\mathbf{x}_{t-1})=\frac{\sum_{i\leq t-1}\sum_{d\in\mathcal{B}_{i}}(\Delta_{i}(d)-\mathbb{E}_{d\sim\mathcal{D}}[\Delta_{i}(d)])}{Bt}$
+By the inductive hypothesis and triangle inequality, this is a vector martingale that is the sum of $Bt$ mean-zero terms each with norm at most $2c_{DP}L$.
+By a vector Azuma inequality [27], w.p.
+$1-\gamma$ (over all timesteps) we have:
+$\left\|\nabla_{t-1}-\nabla F(\mathbf{x}_{t-1})\right\|_{2}\leq\frac{2c_{DP}L\sqrt{3\log(6T/\gamma)}}{\sqrt{Bt}}$
+Plugging these three bounds into Section 5.3 and the resulting bound into Equation 23, we have:
+$\left\|\Delta_{t}(d)\right\|_{2}$ $\leq c_{1}MR\sqrt{\log(T/\gamma)}+c_{1}MR\sqrt{c_{DP}}$
+$+\frac{2c_{DP}L\sqrt{6T\log(6T/\gamma)}}{\sqrt{B}}+Mb_{\max}+2(c_{1}+1)MR+L.$
+$\leq c_{1}MR\sqrt{\log(T/\gamma)}+c_{1}MR\sqrt{c_{DP}}$
+$+\frac{2c_{DP}^{2}L\sqrt{6\log(6T/\gamma)}}{n^{1/4}}+(2c_{1}+3)MR+L.$
+$\leq\left(c_{1}c_{M}\sqrt{\log(T/\gamma)}+c_{1}c_{M}\sqrt{c_{DP}}\right.$
+$\left.+\frac{2c_{DP}^{2}\sqrt{6\log(6T/\gamma)}}{n^{1/4}}+2c_{1}c_{M}+3c_{M}+1\right)\cdot L.$ (25)
+Section 5.3 is at most $c_{Q}L$ and hence the lemma is proven by induction as long as:
+$c_{1}c_{M}\sqrt{c_{DP}\log(T/\gamma)}+c_{1}c_{M}\sqrt{c_{DP}}+\frac{2c_{DP}^{2}\sqrt{6\log(6T/\gamma)}}{n^{1/4}}+2c_{1}c_{M}+3c_{M}+1\leq c_{DP}.$
+Since $c_{1},c_{M}$ are constants, we have $\sqrt{c_{DP}\log(T/\gamma)}=o(c_{DP})$ since $c_{DP}=\omega(\sqrt{\log(T/\gamma)})$ under the condition $\gamma=\Omega(\delta)$, and we have $\frac{2c_{DP}^{2}\sqrt{6\log(6T/\gamma)}}{n^{1/4}}=o(c_{DP})$ under the condition $\delta=2^{-o(n^{1/3})}$.
+Hence the left-hand side is $o(c_{DP})$, so for all sufficiently large $n$ the inductive step holds.
+A crucial aspect of Lemma 5.5 is that conditioned on $b_{max}$, the sensitivity is independent of $t$, even if the norm of the quantity that is being approximated (i.e., $\eta_{t}\nabla f(\mathbf{x}_{t};d)$) scales with $t$.
+Deciding the noise scale for $(\varepsilon,\delta)$-DP: We now show how to sample $\mathbf{b}_{t}$ such that Algorithm 1 satisfies $(\varepsilon,\delta)$-DP while achieving strong utility guarantees.
+Each $\nabla_{t}$ is a (rescaling of a) prefix sum of the $\Delta_{t}$ values.
+So, choosing a noise mechanism that does not add too much noise to the gradients in Algorithm 1 is equivalent to privatizing the list of prefix sums $\{\Delta_{1},\Delta_{1}+\Delta_{2},\ldots,\Delta_{1}+\ldots+\Delta_{T}\}$, where by Claim 5.5, we know the sensitivity of each $\Delta_{t}$ is constant with respect to $t$.
+This is a well studied problem known as private continual counting, and a standard solution to this problem is the binary tree mechanism [19].
+Lemma 5.6.
+Let $\mathbf{g}_{i}=\sum_{j\leq i}\Delta_{i}$ for $i\in[T]$ and suppose the $\Delta_{i}$ are computed such that between any two adjacent datasets, only one $\Delta_{i}$ changes, and it changes by $\ell_{2}$-norm at most $C$.
+Then for any $T\geq 2$ there is a $\mu$-GDP algorithm (i.e., an algorithm which satisfies any DP guarantee satisfied by the Gaussian mechanism with sensitivity $\mu$ and variance 1, including ) that gives an unbiased estimate of each $\mathbf{g}_{i}$ such that with probability $1-\delta$ all $\mathbf{g}_{i}$’s estimates have $\ell_{2}$-error at most $\frac{4C\log^{3/2}T\sqrt{p\ln(2T/\delta)}}{\mu}$.
+Proof.
+For completeness, we define the binary tree mechanism and state its guarantees here.
+For $k=0,1,2,\ldots,\lceil\log T\rceil$, $j=1,3,5,\ldots T/2^{k}-1$, the binary tree mechanism computes $s_{j,k}=\sum_{i=(j-1)\cdot 2^{k}+1}^{j\cdot 2^{k}}\Delta_{i}+\xi_{j,k},\xi_{j,k}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,\sigma^{2}\mathbb{I})$.
+An unbiased estimate of any prefix sum can be formed by adding at most $\lceil\log T\rceil$ of the $s_{j,k}$.
+For example, the sum of $\Delta_{1}$ to $\Delta_{7}$ has an unbiased estimator using at most $\lceil\log 7\rceil=3$ of the $s_{j,k}$, $s_{1,2}+s_{3,1}+s_{7,0}$.
+Each $\Delta_{i}$ contributes to at most one $s_{j,k}$ for each value of $k$, i.e.
+at most a total of $1+\lceil\log T\rceil$ values of $s_{j,k}$.
+So, if $\left\|\Delta_{i}\right\|_{2}\leq C$ for all $i$, then with $\sigma=C\sqrt{1+\lceil\log T\rceil}/\mu$ we can release all $s_{j,k}$ with $\mu$-GDP.
+Furthermore, by a standard Gaussian tail bound with probability $21-\delta/$ 2 we have that $\left\|\xi_{i}\right\|_{2}\leq\sigma\sqrt{2p\ln(2T/\delta)}$ for all $i$.
+Since each $\mathbf{g}_{i}$ has an unbiased estimator formed by adding at most $\lceil\log T\rceil$ of the $s_{j,k}$, the maximum additive error is then at most
+$\lceil\log T\rceil\max_{i}\left\|\xi_{i}\right\|_{2}\leq 2C\log T\sqrt{2p(1+\lceil\log T\rceil)\ln(2T/\delta)}/\mu\leq 4C\log^{3/2}T\sqrt{p\ln(2T/\delta)}/\mu.$
+Corollary 5.7.
+Suppose we compute $\mathbf{b}_{i}$ in Algorithm 1 using the binary tree mechanism (Lemma 5.6) with $\sigma=\frac{c_{DP}L\sqrt{\log(T)\ln(2.5/\delta)}}{\varepsilon B\beta}$ to compute prefix sums of the $\Delta_{i}$.
+Then for the parameter choices in Claim 5.5, Algorithm 1 satisfies $(\varepsilon,\delta)$-DP and we have w.p.
+$1-\delta$:
+$\max_{t}\left\|\mathbf{b}_{t}\right\|_{2}\leq c_{DP}^{2}L\sqrt{p}/\varepsilon B\beta.$
+Proof.
+If we run Algorithm 1 using the binary tree mechanism to generate the $\mathbf{b}_{i}$ but instead clip $\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$ for each $t$ and $d$ in $\mathcal{B}_{t}$ to norm at most $c_{DP}L$ if its norm exceeds $c_{DP}L$, then by Corollary 5.7 Algorithm 1 satisfies $\varepsilon/\sqrt{2\ln(2.5/\delta)}$-GDP which implies it satisfies $(\varepsilon,\delta/2)$-DP [21].
+Algorithm 1 without clipping is within total variation distance $\delta/2$ of the version with clipping by Lemma 5.4 and Lemma 5.5, i.e.
+Algorithm 1 satisfies $(\varepsilon,\delta$)-DP.
+The bound on $\left\|\mathbf{b}_{t}\right\|_{2}$ also follows by Lemma 5.6.
+∎
+Finally, we plug the error guarantees of the binary tree mechanism into Theorem 5.1 to get our main result:
 Theorem 5.8.
-Algorithm 1 with batch size
-$B\leq\min\left\{\sqrt{n},\frac{(L+2M\left\|\mathcal{C}\right\|_{2})n^{1.5}\varepsilon}{4\sqrt{2}\left\|\mathcal{C}\right\|_{2}\sqrt{d}\log^{1.5}(T)\sqrt{\ln(4T/\delta)\ln(2.5)/\delta}}\right\}$
-and $T=n/B$ (i.e., a single pass over a dataset of size $n$), $\eta_{t}=t+1$, $\beta=M+\frac{(8L+16M\left\|\mathcal{C}\right\|_{2})n^{3/2}}{\left\|\mathcal{C}\right\|_{2}B^{2}}$, $\sigma=\frac{(8\sqrt{2}L+16\sqrt{2}M\left\|\mathcal{C}\right\|_{2})\sqrt{\log(T)\ln(2.5/\delta)}}{\varepsilon B\beta}$ is $(\varepsilon,\delta)$-DP and has the following excess population loss guarantee:
+For $2^{-o(n^{1/6})}\leq\delta\leq 1/n,p\leq\varepsilon^{2}n^{2}$, Algorithm 1 with parameters given by Section 5.2 and using the binary tree mechanism to sample $\mathbf{b}_{i}$ is $(\varepsilon,\delta)$-DP and has the following excess population risk guarantee:
+$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left(LR\cdot\text{polylog}(n,1/\delta)\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{p}}{\varepsilon n}\right)\right).$
+Before we prove the theorem, we note that both the lower bound on $\delta$ and upper bound on $p$ are vacuous: If $p\geq\varepsilon^{2}n^{2}$, then $LR\sqrt{p}/\varepsilon n\geq LR$, and $LR$ is the excess risk achieved by picking any point in $\mathcal{C}$, i.e.
+the excess risk guarantee in the Theorem is trivially achieved.
+Similarly, if $\delta=2^{-\Omega(n^{1/6})}$, then $\sqrt{n}\leq{\rm polylog}\left(1/\delta\right)$ and again the excess risk guarantee in the Theorem is trivially achieved.
+Proof.
+In Lemma 5.5 we chose $\gamma=\delta/4\leq 1/4n$, in which case the high probability events of Theorem 5.1, Lemma 5.4 and Lemma 5.5 happen w.p.
+at least $1-1/2n$.
+We give a bound conditioned on these events, which can differ from the unconditional final bound by at most $O(LR/n)$, which is a lower-order term.
+From Theorem 5.1, using $Q=2c_{DP}L$, $\left\|\mathbf{b}_{t}\right\|_{2}\leq b_{max}$, and the bound from Lemma 5.4 on $\left\|\nabla F(\mathbf{x}_{t})\right\|_{2}$, simplifying by suppressing terms logarithmic in $n,1/\delta$ we have conditioned on the event of Lemma 5.5:
+$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})=$
+$\widetilde{O}_{n,1/\delta}\left(\frac{1}{T^{2}}\sum\limits_{t=1}^{T}t^{2}\left(\frac{L\left(\sqrt{M}\left(\sqrt{\frac{LR}{\sqrt{Bt}}}+\sqrt{\frac{\beta b_{max}R}{t}}+\frac{\sqrt{\beta}R}{t}\right)+\frac{L}{\sqrt{Bt}}\right)}{\sqrt{Bt}\beta}\right.\right.$
+$\left.\left.+\frac{LR}{\sqrt{B}t^{3/2}}+\frac{\beta b_{max}R}{t^{2}}\right)+\frac{\beta}{T^{2}}R^{2}\right)$
+In (21) and (22) we bounded the first term inside the sum, we can reuse the bound to get:
+$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})$
+$=\widetilde{O}_{n,1/\delta}\left(\frac{LR}{\sqrt{n}}+\frac{1}{T^{2}}\sum\limits_{t=1}^{T}t^{2}\left(\frac{LR}{\sqrt{B}t^{1.5}}+\frac{\beta b_{max}R}{t^{2}}\right)+\frac{\beta}{T^{2}}R^{2}\right)$
+$=\widetilde{O}_{n,1/\delta}\left(\frac{LR}{\sqrt{n}}+\frac{LR}{\sqrt{BT}}+\frac{\beta b_{max}R}{T}+\frac{\beta}{T^{2}}R^{2}\right)$
+$=\widetilde{O}_{n,1/\delta}\left(\frac{LR}{\sqrt{n}}+\frac{\beta b_{max}R}{T}+\frac{\beta}{T^{2}}R^{2}\right).$
+Finally, we plug in the parameter values from (5.2) and use the fact that $\beta b_{max}\leq c_{DP}^{2}L\sqrt{p}/\varepsilon B$ to get:
+$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})=\widetilde{O}_{n,1/\delta}\left(LR\left({\sqrt{n}}+\frac{\sqrt{p}}{\varepsilon n}\right)\right).$
+5.4 Removing the Assumption $\mathbf{x}^{\dagger}\in\mathcal{C}$
+In this section we show how to modify our analysis to the setting where $\mathbf{x}^{\dagger}$ may lie outside the constraint set, i.e.
+the gradient at $\mathbf{x}^{*}$ is not necessarily 0.
+In this setting the gradient norm is no longer guaranteed to be decreasing in $t$, which causes our sensitivity to increase and also worsens our final utility bound.
+To undo both of these effects, we increase $\beta$ to be $\widetilde{\Theta}(MT)$ instead of $\widetilde{\Theta}(M)$, and consequently need to raise $T$ to be $\sqrt{n}$ instead of $\widetilde{\Theta}(n^{1/4})$ to ensure the optimization error $\beta R^{2}/T^{2}$ is $O(1/\sqrt{n})$.
+Lemma 5.9 (Bounding $\ell_{2}$-sensitivity).
+In Algorithm 1, suppose for all $t$, $\left\|\mathbf{b}_{t}\right\|_{2}\leq b_{\max}$. Then for $\eta_{t}=t+1$, $\beta\geq 2MT$ for any $d\in\mathcal{D}$, $t\in\{0,1,\ldots,T-1\}$
+$\left\|\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)\right\|_{2}\leq 2(Mb_{\max}+2MR+L)\leq 2Mb_{max}+(2c_{M}+1)L.$
+Proof.
+For convenience let $\Delta_{t}(d)=\eta_{t}\nabla f(\mathbf{x}_{t},d)-\eta_{t-1}\nabla f(\mathbf{x}_{t-1},d)$.
+The bound holds by Lipschitzness for $t=0$.
+We prove the bound for $t>0$ by induction.
+We have:
+$\Delta_{t}(d)$ $=\eta_{t}\left(\nabla f(\mathbf{x}_{t},d)-\nabla f(\mathbf{x}_{t-1},d)\right)+\left(\eta_{t}-\eta_{t-1}\right)\nabla f(\mathbf{x}_{t},d)$
+$\Rightarrow\left\|\Delta_{t}(d)\right\|_{2}$ $\leq\eta_{t}\cdot M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}+\left|\eta_{t}-\eta_{t-1}\right|\cdot L.$ (26)
+We now bound each of the terms in the R.H.S.
+of Equation 26 independently.
+Since $\eta_{t}=t+1$, $\eta_{t}-\eta_{t-1}=1$.
+In the following, we bound $\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$.
+We have,
+$\eta_{0:t}\cdot\mathbf{x}_{t}$ $=\eta_{0:t-1}\cdot\mathbf{y}_{t}+\eta_{t}\cdot\mathbf{z}_{t}$
+$\Leftrightarrow\eta_{0:t}\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $=\left\|\eta_{0:t-1}(\mathbf{y}_{t}-\mathbf{x}_{t-1})+\eta_{t}(\mathbf{z}_{t}-\mathbf{x}_{t-1})\right\|_{2}$
+$\leq\eta_{0:t-1}\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|_{2}+\eta_{t}\left\|\mathbf{z}_{t}-\mathbf{x}_{t-1}\right\|_{2}$
+$\leq\eta_{0:t-1}\left\|\frac{\nabla_{t-1}}{\beta}+\frac{\mathbf{b}_{t-1}}{\eta_{t}}\right\|_{2}+\eta_{t}R$
+$\Rightarrow\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{\eta_{0:t-1}}{\eta_{0:t}}\cdot\frac{\left\|\nabla_{t-1}\right\|_{2}}{\beta}+\frac{\eta_{0:t-1}\left\|\mathbf{b}_{t-1}\right\|_{2}}{\eta_{t}\cdot\eta_{0:t}}+\frac{\eta_{t}}{\eta_{0:t}}R$
+$\leq\frac{\sum_{i\leq t-1}\sum_{d\in\mathcal{B}_{i}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta(t+1)B}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}R$
+$\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta}+\frac{1}{t+1}b_{\max}+\frac{2}{t+2}R$
+$\Rightarrow\eta_{t}\cdot M\left\|\mathbf{x}_{t}-\mathbf{x}_{t-1}\right\|_{2}$ $\leq\frac{\eta_{t}M\cdot\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{\beta}+Mb_{\max}+2MR$
+$\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{2}+Mb_{\max}+2MR.$ (27)
+Plugging Section 5.4 and the fact that $\eta_{t}-\eta_{t-1}=1$ into Equation 26, we have the following:
+$\left\|\Delta_{t}(d)\right\|_{2}\leq\frac{\max_{i\leq t-1,d\in\mathcal{D}}\left\|\Delta_{i}(d)\right\|_{2}}{2}+Mb_{max}+2MR+L.$ (28)
+By induction, this completes the proof of Lemma 5.9. ∎
+By using the high-probability upper bound on $\mathbf{b}_{t}$ combined with Lemma 5.9, we can show that $Mb_{max}$ is a lower-order term in Lemma 5.9:
+Corollary 5.10.
+Suppose we compute $\mathbf{b}_{i}$ in Algorithm 1 using the binary tree mechanism (Lemma 5.6) with $C=\frac{(8c_{M}+4)L}{B\beta}$.
+Then for $p\leq\frac{B^{2}\beta^{2}\varepsilon^{2}}{32M^{2}\log^{3}(T)\ln(2T/\delta)\ln(2.5/\delta)}$ with probability $21-\delta/$ 2, $\left\|\Delta_{t}(d)\right\|_{2}\leq(8c_{M}+4)L$ for all $d\in supp(\mathcal{D}),t$.
+Proof.
+The corollary follows from Lemma 5.9 if $b_{max}\leq(2c_{M}+1)L/M$. Plugging $C$ into the bound on $\max_{t}\left\|\mathbf{b}_{t}\right\|_{2}$ from Lemma 5.6, we have:
+$b_{max}\leq\frac{(8c_{M}+4)L\log^{3/2}T\sqrt{p\ln(2T/\delta)}}{B\beta\mu}.$
+Hence it suffices if:
+$\frac{(8c_{M}+4)L\log^{3/2}T\sqrt{p\ln(2T/\delta)}}{B\beta\mu}\leq(2c_{M}+1)L/M.$
+Rearranging, we get that it suffices if
+$p\leq\left(\frac{(2c_{M}+1)B\beta\mu}{M(8c_{M}+4)\log^{3/2}T\sqrt{\ln(2T/\delta)}}\right)^{2}=\frac{B^{2}\beta^{2}\mu^{2}}{16M^{2}\log^{3}(T)\ln(2T/\delta)}.$
+Then by a similar argument to Corollary 5.7, we have:
+Corollary 5.11.
+Suppose we compute $\mathbf{b}_{i}$ in Algorithm 1 using the binary tree mechanism (Lemma 5.6) with $\sigma=\frac{(16\sqrt{2}c_{M}+8\sqrt{2})L\sqrt{\log(T)\ln(2.5/\delta)}}{\varepsilon B\beta}$ on $\Delta_{i}$.
+Then for $p\leq\frac{B^{2}\beta^{2}\mu^{2}}{16M^{2}\log^{3}(T)\ln(2T/\delta)}$, Algorithm 1 satisfies $(\varepsilon,\delta)$-DP.
+While we require a bound on the dimension to ensure the norm of the noise is bounded, this bound will end up being nearly vacuous: Our choice of $\beta$ is proportional to $n^{3/2}B^{2}$, i.e.
+the bound we get is roughly $p\leq n^{3}\varepsilon^{2}/B^{2}$.
+We will choose $B\leq\sqrt{n}$ so it suffices if $p\leq\varepsilon^{2}n^{2}$; this is the range of dimensions where DP-SCO is non-trivial, as for $p>\varepsilon^{2}n^{2}$ any point in $\mathcal{C}$ achieves the (asymptotic) optimal loss bound.
+We now give our main result without the assumption $\mathbf{x}^{\dagger}\in\mathcal{C}$:
+Theorem 5.12.
+Algorithm 1 with batch size $B\leq\sqrt{n}$, $T=n/B$, $\eta_{t}=t+1$, $\beta=\frac{(16c_{M}+8)Ln^{3/2}}{RB^{2}}$, $\sigma=\frac{((16\sqrt{2}c_{M}+8\sqrt{2})L)\sqrt{\log(T)\ln(2.5/\delta)}}{\varepsilon B\beta}$ is $(\varepsilon,\delta)$-DP and has the following excess population loss guarantee:
 $\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=$
-$O\left(\frac{M\left\|\mathcal{C}\right\|_{2}^{2}B^{2}}{n^{2}}+\frac{(L+M\left\|\mathcal{C}\right\|_{2})\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}+\frac{(L+\frac{M\left\|\mathcal{C}\right\|_{2}}{\sqrt{B}})\left\|\mathcal{C}\right\|_{2}\sqrt{d}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right).$
-In particular, if we choose $B=\min\{\sqrt{n},Mn^{1.5}\varepsilon/(\sqrt{d}\cdot\text{polylog}(n/B,1/\delta))\}$ then Algorithm 1 requires
-$\max\{\sqrt{n},\sqrt{d}\cdot\text{polylog}(n/B,1/\delta)/M\varepsilon\sqrt{n}\}\}$
+$O\left(\frac{MR^{2}B^{2}}{n^{2}}+\frac{(L+MR)R}{\sqrt{n}}+\frac{(L+\frac{MR}{\sqrt{B}})R\sqrt{p}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right).$
+In particular, if we choose $B=\min\{\sqrt{n},Mn^{3/2}\varepsilon/(\sqrt{p}\cdot\text{polylog}(n/B,1/\delta))\}$ then Algorithm 1 requires
+$\max\{\sqrt{n},\sqrt{p}\cdot\text{polylog}(n/B,1/\delta)/M\varepsilon\sqrt{n}\}\}$
 batch gradients and:
-$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left((L\left\|\mathcal{C}\right\|_{2}+M\left\|\mathcal{C}\right\|_{2}^{2})\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right)\right).$
+$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left(LR\cdot\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{p}}{\varepsilon n}\cdot\text{polylog}(n/B,1/\delta)\right)\right).$
 Proof.
-Note that the upper bound on $B$ gives $\beta\geq 16Mn/B\geq 2MT$, satisfying the assumption in Claim 4.2, and also gives $d$ satisfying the assumption in Corollary 4.5.
-Let $Q=8L+16M\left\|\mathcal{C}\right\|_{2}=O(L+M\left\|\mathcal{C}\right\|_{2})$.
-We plug the tail bound of Lemma 4.3 into the error bound of Theorem 4.7, use a failure probability of $1/T^{3}$ in that bound (the increase in the expected error due to this low-probability event is $O(L\left\|\mathcal{C}\right\|_{2}/T^{3})$, which is a lower order term in our error bound) and choose $\eta_{t}=t+1,\beta=\frac{QT^{3/2}}{\left\|\mathcal{C}\right\|_{2}\sqrt{B}}$, let:
-$c_{DP}=\frac{\sqrt{d\log^{3}(T)\log(T/\delta)\log(1/\delta)}}{\varepsilon}$
-We can assume that $d=O(n^{2}\varepsilon^{2})$, i.e.
-$c_{DP}=O(n\cdot{\rm polylog}\left(T,1/\delta\right)$, as if $d=\omega(n^{2}\varepsilon^{2})$ then any point in $\mathcal{C}$ achieves the desired excess loss bound.
-Now we split the bound in Theorem 4.7 into six terms and bound them individually.
-Specifically, from Theorem 4.7 we have:
-$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq$
-$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\underbrace{\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}}_{E_{1}(t)}+\underbrace{\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}}_{E_{2}(t)}+\underbrace{\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}\eta_{0:t}}}_{E_{3}(t)}+\underbrace{\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}}_{E_{4}(t)}+\underbrace{\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}}_{E_{5}(t)}\right)$
-$+\underbrace{\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}}_{E_{6}}.$
-Now, using (i) $Q/\left\|C\right\|_{2}=O(M)$, (ii) $n=BT$, (iii) $B\leq\sqrt{n}$ (and thus $T\geq\sqrt{n}$) and (iv) $\left\|\mathbf{b}_{t}\right\|_{2}=O\left(\frac{Qc_{DP}}{B\beta}\right)$ by Lemma 4.3 and our choice of $\sigma$, we can bound the contribution of each of $E_{1},E_{2},E_{3},E_{4},E_{5},E_{6}$ individually:
-$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{1}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{2Q\sqrt{t+1}(L+\frac{Q(t+1)}{\sqrt{B}\eta_{t}})}{\sqrt{B}\eta_{t}\beta}\right)=O\left(\frac{Q\sqrt{T}(L+\frac{Q}{\sqrt{B}})}{\sqrt{B}\beta}\right)$
-$=O\left(\frac{\left\|\mathcal{C}\right\|_{2}(L+\frac{Q}{\sqrt{B}})}{T}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{T}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}\right),$
-$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{2}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{(\frac{3Q\sqrt{t+1}}{\sqrt{B}\eta_{t}}+L)\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}\right)=O\left(Q\left\|\mathbf{b}_{t}\right\|_{2}\right)=O\left(\frac{Q^{2}c_{DP}}{B\beta}\right)$
-$=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{\sqrt{B}T^{3/2}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{T\sqrt{n}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{n}\right),$
-$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{3}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}\eta_{0:t}}\right)=\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}Q(t+1)}{\sqrt{B}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}\right),$
-$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{4}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{3\beta}{\eta_{0:t}}\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}\right)=O\left(\frac{1}{T^{2}}\sum\limits_{t=0}^{T-1}\frac{Qc_{DP}\left\|\mathcal{C}\right\|_{2}}{B}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{n}\right),$
-$\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}E_{5}(t)=$ $\frac{1}{\eta_{0:T}}\sum\limits_{t=0}^{T-1}\eta_{0:t}\left(\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)=O\left(\frac{1}{T^{2}}\sum\limits_{t=0}^{T-1}\beta\left(\frac{Qc_{DP}}{B\beta}\right)^{2}\right)=O\left(\frac{Q^{2}c_{DP}^{2}}{B^{2}T\beta}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}^{2}}{B^{3/2}T^{5/2}}\right)$
-$=$ $O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}B}{n^{3/2}}{\rm polylog}\left(T,1/\delta\right)\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}c_{DP}}{n}{\rm polylog}\left(n/B,1/\delta\right)\right),$
-$E_{6}=$ $\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}=O\left(\frac{QT^{3/2}}{\left\|\mathcal{C}\right\|_{2}\sqrt{B}}\cdot\frac{1}{T^{2}}\cdot\left\|\mathcal{C}\right\|_{2}^{2}\right)=O\left(\frac{M\left\|\mathcal{C}\right\|_{2}^{2}}{\sqrt{n}}\right)=O\left(\frac{Q\left\|\mathcal{C}\right\|_{2}}{\sqrt{n}}\right).$
-Combining the asymptotic bounds, we get our final bound:
-$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})=O\left(Q\left\|\mathcal{C}\right\|_{2}\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{d}}{\varepsilon n}\text{polylog}(n/B,1/\delta)\right)\right).$
-4.4.3 Analysis of Stochastic Accelerated Gradient Descent using Independent Gradients
+We assume $p=O\left(\frac{\varepsilon^{2}n^{2}}{{\rm polylog}\left(n,1/\delta\right)}\right)$; if $p=\omega\left(\frac{\varepsilon^{2}n^{2}}{{\rm polylog}\left(n,1/\delta\right)}\right)$, then any point in $\mathcal{C}$ achieves the desired excess risk.
+Under this assumption, the upper bound on $B$ gives
+$\beta\geq 16Mn^{3/2}/B^{2}\geq 16Mn/B\geq 2MT,$
+satisfying the assumption in Claim 5.9, and also gives that $d$ satisfies the assumption in Corollary 5.11.
+Let $Q=O(L)$ in Theorem 5.1 by Corollary 5.10 and $b_{max}=\widetilde{O}_{n,1/\delta}\left(\frac{L\sqrt{p}}{\varepsilon B\beta}\right)$ by Lemma 5.6.
+We plug the tail bound of Lemma 5.6 into the high-probability error bound of Theorem 5.1, and use a failure probability of $1/T^{2}$ in both Lemma 5.6 and Theorem 5.1 (the increase in the expected error due to this low-probability event is $O(LR/T^{2})$, which is a lower order term in our error bound).
+Then we have:
+$F(\mathbf{y}_{T})-F(\mathbf{x}^{*})$
+$=\widetilde{O}_{n,1/\delta}\left(\frac{1}{T^{2}}\sum\limits_{t=1}^{T}t^{2}\left(\frac{L^{2}}{\sqrt{Bt}\beta}+\frac{Lb_{max}}{\sqrt{B}t^{1.5}}+\frac{LR}{\sqrt{B}t^{1.5}}+\frac{\beta b_{max}R}{t^{2}}+\frac{\beta b_{max}^{2}}{t^{2}}\right)+\frac{\beta R^{2}}{T^{2}}\right)$
+$=\widetilde{O}_{n,1/\delta}\left(\frac{L^{2}\sqrt{T}}{\sqrt{B}\beta}+\frac{Lb_{max}}{\sqrt{BT}}+\frac{LR}{\sqrt{BT}}+\frac{\beta b_{max}R}{T}+\frac{\beta b_{max}^{2}}{T}+\frac{\beta R^{2}}{T^{2}}\right)$
+$=\widetilde{O}_{n,1/\delta}\left(\frac{LRB^{1.5}\sqrt{T}}{n^{3/2}}+\frac{L^{2}\sqrt{p}}{\varepsilon\beta B^{1.5}\sqrt{T}}+\frac{LR}{\sqrt{BT}}+\frac{LR\sqrt{p}}{\varepsilon BT}+\frac{L^{2}p}{\varepsilon^{2}B^{2}\beta T}+\frac{n^{3/2}LR}{B^{2}T^{2}}\right).$
+$=\widetilde{O}_{n,1/\delta}\left(\frac{LRB}{n}+\frac{LR\sqrt{Bp}}{\varepsilon n^{3/2}\sqrt{T}}+\frac{LR}{\sqrt{n}}+\frac{LR\sqrt{p}}{\varepsilon n}+\frac{LRp}{\varepsilon^{2}n^{3/2}T}+\frac{LR}{\sqrt{n}}\right).$
+=$\stackrel{{\scriptstyle(\ast_{1})}}{{=}}\widetilde{O}_{n,1/\delta}\left(\frac{LR}{\sqrt{n}}+\frac{LR\sqrt{p}}{\varepsilon n}+\frac{LRp}{\varepsilon^{2}n^{2}}\right).$
+=$\stackrel{{\scriptstyle(\ast_{2})}}{{=}}\widetilde{O}_{n,1/\delta}\left(\frac{LR}{\sqrt{n}}+\frac{LR\sqrt{p}}{\varepsilon n}\right).$
+In $(\ast_{1})$, we use the fact that $B\leq\sqrt{n}$ and $T\geq\sqrt{n}$, and in $(\ast_{2})$ we use the assumption $p=O\left(\frac{\varepsilon^{2}n^{2}}{{\rm polylog}\left(n,1/\delta\right)}\right)$, which implies $p/\varepsilon^{2}n^{2}\leq\sqrt{p}/\varepsilon n$.
+5.5 Analysis of Stochastic Accelerated Gradient Descent using Independent Gradients
 Theorem 5.13.
 If we use $\nabla_{t}=\frac{1}{B}\sum_{d\in\mathcal{B}_{t}}\nabla f(\mathbf{x}_{t},d)$ instead in Accelerated-DP-SRGD , if $\eta_{t}^{2}\leq 4\eta_{0:t}$ and $\beta\geq M$ then we have:
-$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq\frac{1}{\eta_{0:T}}\sum_{i=0}^{T-1}\left[\eta_{0:t}\left(\frac{8L^{2}}{\sqrt{B}\beta}+\frac{5L\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)\right]+\frac{\beta}{\eta_{0:T}}\left\|\mathcal{C}\right\|_{2}^{2}.$
+$\mathbb{E}\left[F(\mathbf{y}_{T})\right]-F(\mathbf{x}^{*})\leq\frac{1}{\eta_{0:T}}\sum_{i=0}^{T-1}\left[\eta_{0:t}\left(\frac{8L^{2}}{\sqrt{B}\beta}+\frac{5L\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\cdot R+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)\right]+\frac{\beta}{\eta_{0:T}}R^{2}.$
 Proof.
@@
 Using the fact that $\mathbf{q}_{t}$ is mean-zero and $\mathbb{E}_{\mathbf{q}_{t}}\left[\left\|\mathbf{q}_{t}\right\|_{2}|\mathbf{x}_{t}\right]\leq\frac{2L}{\sqrt{B}}$ we then have the following bound on the potential change by plugging (30) into (10):
-$\mathbb{E}_{\mathbf{q}_{t}}\left[\Delta\Phi_{t}|\mathbf{x}_{t}\right]\leq\eta_{0:t}\left(\frac{8L^{2}}{\sqrt{B}\beta}+\frac{5L\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\cdot\left\|\mathcal{C}\right\|_{2}+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)$ (23)
+$\mathbb{E}_{\mathbf{q}_{t}}\left[\Delta\Phi_{t}|\mathbf{x}_{t}\right]\leq\eta_{0:t}\left(\frac{8L^{2}}{\sqrt{B}\beta}+\frac{5L\left\|\mathbf{b}_{t}\right\|_{2}}{\eta_{t}}+3\beta\left\|\mathbf{b}_{t}\right\|_{2}\cdot R+\frac{\beta\left\|\mathbf{b}_{t}\right\|_{2}^{2}}{\eta_{t}^{2}}\right)$ (31)
 We now proceed similarly to Theorem 5.1. ∎
-5 Empirical Evaluation
-To evaluate the practicality of DP-SRG algorithms, we perform empirical evaluations comparing DP-SRG algorithms to their analogs using batch gradients.
-We consider the multi-pass streaming setting where we can take multiple passes over the dataset in a fixed order, but this order is treated as public information, i.e.
-we cannot randomize the order and assume privacy amplification.
-This aligns with many practical settings for DP optimization, where the standard amplification techniques of Poisson sampling / uniformly random shuffling are difficult to implement [38, 16].
-In all settings, we compare to the baseline DP-MEMF algorithm (Algorithm 2) of [15] which is the SoTA in the absence of privacy amplification by sampling.
-To combine SRG with DP-MEMF , we give an algorithm DP-SRG-MEMF as Algorithm 3.
-In DP-SRG-MEMF we take the underlying first-order optimization algorithm as a black-box.
-For simplicity of presentation we state DP-SRG-MEMF for a one-dimensional optimization problem, but it is easy to extend to a multi-dimensional problem.
-As long as $\max_{j\in[b]}\left\|\sum_{i=0}^{k-1}\mathbf{C}_{:,ib+j}\right\|_{2}\leq 1$, DP-SRG-MEMF satisfies $\rho$-zCDP, i.e.
-$(\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP [15].
-[15] provides guidance for choosing a matrix $\mathbf{C}$.
-When clipping and noising are no-ops, DP-SRG-MEMF computes gradients using the SRG framework and passes them to the first-order optimization method.
-DP-SRG-MEMF privatizes the gradients by using correlated noise like DP-MEMF , but we apply the correlated noise to $\Delta_{t}$, the “update” term in SRG, rather than the gradients themselves (as was done in DP-MEMF ).
-Algorithm 2 Multi-Epoch Matrix Factorization (DP-MEMF )
-Inputs: Epochs $k$, batches $\mathcal{B}_{1},\mathcal{B}_{2},\ldots,\mathcal{B}_{b}$ of size $B$, lower-triangular matrix $\mathbf{C}\in\mathbb{R}_{\geq 0}^{kb\times kb}$ such that $\max_{j\in[b]}\left\|\sum_{i=0}^{k-1}\mathbf{C}_{:,ib+j}\right\|_{2}\leq 1$, target zCDP $\rho$, clip norm $C$.
-1:$\mathbf{Z}\sim N(0,\frac{1}{2\rho}\mathbb{I}_{kb\times kb})$
-2:for $i$ in $\{0,1,\ldots,k-1\}$ do
-3:     for $j$ in $[b]$ do
-4:         $\nabla_{ik+j}\leftarrow\frac{1}{B}\sum_{d\in\mathcal{B}_{j}}\text{clip}(\nabla\ell(\mathbf{x}_{ik+j};d),C)+(\mathbf{C}^{-1}\mathbf{Z})_{ik+j}$
-5:         Use $\nabla_{ik+j}$ as gradient in first-order optimization algorithm to update $\mathbf{x}_{ik+j}$ to $\mathbf{x}_{ik+j+1}$.
-6:     end for
-7:end for
-Algorithm 3 SRG-Multi-Epoch Matrix Factorization (DP-SRG-MEMF )
-Inputs: Epochs $k$, batches $\mathcal{B}_{1},\mathcal{B}_{2},\ldots,\mathcal{B}_{b}$ of size $B$, lower-triangular matrix $\mathbf{C}\in\mathbb{R}_{\geq 0}^{kb\times kb}$ such that $\max_{j\in[b]}\left\|\sum_{i=0}^{k-1}\mathbf{C}_{:,ib+j}\right\|_{2}\leq 1$, target zCDP $\rho$, clip norm $C$, decay sequence $\{c_{t}\}_{t\in[kb]}$.
-1:$\mathbf{Z}\sim N(0,\frac{1}{2\rho}\mathbb{I}_{kb\times kb})$
-2:for $i$ in $\{0,1,\ldots,k-1\}$ do
-3:     for $j$ in $[b]$ do
-4:         $\Delta_{ik+j}\leftarrow\frac{1}{B}\sum_{d\in\mathcal{B}_{j}}\text{clip}(\nabla\ell(\mathbf{x}_{ik+j};d)-c_{ik+j}\nabla\ell(\mathbf{x}_{ik+j-1};d),C)$
-5:         $\widetilde{\Delta}_{ik+j}\leftarrow\Delta_{ik+j}+(\mathbf{C}^{-1}\mathbf{Z})_{ik+j}$
-6:         $\nabla_{ik+j}=c_{ik+j}\nabla_{ik+j-1}+\widetilde{\Delta}_{ik+j}$
-7:         Use $\nabla_{t}+(\mathbf{C}^{-1}\mathbf{Z})_{ik+j}$ as gradient in first-order optimization algorithm to update $\mathbf{x}_{ik+j}$ to $\mathbf{x}_{ik+j+1}$.
-8:     end for
-9:end for
-5.1 Logistic Regression
-We compare DP-MEMF and DP-SRG-MEMF on multiclass logistic regression on the MNIST [20] and CIFAR10 [39] datasets.
-For both we load from tensorflow_datasets and use the pre-defined train/test split.
-We fix a batch size of 500 for each dataset, and train for $k=1$ and $6$ epochs respectively, the number of epochs we observed were needed for convergence in the non-private setting.
-We target $(0.1,10^{-6})$ and $(2,10^{-6})$-DP respectively.
-We use SGD with a standard momentum of $0.9$ as the underlying first-order algorithm for both DP-MEMF and DP-SRG-MEMF .
-For both datasets and both algorithms, we sweep the learning rate and clip norm, and pick the best performing parameters per algorithm.
-For the choice of $c_{t}$ in DP-SRG-MEMF , we fix $c_{t}=c$ for a constant $c$ less than 1.
-This is analogous to choosing $\eta_{t}=c^{t}$ in Accelerated-DP-SRGD .
-We sweep over the values of $c$ and find that $c=e^{-5/2}\approx 1/12$ works well.
-For the choice of $\mathbf{C}$, in [15], $\mathbf{C}$ is chosen to minimize $\|\mathbf{A}\mathbf{C}^{-1}\|_{F}$ under the constraint
-$\max_{j\in[b]}\left\|\sum_{i=0}^{k-1}\mathbf{C}_{:,ib+j}\right\|_{2}\leq 1$, where $\mathbf{A}$ is the lower-triangular all ones matrix.
-The motivation for this choice of $\mathbf{A}$ is that in SGD, $\mathbf{x}_{i}$ are equal to $\mathbf{x}_{0}$ plus the $i$th row of $\mathbf{A}\mathbf{G}$ where $\mathbf{G}$ is the matrix whose rows are the gradients.
-Since we add noise $\mathbf{C}^{-1}\mathbf{Z}$ to the gradients, $\|\mathbf{A}\mathbf{C}^{-1}\|_{F}$ gives some measure of the total error of the algorithm due to noise.
-However, in SGD with momentum, we instead have that $\mathbf{x}_{i}$ is equal to $\mathbf{x}_{0}$ plus the $i$th row of $\mathbf{M}\mathbf{A}\mathbf{G}$, where $\mathbf{M}$ is some function of the momentum parameter.
-Taking this one step further, if $\mathbf{D}$ is the matrix whose rows are the $\Delta_{i}$ in DP-SRG-MEMF , the $i$th row of $\mathbf{x}_{i}$ is now equal to $\mathbf{x}_{0}$ plus the $i$th row of $\mathbf{M}\mathbf{A}\mathbf{L}\mathbf{D}$, where $\mathbf{L}$ is some function of the decay parameter $c_{t}$.
-While [15] observed that using $\mathbf{C}$ in DP-MEMF optimized for $\mathbf{A}$ is usually as good as or better than optimizing for $\mathbf{M}\mathbf{A}$ in practice, it is unclear whether the same is true for DP-SRG-MEMF .
-So, we try two different choices of $\mathbf{C}$ for both algorithms: $\mathbf{C}$ optimized to minimize $\|\mathbf{A}\mathbf{C}^{-1}\|_{F}$the “ones” workload), and $\mathbf{C}$ optimized to minimize $\|\mathbf{M}\mathbf{A}\mathbf{C}^{-1}\|_{F}$ for DP-MEMF and $\|\mathbf{M}\mathbf{A}\mathbf{L}\mathbf{C}^{-1}\|_{F}$ for DP-SRG-MEMF ( the “true” workload).
-Algorithm, Workload MNIST Acc. @ $(0.1,10^{-6})$-DP CIFAR Acc. @ $(2.0,10^{-6})$-DP
-DP-MEMF , ones 82.555 $\pm$ .144 35.680 $\pm$ .166
-DP-MEMF , true 83.593 $\pm$ .123 35.272 $\pm$ .246
-DP-SRG-MEMF , ones 82.456 $\pm$ .143 36.854 $\pm$ .123
-DP-SRG-MEMF , true 83.753 $\pm$ .121 36.816 $\pm$ .113
-Non-private 91.530 $\pm$ .042 37.971 $\pm$ .076
-Figure 1: Average test accuracy on MNIST/CIFAR for the different approaches with 96% CI.
-In Figure 1 we compare the test accuracy of DP-MEMF and DP-SRG-MEMF , using the two different choices of $\mathbf{C}$, on the two tasks.
-Using the code base of [15], we are able to generate the $\mathbf{C}$ matrices in under 2 minutes per matrix on CPU.
-For each task, one training run takes less than 10 minutes to run on CPU.
-Each reported accuracy is the average over 100 training runs.
-We see that DP-SRG-MEMF always outperforms DP-MEMF .
-For CIFAR the absolute improvement is reasonably large for the given task, despite the modest choice of $c$.
-Furthermore, depending on the task, the choice of $\mathbf{C}$ can heavily affect the performance of the algorithm.
-Interestingly, even though we look at logistic regression tasks for both datasets, for MNIST choosing $\mathbf{C}$ optimized for the “true” workload is much better for both algorithms, whereas for CIFAR choosing $\mathbf{C}$ optimized for the “ones” workload is slightly better for both algorithms.
-It is an interesting question to better understand how the dataset affects the best choice of $\mathbf{C}$ in DP-MEMF and its variants.
-Note on not comparing to methods specific to convex losses: The task we consider is technically a convex problem, and some of the methods specific to convex losses are (potentially) applicable to this problem [40, 52, 28].
-We do not compare to them because of the following reason: Because we do not assume any apriori bound on the norm of the gradients, we clip the gradients to a fixed norm.
-Unfortunately, once clipping is involved, multi-class logistic loss does not any remain a well-defined function [48, Section 5.3], let alone a convex function.
-6 Conclusion and Future Work
-In this work, we demonstrated an optimal DP-SCO algorithm (Accelerated-DP-SRGD ); a single epoch first-order method that can operate on much larger batch sizes ($B\approx\sqrt{n}$ or larger) with $T=n/B$ batch gradient steps.
-It is interesting to note that our algorithm does not match the optimal non-private batch gradient complexity for SCO, achieved via accelerated SGD method [53, 41], even when $\varepsilon\to\infty$.
-In Accelerated-DP-SRGD , we use stochastic recursive gradients (SRGs) instead of the standard minibatch gradient to reduce the $\ell_{2}$-sensitivity by factor of $\approx 1/t$ for privacy purposes.
-The variance introduced by SRGs become the dominant effect even when $\varepsilon\to\infty$, and lower bounds the number of batch gradient steps the algorithm needs to make.
-An important question that evolve from our work is whether we can match the optimal non-private batch gradient steps, and still achieve optimal DP-SCO error?
-A related question is whether our rates can be matched or beaten by an algorithm that uses standard gradients, and in particular one gradient query per example.
-Such an algorithm is more likely to match the non-private number of gradient steps as $\varepsilon\rightarrow\infty$.
-In the current paper, we did not explore the applicability of the stochastic recursive gradients approach beyond standard (accelerated) SGD based methods.
-We leave the exploration of the possible gains in adaptive optimization settings like AdaGrad [22, 4] as an open problem.
-As mentioned earlier, our algorithm is applicable to non-convex settings without compromising on privacy guarantees.
-In this work, we experimented with (clipped) logistic regression as a proof-of-concept.
-We leave a detailed evaluation along the lines of [15] for future exploration.